### v1

Abstract
Determinantal varieties—the sets of bounded-rank matrices or tensors—have attracted growing interest in low‑rank optimization. The tangent cone to low-rank sets is widely studied and underpins a range of geometric methods. The second-order geometry, which encodes curvature information, is more intricate. In this work, we develop a unified framework to derive explicit formulas for both first- and second-order tangent sets to various low-rank sets, including low-rank matrices, tensors, symmetric matrices, and positive semidefinite matrices. The framework also accommodates the intersection of a low-rank set and another set satisfying mild assumptions, thereby yielding a tangent intersection rule. Through the lens of tangent sets, we establish a necessary and sufficient condition under which a nonsmooth problem and its smooth parameterization share equivalent second-order stationary points. Moreover, we exploit tangent sets to characterize optimality conditions for low-rank optimization and prove that verifying second-order optimality is NP-hard. In a separate line of analysis, we investigate variational geometry of the graph of the normal cone to matrix varieties, deriving the explicit Bouligand tangent cone, Fréchet and Mordukhovich normal cones to the graph. These results are further applied to develop optimality conditions for low-rank bilevel programs.
1 Introduction
The low-rank structure of matrix data is widely exploited in various applications to improve memory and computation efficiency markovsky2008systemide; udell2019big, which leads to the consideration of the set of bounded-rank matrices:
$\mathcal{M}_{\leq r}:=\left\{{X\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(X)\leq r}\right\},$ (1.1)
where $r\leq\min\{m,n\}$ is the rank parameter. In fact, $\mathcal{M}_{\leq r}$ is a real algebraic variety, commonly referred to as the determinantal variety harris2013algebraic, and the concepts can be extended to tensor determinantal varieties kutschan2018tangentTT; gao2025lowranktucker.
First-order variational results on $\mathcal{M}_{\leq r}$ luke2013Mordukhovichcone; cason2013iterative; hosseini2019MordukhovichClarke; olikier2022continuity, including the characterization of its tangent and normal cones, have contributed to low-rank optimization by allowing the derivation of first-order optimality conditions and supporting the development of geometric methods schneider2015Lojaconvergence; olikier2023RFDR; olikier2025PGD; olikier2025gauss. Furthermore, imposing an additional constraint $\mathcal{H}$ on low-rank optimization has attracted growing interest in the geometry of $\mathcal{M}_{\leq r}\cap\mathcal{H}$, which, however, was previously treated on a case-by-case basis for different $\mathcal{H}$ in the literature cason2013iterative; tam2017sparsesdp; li2020jotaspectral; li2023normalboundedaffine; yang2025spacedecouple.
Moreover, the second-order geometry of $\mathcal{M}_{\leq r}$ remains unclear, mainly due to two obstacles: 1) second-order analysis in essence captures curvature information bonnans2000perturbationanalysis; gfrerer2017robinsonstability, thereby appearing more involved than the first-order counterpart; 2) the determinantal variety is both nonconvex and nonsmooth olikier2022continuity; levin2023remedy, hindering the direct use of standard analysis tools.
The following two problems underpin the importance of variational analysis of $\mathcal{M}_{\leq r}$.
Optimization problem over bounded-rank matrices.
By imposing a low-rank constraint (possibly together with an additional constraint $\mathcal{H}$) on the matrix variable, the following formulation has been successfully used in a multitude of applications markovsky2008systemide; zhu2022learningmarkov,
$\min_{X\in\mathbb{R}^{m\times n}}$ $f(X)$ (1.2)
$\mathrm{s.\,t.}$ $X\in\mathcal{M}_{\leq r}\cap\mathcal{H}.$
When $\mathcal{H}=\mathbb{R}^{m\times n}$, the problem reduces to minimizing a function over bounded-rank matrices. The first-order optimality conditions are well understood cason2013iterative; hosseini2019MordukhovichClarke; levin2023remedy, whereas the second-order counterparts remain ambiguous, since the second-order geometry of $\mathcal{M}_{\leq r}$, specifically its second-order tangent set (see section 2.2), has yet to be fully explored. When the data exhibits an additional structure, the constraint evolves into $X\in\mathcal{M}_{\leq r}\cap\mathcal{H}$ with a nontrivial $\mathcal{H}$ cason2013iterative; tam2017sparsesdp; li2020jotaspectral; li2023normalboundedaffine; yang2025spacedecouple. More specifically, when $\mathcal{H}$ enforces the semidefinite constraint, problem (1.2) gives rise to a line of low-rank semidefinite programming problems boumal2016BMSDP; wang2023decomposition; tang2024feasibleSDP; levin2025effect. The coupled structure renders the geometry of the feasible region more intricate, impeding the development of a unified analysis of first- and second-order tangent sets to a general $\mathcal{M}_{\leq r}\cap\mathcal{H}$.
Bilevel programming with low-rank structure.
We consider a bilevel programming problem where the lower-level task seeks a low-rank solution, modeling applications across various fields shaban2019truncated; grangier2023LLMshift; shen2025seal; zangrando2025debora (see section 8.1):
$\min_{x\in\mathbb{R}^{q},X^{*}\in\mathbb{R}^{m\times n}}$ $\ \mathcal{L}(x,X^{*})$ (LRBP)
$\mathrm{s.\,t.}$ $\ G(x)\leq 0,$
$\ X^{*}\in\operatorname*{arg\,min}_{X\in\mathbb{R}^{m\times n}}\ \ F(x,X),$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \mathrm{s.\,t.}\ \ \ X\in\mathcal{M}_{\leq r}.$
Generally, finding a global minimizer of a function subject to the bounded-rank constraint is NP-hard gillis2011NPlowrank. Nevertheless, existing literature schneider2015Lojaconvergence; levin2023remedy; jia2023convergencePGDKL; aragon2024coderivativeNewton is able to compute an M-stationary point, at which the antigradient belongs to the Mordukhovich normal cone of $\mathcal{M}_{\leq r}$. Therefore, it is reasonable to introduce the following relaxation for (LRBP), by replacing the lower-level global optimality with the M-stationarity,
$\min_{x\in\mathbb{R}^{q},X\in\mathbb{R}^{m\times n}}$ $\ \mathcal{L}(x,X)$ (1.3)
$\mathrm{s.\,t.}$ $\ G(x)\leq 0,$
$\ -\nabla_{X}F(x,X)\in\mathrm{N}_{\mathcal{M}_{\leq r}}(X),$
where $\nabla_{X}$ denotes the partial gradient with respect to $X$. Since the Mordukhovich normal cone mapping $\mathrm{N}_{\mathcal{M}_{\leq r}}(\cdot)$ arises in the constraints, its coderivative—a generalized subdifferential mordukhovich2006variationalI—would be involved in deriving the first-order optimality condition for (1.3). Note that the normal cone mapping corresponds to the subdifferential of the indicator function associated with the set $\mathcal{M}_{\leq r}$, and thus the coderivative naturally enters the scope of the second-order variational analysis of $\mathcal{M}_{\leq r}$.
Corresponding to the above two examples, we are concerned with variational analysis of determinantal varieties from two aspects: the first- and second-order tangent sets to an array of low-rank sets, and the coderivative of the normal cone mapping $\mathrm{N}_{\mathcal{M}_{\leq r}}(\cdot)$. Next, we provide an overview of the existing literature and main challenges.
1.1 Related work and main challenges
We begin by summarizing the first-order variational results on $\mathcal{M}_{\leq r}$, which enlighten a range of geometric algorithms for low-rank optimization. The term first-order tangent set refers to the Bouligand tangent cone.
First-order variational analysis of the determinantal variety.
Research on the geometry of $\mathcal{M}_{\leq r}$, especially the associated tangent and normal cones, has flourished in low-rank optimization. Typical characterizations include the Mordukhovich normal cone luke2013Mordukhovichcone, the Bouligand tangent cone cason2013iterative; schneider2015Lojaconvergence; olikier2025fourtangentproof, and the Clarke tangent and normal cones hosseini2019MordukhovichClarke; li2019optimalitylowrank. The optimality conditions are derived: the projection of the antigradient onto the corresponding tangent cone vanishes. Furthermore, Olikier and Absil olikier2022continuity investigated the continuity of the cone mappings, which underlies the so-called apocalypse phenomenon observed in optimization problems over the bounded-rank matrices levin2023remedy.
The developed geometry has given rise to numerous algorithms for low-rank optimization, i.e., problem (1.2) with $\mathcal{H}=\mathbb{R}^{m\times n}$. One class builds on the projected gradient descent framework schneider2015Lojaconvergence; olikier2022P2GDR; olikier2025PGD, where each iteration proceeds by taking a descent step followed by a projection onto the feasible region $\mathcal{M}_{\leq r}$. Another line of work embraces a retraction-free approach schneider2015Lojaconvergence; olikier2023RFDR; olikier2024ERFDR: instead of performing projections, one adopts search directions from the so-called restricted tangent cone to $\mathcal{M}_{\leq r}$ (see olikier2023RFDR) and iterates along straight lines. The third class leverages a smooth parameterization of $\mathcal{M}_{\leq r}$ by constructing a manifold $\overline{\mathcal{M}}$ and a mapping $\phi$ such that $\phi(\overline{\mathcal{M}})=\mathcal{M}_{\leq r}$, thereby implementing optimization algorithms over the smooth manifold instead of the nonsmooth determinantal variety khrulkov2018desingularization; rebjock2024boundedrank; levin2023remedy; levin2025effect; olikier2025gauss.
When we introduce an additional constraint set $\mathcal{H}$ to study the coupled region $\mathcal{M}_{\leq r}\cap\mathcal{H}$, the geometry becomes more complicated. Specifically, for the case $m=n$ and $\mathcal{H}=\mathcal{S}(n)$, the set of $n\times n$ symmetric matrices, Tam et al. tam2017sparsesdp established the Mordukhovich normal cone to $\mathcal{M}_{\leq r}\cap\mathcal{H}$. Subsequently, Li et al. li2020jotaspectral provided formulations for the Fréchet normal cones when $\mathcal{H}$ represents the intersection of $\mathcal{S}(n)$ with the closed unit Frobenius ball, the symmetric box, or the spectrahedron. Moreover, Levin et al. levin2025effect explicitly computed the Bouligand tangent cone to $\mathcal{S}^{+}_{\leq r}(n):=\mathcal{M}_{\leq r}\cap\mathcal{S}^{+}(n)$, where $\mathcal{S}^{+}(n)$ denotes the closed convex cone of all positive semidefinite matrices in $\mathcal{S}(n)$. The above advancements require the matrix to be square and symmetric. When $m\neq n$ breaks this symmetry, more challenges arise, and new techniques tailored to the geometry of the coupled region are needed. A closed-form expression of the Bouligand tangent cone to the intersection of $\mathcal{M}_{\leq r}$ and the Frobenius sphere was developed in cason2013iterative. Recently, Li and Luo li2023normalboundedaffine characterized the Fréchet normal cone to $\mathcal{M}_{\leq r}\cap\mathcal{H}$ with $\mathcal{H}$ as an affine manifold. In addition, Yang et al. yang2025spacedecouple derived the Bouligand tangent cone and the Fréchet normal cone to $\mathcal{M}_{\leq r}\cap\mathcal{H}$ for $\mathcal{H}=\{X\in\mathbb{R}^{m\times n}\mid h(X)=0\}$ with a differentiable and orthogonally invariant mapping $h$, i.e., $h(X)=h(XQ)$ for all orthogonal matrices $Q$.
Notice that second-order results of the determinantal variety remain limited. Therefore, we then revisit relevant developments in second-order variational analysis for general sets; details are referred to rockafellar2009variationalanalysis; mordukhovich2024secondorder and the references therein. More specifically, regarding certain structured sets, we discuss some existing techniques for deriving the associated second-order tangents set and coderivatives of normal cone mappings.
Second-order tangent set.
Analogous to how the second-order derivative of a mapping refines the linear approximation into the quadratic one, the second-order tangent set to a given set provides a more accurate local approximation than the Bouligand tangent cone bonnans2000perturbationanalysis; chen2019exactSOC. Therefore, it serves as an important tool for analyzing optimality conditions bonnans1999parabolicset; gfrerer2022secondnonconvex, metric subregularity gfrerer2011subregularity, and system stability gfrerer2017robinsonstability. More relevant to this work, Levin et al. levin2025effect tackled optimization problems over a nonsmooth set (e.g., the determinantal variety) by studying the parameterization technique that recasts the problem on a smooth manifold. Specifically, the second-order tangent set to the manifold is exploited to establish the equivalence between the smooth and nonsmooth problems.
There are two classes of sets to which the second-order tangent set is well understood. The first class considers sets of the form $\mathcal{X}=\{X\in\mathbb{R}^{q}\mid h(X)\in\mathcal{C}\}$, where $h$ is a smooth mapping and $\mathcal{C}$ is a closed set. A characterization of the second-order tangent set to $\mathcal{X}$ was given in (rockafellar2009variationalanalysis, Proposition 13.13). As a special case, when $\mathcal{X}$ is a smooth manifold, an interpretation through the lens of smooth curves was provided in levin2025effect. The second class consists of sets of the form $\mathcal{X}=\{X\in\mathbb{R}^{q}\mid h(X)\leq 0\}$, where $h$ is convex. Under the Slater condition, Bonnans et al. computed the second-order tangent set by relating it to the second-order subdifferential of $h$ (bonnans1999parabolicset, Proposition 2.1). Applying this theory, the second-order tangent sets to the convex cone $\mathcal{S}^{+}(n)$ and to the second-order cone [^1] The second-order cone is defined as $\mathcal{X}:=\{(X_{1},X_{2})\in\mathbb{R}\times\mathbb{R}^{q}:\left\|{X_{2}}\right\|_{2}\leq X_{1}\}$. (SEC) were characterized in (bonnans2000perturbationanalysis, Example 3.40) and (bonnans2005perturbation, Lemma 27), respectively.
However, when the set $\mathcal{X}$ of interest is both nonsmooth and nonconvex, the analysis should resort to the specific structure of $\mathcal{X}$. An example of $\mathcal{X}$ is the SEC complementarity set. In this case, Chen et al. chen2019exactSOC noticed that the metric projection operator onto the SEC is well-defined and admits second-order directional derivatives, based on which they gave the exact formula for the second-order tangent set to $\mathcal{X}$.
Coderivative of normal cone mappings.
Given a set $\mathcal{X}$, the coderivative of the associated normal cone mapping, denoted by $\mathrm{D}^{*}\mathrm{N}_{\mathcal{X}}$, can be treated as the second-order subdifferential of the indicator function of $\mathcal{X}$ mordukhovich2015secondvarconic. The concept of coderivative plays a pivotal role in investigating the stability and sensitivity of variational systems poliquin1998tilt; dontchev2009implicitsolutionmappings, regularity properties of set-valued mappings dontchev2009implicitsolutionmappings, and optimality conditions of bilevel programming problems ding2014SDCMPCC; dempe2018optimality. As pointed out in chieu2017coderivativeweakcondition, computing explicitly the coderivative of a given set-valued mapping is generally a demanding task.
When the considered $\mathcal{X}$ is a polyhedral convex set, Dontchev and Rockafellar dontchev1996polyhedralconvex characterized the associated $\mathrm{D}^{*}\mathrm{N}_{\mathcal{X}}$; and then a line of works gfrerer2015weakest; gfrerer2016computationgeneralizedderivatives; chieu2017coderivativeweakcondition extended the results to the case of $\mathcal{X}=\{X\in\mathbb{R}^{q}\mid h(X)\in\mathcal{C}\}$ with a twice continuously differentiable mapping $h$ and a polyhedral convex set $\mathcal{C}$ satisfying some qualification conditions. Additionally, for $\mathcal{X}=\mathcal{S}^{+}(n)$ as a closed convex cone, the directional derivative of the projection operator onto $\mathcal{S}^{+}(n)$ was exploited to obtain the explicit formula of $\mathrm{D}^{*}\mathrm{N}_{\mathcal{S}^{+}(n)}$ ding2014SDCMPCC; wu2014SDCMPCC. This identification facilitates the derivation of optimality conditions for bilevel programming problems where $\mathcal{S}^{+}(n)$ appears as a constraint set in the lower-level problem dempe2018optimality.
In summary, the first-order geometry of low-rank sets $\mathcal{M}_{\leq r}\cap\mathcal{H}$ has been treated in the literature only on a case-by-case basis for different choices of $\mathcal{H}$—there is currently no framework that both unifies existing results and guides new developments. In addition, the second-order analysis of $\mathcal{M}_{\leq r}$ is even more challenging. Essentially, the determinantal variety is nonconvex and nonsmooth: the recent work olikier2022continuity showed that the first-order geometric objects associated with $\mathcal{M}_{\leq r}$—such as the Mordukhovich normal cone mapping—are discontinuous, further underscoring the difficulty of developing second-order theory. Moreover, the lack of regularity precludes the straightforward employment of existing techniques to low-rank sets. For instance, in the cases of the aforementioned SEC and $\mathcal{S}^{+}(n)$, analyzing the directional derivative of the associated projection operator is able to identify the second-order tangent set or the coderivative of the normal cone mapping. However, for the determinantal variety, the projection $\mathcal{P}_{\mathcal{M}_{\leq r}}(X)$ becomes set-valued when the nonzero $r$-th and $(r+1)$-th largest singular values of $X$ coincide. This non-uniqueness breaks the differentiability of the projection operator, thereby restricting the translation of the existing projection-based techniques.
1.2 Contributions
In this paper, we investigate variational properties of low-rank sets. The contributions unfold along two lines: I. a unified framework for developing first- and second-order tangent sets to low-rank sets; II. the geometry of the graph of the normal cone mapping induced by the determinantal variety, which facilitates the characterization of the associated coderivatives; see the roadmap in Fig. 1.
Problem on intersection of sets $\min\ \ f(X)\ \ \ \mathrm{s.\,t.}\ \ X\in\mathcal{M}\cap{\mathcal{K}}$ I. Tangent sets and optimization TableTheorem 3.1: tangent sets to $\mathcal{M}$
TableTheorem 4.1: intersection rule for $\mathcal{M}\cap\mathcal{K}$
TheoremTable 1: applications to low-rank sets
Second-orderoptimality (SOC)Sufficient and necessarycondition for $2\!\Rightarrow\!2$NP-hardness ofverifying SOCParameterizations yield SOC on $\mathcal{M}_{\leq r}$only at rank-$r$ pointsPropositions 6-7Theorem 5.1Theorems 6.1-6.2Propositions 8-9 Low-rank bilevel programming problem (LRBP) $\Downarrow$ Relaxed version involving $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$
as a constraint (M-LRBP)
II. Geometry of $\mathrm{gph}\mathrm{N}_{\!\mathcal{M}_{\leq r}}$ and bilevel programBouligand tangent cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$}(Theorem 7.1)Fréchet normal cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$}(Corollaries 1-2)Mordukhovich normal cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$}(Theorem 7.2)Optimality conditions for (M-LRBP)(Proposition 11)
Figure 1: Roadmap of the contributions. Part I: sections 3-6; Part II: sections 7-8.
I. Tangent sets and optimization. The main goal of this part is to identify the tangent sets to low-rank sets, thereby providing the optimality analysis for rank-constrained optimization problems.
We endeavor to relate the tangent sets of the determinantal variety to appropriate (generalized) differentials of mappings that capture the low-rank structure. To this end, we consider the singular value mappings which are non-negative and are ordered in a non-increasing fashion: $\sigma_{1}(X)\geq\sigma_{2}(X)\geq\cdots\geq\sigma_{\min\{m,n\}}(X)\geq 0$. Subsequently, $\mathcal{M}_{\leq r}$ can be treated as the zeros of the $(r+1)$-th singular value mapping $\sigma_{r+1}$:
$\mathcal{M}_{\leq r}=\{X\in\mathbb{R}^{m\times n}\mid\sigma_{r+1}(X)=0\}.$
Specifically, the mapping $\sigma_{r+1}$ comes with two benefits: 1) it is Lipschitz continuous by Weyl’s inequality weyl1912inequality; 2) it satisfies the error bound condition according to the truncated SVD, namely, $\operatorname{dist}(X,\mathcal{M}_{\leq r})\leq(\min\{m,n\}-r)^{1/2}\sigma_{r+1}(X)$ for any $X\in\mathbb{R}^{m\times n}$. We generalize these two conditions into Assumption 1, under which Theorem 3.1 is established: given a general set $\mathcal{M}$ realized as the zeros of a mapping $c$, the first- and second-order tangent sets to $\mathcal{M}$ coincide with the zeros of the first- and second-order directional differentials of $c$, respectively. Applying the developed results to $\mathcal{M}=\mathcal{M}_{\leq r}$ and substituting the directional differentials of $\sigma_{r+1}$ given in zhang2013secondordersingular, we recover the first-order geometry of $\mathcal{M}_{\leq r}$ schneider2015Lojaconvergence, and compute explicitly the second-order tangent set to $\mathcal{M}_{\leq r}$ in Proposition 1.
Interestingly, the perspectives can be extended to broader scenarios: we view low-rank sets of tensors, symmetric matrices, and positive semidefinite (PSD) matrices as the zeros of singular or eigenvalue mappings (see (4.3), (4.12), and (4.16), respectively), which allows us to invoke Theorem 3.1 to obtain the associated tangent sets.
Moreover, when an additional structured set $\mathcal{K}$ is imposed, we consider the geometry of the intersection $\mathcal{M}\cap\mathcal{K}$, and employ the idea of smooth parameterization levin2025effect, which introduces a smooth manifold $\overline{\mathcal{M}}$ embedded in another Euclidean space $\mathcal{W}$ and a smooth mapping $\phi$ with $\phi(\overline{\mathcal{M}})=\mathcal{M}$; see Fig. 2 for an illustration. We pull back $\mathcal{K}$ through $\phi$ to obtain $\overline{\mathcal{K}}:=\phi^{-1}(\mathcal{K})$, and then take into account the intersection $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ in the “auxiliary space” $\mathcal{W}$. Theorem 4.1 reveals that, under Assumption 2, some benign properties of the lift $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ can be inherited by $\mathcal{M}\cap\mathcal{K}$, therefore decoupling the computation of the tangent sets to $\mathcal{M}\cap\mathcal{K}$ into the intersection of those to each component,
$\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)$ $=\mathrm{T}_{\mathcal{M}}(X)\cap\mathrm{T}_{\mathcal{K}}(X),$
$\mathrm{T}^{2}_{\mathcal{M}\cap\mathcal{K}}(X;\eta)$ $=\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{K}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X).$
Theorem 4.1 serves as an extension of Theorem 3.1 to the intersection of sets.
Applications of Theorems 3.1 and 4.1 to the (structured) low-rank sets—realized with different choices of $\mathcal{M}$ and $\mathcal{K}$—are summarized in Table 1. Notably, the Bouligand tangent cones to hierarchical Tucker varieties and to $\mathcal{M}_{\leq r}\cap\mathcal{H}$ with hyperbolic $\mathcal{H}$ are new results, and all the second-order tangent sets reported in the table are also novel, to our knowledge.
Indeed, first- and second-order tangent sets play a crucial role in bridging optimization landscapes. Specifically, as demonstrated in Fig. 2, the smooth parameterization $(\overline{\mathcal{M}},\phi)$ reformulates the original nonsmooth problem (P) as a smooth optimization problem (P-M). In light of levin2025effect, Theorem 5.1 exploits the structures of tangent sets to $\overline{\mathcal{M}}$ and $\mathcal{M}$, and provides a sufficient and necessary condition to characterize when second-order stationary points of (P-M) map to those of (P), which partially answers a question raised in (levin2025effect, §6).
[…]
Figure 2: Illustration of optimization through a smooth parameterization, where $\mathcal{M}$ is a possibly nonsmooth set, $\overline{\mathcal{M}}$ is a smooth manifold, and $\phi$ is a smooth mapping between the two Euclidean spaces $\mathcal{W}$ and $\mathcal{E}$.
Finally, the developed framework is applied to low-rank optimization in section 6. Specifically, substituting the characterizations of tangent sets to the low-rank sets, we derive in Proposition 6 the first- and second-order optimality conditions for the low-rank problem (1.2). Building on these results, we reveal in Theorems 6.1 and 6.2 that verifying second-order optimality for low-rank optimization is NP-hard in general. Nevertheless, in certain special cases, it is still possible to achieve second-order stationarity on $\mathcal{M}_{\leq r}$ at rank-$r$ points, by adopting specific smooth parameterizations; see Propositions 8 and 9. Therefore, the NP-hardness identified in Theorems 6.1 and 6.2 essentially stems from the singular points on the determinantal variety.
Table 1: Summary of Bouligand tangent cones (first-order) and second-order tangent sets to the low-rank sets.
Set Format First-order Second-order
    $\mathcal{M}_{\phantom{\leq r}}$ Assumption 1 Theorem 3.1 Theorem 3.1
    $\mathcal{M}_{\leq r}$ (1.1) matrix cason2013iterative; schneider2015Lojaconvergence Proposition 1
    $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ (4.1) hierarchical Tucker Proposition 2 Proposition 2
    $\mathcal{M}^{\mathrm{tc}}_{\leq\mathbf{r}}$ (A.1) Tucker gao2025lowranktucker Proposition 2
    $\mathcal{M}^{\mathrm{tt}}_{\leq\mathbf{r}}$ (A.2) tensor train kutschan2018tangentTT Proposition 2
    $\mathcal{S}_{\leq r}(n)$ (4.11) symmetric matrix li2020jotaspectral Proposition 3
    $\mathcal{S}^{+}_{\leq r}(n)$ (4.15) PSD matrix levin2025effect Proposition 4
Intersection of sets Structured set First-order Second-order
$\mathcal{M}\cap\mathcal{K}$ Assumption 2 Theorem 4.1 Theorem 4.1
$\mathcal{M}_{\leq r}\cap\mathcal{H}$ $\mathcal{H}=\{X\in\mathbb{R}^{m\times n}\mid\mathcal{A}(X)=b\}$ li2023normalboundedaffine Appendix C.1
$\mathcal{M}_{\leq r}\cap\mathcal{H}$ $\mathcal{H}$ is orthogonally invariant (C.11) yang2025spacedecouple Appendix C.2
$\mathcal{M}_{\leq r}\cap\mathcal{H}$ $\mathcal{H}$ is hyperbolic (C.14) Appendix C.3 Appendix C.3
$\mathcal{S}_{\leq r}(n)\cap\mathcal{U}$ $\mathcal{U}=\{X\in\mathcal{S}(n)\mid\|X\|^{2}_{\mathrm{F}}=1\}$ Appendix D.2 Appendix D.2
$\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ $\mathcal{U}=\{X\in\mathcal{S}(n)\mid\mathcal{A}(X)=b\}$ levin2025effect Appendix D.3
II. Geometry of $\bm{\mathrm{gph}}\mathrm{N}_{\mathcal{M}_{\leq r}}$ and bilevel program. We investigate, in section 7, the geometry of the graph of the Mordukhovich normal cone mapping associated with the determinantal variety $\mathcal{M}_{\leq r}$, along the right side of Fig. 1. In detail, the Bouligand tangent cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ is characterized in Theorem 7.1, and then we take the polar operation to obtain the Fréchet normal cone in Corollaries 1-2. Consequently, in Theorem 7.2, the Mordukhovich normal cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ is identified as the outer limit of the developed Fréchet normal cone, which induces the calculation of the coderivative of the Mordukhovich normal cone mapping $X\mapsto\mathrm{N}_{\mathcal{M}_{\leq r}}(X)$.
The geometry of $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ is important in analyzing bilevel programs with low-rank constraints at the lower level. Specifically, relaxing the bilevel problem (LRBP), we propose a formulation (M-LRBP) equivalent to (1.3), which involves $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ in the constraints. Consequently, Proposition 11 applies the results of Theorem 7.2 to give an optimality condition for problem (M-LRBP).
1.3 Organization
Section 2 reviews some notation and preliminaries from variational analysis. In section 3, we present a rule to compute the first- and second-order tangent sets to a general set, and then apply the results to the determinantal variety $\mathcal{M}_{\leq r}$. Section 4 extends the analysis to more structured low-rank sets, including sets of low-rank matrices, tensors, symmetric matrices, PSD matrices, and intersections of sets; details are organized in appendices. Section 5 unveils that the concept of tangent sets indeed bridges optimization landscapes under smooth parameterization. The developed framework is then applied to low-rank optimization in section 6. In section 7, we investigate the geometry of the graph of the normal cone to the determinantal variety, which is applied to give an optimality condition in low-rank bilevel programs; see section 8. Finally, we draw the conclusion in section 9.
2 Notation and preliminaries
This section summarizes the notation and reviews some preliminaries in variational analysis. For further background and references, see bonnans2000perturbationanalysis; rockafellar2009variationalanalysis.
2.1 Notation
Let $\mathrm{St}(n,p)=\{X\in\mathbb{R}^{n\times p}\mid X^{\top}X=I_{p}\}$ be the Stiefel manifold, $\mathcal{O}(n)=\mathrm{St}(n,n)$ be the orthogonal group, $\mathcal{M}_{s}=\{X\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(X)=s\}$ be the set of fixed-rank matrices, and $\mathrm{Skew}(n)=\{\Omega\in\mathbb{R}^{n\times n}\mid\Omega^{\top}=-\Omega\}$ be the set of skew-symmetric matrices. We use $\operatorname{Diag}(x)$ to denote the diagonal matrix with diagonal entries given by the vector $x$, and $\operatorname{diag}(X)$ to denote the vector extracting the diagonal from a square matrix $X$. Given a smooth manifold $\mathcal{X}$, $\mathrm{T}_{\mathcal{X}}(X)$ denotes the tangent space at $X\in\mathcal{X}$. Given a mapping $F:\mathcal{X}_{1}\to\mathcal{X}_{2}$ between two manifolds, $\mathrm{D}F_{X}:\mathrm{T}_{\mathcal{X}_{1}}(X)\to\mathrm{T}_{\mathcal{X}_{2}}(F(X))$ denotes the differential of $F$ at $X\in\mathcal{X}_{1}$. The standard inner product in an Euclidean space is given by $\langle{X_{1},X_{2}}\rangle:=\mathrm{tr}(X_{1}^{\top}X_{2})$. The distance from $Y$ to $\mathcal{X}$ is defined as $\operatorname{dist}(Y,\mathcal{X})=\inf_{X\in\mathcal{X}}\|X-Y\|$. Let $\mathcal{P}_{\mathcal{X}}$ denote the projection onto the set $\mathcal{X}$. Given a matrix $\in\mathrm{St}(n,r), VV_{\bot}$ is an orthogonal complement of it in the sense of $V\ V_{\bot}]\in\mathcal{O}(n)$. Throughout the paper, whenever the rank of a matrix $X\in\mathbb{R}^{m\times n}$ is explicitly specified—e.g., $\operatorname{rank}(X)=s$ or $X\in\mathcal{M}_{s}$—and an SVD of $X$ is invoked, we use the following conventions: for the compact SVD, $X=U\varSigma V^{\top}$ with $U\in\mathrm{St}(m,s)$, $\varSigma\in\mathbb{R}^{s\times s}$, and $V\in\mathrm{St}(n,s)$; for the full SVD, $X=\bar{U}[\bar{\varSigma}\ 0]\bar{V}^{\top}$ with $\bar{U}\in\mathcal{O}(m)$, $\bar{\varSigma}\in\mathbb{R}^{m\times m}$, and $\bar{V}\in\mathcal{O}(n)$ (when $m\leq n$). Based on the compact SVD, the Moore–Penrose inverse of $X$ is $X^{\dagger}=V\varSigma^{-1}U^{\top}$. Additionally, we use $\sigma_{i}(X)$ to denote the $i$-th largest singular value of $X\in\mathbb{R}^{m\times n}$; when $X$ is symmetric, i.e., $X\in\mathcal{S}(n)$, we use $\lambda_{i}(X)$ to denote its $i$-th largest eigenvalue. Matrices or vectors are denoted by usual roman letters (e.g., $X$, $\eta$), while higher-order tensors are written in boldface (e.g., $\mathbf{X}$, $\bm{\eta}$).
2.2 Background in variational analysis
Let $\mathcal{X}$ be a nonempty and closed subset of a finite-dimensional Euclidean space $\mathcal{E}$. The Bouligand tangent cone to $\mathcal{X}$ at a point $X\in\mathcal{X}$ is
$\mathrm{T}_{\mathcal{X}}(X)=\left\{\eta\in\mathcal{E}\mid\exists t_{i}\to 0,\,\text{such that}\,\operatorname{dist}(X+t_{i}\eta,\mathcal{X})=o(t_{i})\right\}.$ (2.1)
The second-order tangent set to $\mathcal{X}$ at $X$ in the direction $\eta\in\mathrm{T}_{\mathcal{X}}(X)$ is defined by
$\mathrm{T}^{2}_{\mathcal{X}}(X;\eta)=\{\zeta\in\mathcal{E}\mid\exists t_{i}\to 0,\,\text{such that}\,\operatorname{dist}(X+t_{i}\eta+\frac{1}{2}t_{i}^{2}\zeta,\mathcal{X})=o(t_{i}^{2})\}.$ (2.2)
In fact, the tangent cone $\mathrm{T}_{\mathcal{X}}(X)$ provides a linear approximation of the set $\mathcal{X}$ around $X$, whereas the second-order tangent set further captures curvature information and thus facilitates a more precise local approximation, as illustrated by the example in Fig. 3.
XXFirst-order approximation Second-order approximation $\mathcal{X}$
Figure 3: Illustration of approximations to the set $\mathcal{X}=\{(x,y)\in\mathbb{R}^{2}\mid y=x^{2}+x^{3}\}$ around $X=(0,0)$. The blue dashed line corresponds to the first-order path $\gamma_{1}(t)=X+t\eta$ with $\eta=(1,0)\in\mathrm{T}_{\mathcal{X}}(X)=\mathbb{R}\times\{0\}$. The red parabola represents the second-order path $\gamma_{2}(t)=X+t\eta+\frac{1}{2}t^{2}\zeta$ with $\zeta=(0,2)\in\mathrm{T}^{2}_{\mathcal{X}}(X;\eta)=\{(\zeta_{1},2)\mid\zeta_{1}\in\mathbb{R}\}$, capturing the curvature of $\mathcal{X}$ and yielding a quadratic approximation of $\mathcal{X}$.
Taking the polar operation on $\mathrm{T}_{\mathcal{X}}(X)$ yields the Fréchet normal cone (also known as the regular normal cone),
$\hat{\mathrm{N}}_{\mathcal{X}}(X):=\left({\mathrm{T}_{\mathcal{X}}(X)}\right)^{\circ}=\left\{{Y\in\mathcal{E}\mid\langle Y,\eta\rangle\leq 0,\ \text{for all}\ \eta\in\mathrm{T}_{\mathcal{X}}(X)}\right\}.$
The Mordukhovich normal cone can be obtained as the outer limit of $\hat{\mathrm{N}}_{\mathcal{X}}$, i.e.,
$\mathrm{N}_{\mathcal{X}}(X)=\{Y\in\mathcal{E}\mid\exists\,X_{k}\rightarrow X,\ Y_{k}\rightarrow Y\ \text{such that}\ X_{k}\in\mathcal{X},\ Y_{k}\in\hat{\mathrm{N}}_{\mathcal{X}}(X_{k})\},$ (2.3)
which is also referred to as the limiting normal cone. When $\mathcal{X}$ is a smooth manifold, the two normal cones coincide with the normal space.
We revisit the basics for tangent and normal sets to the union (or intersection) of finite sets; see rockafellar2009variationalanalysis; lee2012manifolds. Given $X\in\bigcup_{i=1}^{d}{\mathcal{X}_{i}}$, it holds that
$\mathrm{T}_{\bigcup\nolimits_{i=1}^{d}{\mathcal{X}_{i}}}(X)$ $=\bigcup\nolimits_{i=1}^{d}\mathrm{T}_{\mathcal{X}_{i}}(X),$ (2.4)
$\mathrm{T}^{2}_{\bigcup\nolimits_{i=1}^{d}{\mathcal{X}_{i}}}(X;\eta)$ $=\bigcup\nolimits_{i=1}^{d}\mathrm{T}^{2}_{\mathcal{X}_{i}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\bigcup\nolimits_{i=1}^{d}{\mathcal{X}_{i}}}(X),$
where we denote $\mathrm{T}_{\mathcal{X}_{i}}(X)=\emptyset$ and $\mathrm{T}^{2}_{\mathcal{X}_{i}}(X;\eta)=\emptyset$ if $X\notin\mathcal{X}_{i}$. Moreover, given $X\in\mathcal{X}_{1}\cap\mathcal{X}_{2}$, it has
$\mathrm{T}_{\mathcal{X}_{1}\cap\mathcal{X}_{2}}(X)\subseteq\mathrm{T}_{\mathcal{X}_{1}}(X)\cap\mathrm{T}_{\mathcal{X}_{2}}(X)\quad\text{and}\quad\ \mathrm{N}_{\mathcal{X}_{1}\cap\mathcal{X}_{2}}(X)\supseteq\mathrm{N}_{\mathcal{X}_{1}}(X)+\mathrm{N}_{\mathcal{X}_{2}}(X).$ (2.5)
[…]
$\mathrm{T}_{\mathcal{X}_{1}\cap\mathcal{X}_{2}}(X)=\mathrm{T}_{\mathcal{X}_{1}}(X)\cap\mathrm{T}_{\mathcal{X}_{2}}(X)\quad\text{and}\quad\ \mathrm{N}_{\mathcal{X}_{1}\cap\mathcal{X}_{2}}(X)=\mathrm{N}_{\mathcal{X}_{1}}(X)+\mathrm{N}_{\mathcal{X}_{2}}(X).$ (2.6)
The attention then turns to directional derivatives of a mapping $h:\mathcal{E}_{1}\to\mathcal{E}_{2}$. Specifically, we say that $h$ is directionally differentiable at $X\in\mathcal{X}$ in a direction $\eta\in\mathcal{E}_{1}$ if the following limit exists,
$h^{\prime}(X;\eta):=\lim_{t\to 0}\frac{h(X+t\eta)-h(X)}{t}.$
If $h$ is directionally differentiable at $X$ in any direction $\eta\in\mathcal{E}_{1}$, then $h$ is termed directionally differentiable at $X$. Additionally, its parabolic second-order directional derivative is defined by
$h^{\prime\prime}(X;\eta,\zeta):=\lim_{t\to 0}\frac{h(X+t\eta+\frac{1}{2}t^{2}\zeta)-h(X)-th^{\prime}(X;\eta)}{\frac{1}{2}t^{2}},$
provided that the above limit exists. When $h$ is locally Lipschitz around $X$, we have
$h\left({X+t\eta+o(t)}\right)=h(X)+th^{\prime}(X;\eta)+o(t),$ (2.7)
$h(X+t\eta+\frac{1}{2}t^{2}\zeta+o(t^{2}))=h(X)+th^{\prime}(X;\eta)+\frac{1}{2}t^{2}h^{\prime\prime}(X;\eta,\zeta)+o(t^{2}).$ (2.8)
2.3 Variational geometry of low-rank sets
The first-order geometry of the low-rank sets is well developed; see luke2013Mordukhovichcone; vandereycken2013lowrankcompletion; schneider2015Lojaconvergence. As a fixed-rank layer of $\mathcal{M}_{\leq r}=\bigcup_{s=0}^{s=r}\mathcal{M}_{s}$, $\mathcal{M}_{s}$ is indeed an analytic manifold. Given $X\in\mathcal{M}_{s}$ with the singular value decomposition $X=U\varSigma V^{\top}$, the tangent and normal spaces are outlined below,
$\mathrm{T}_{\mathcal{M}_{s}}(X)=\left\{{[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&0\\ \end{matrix}\right][V\ V_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times(n-s)},\\ W_{3}\in\mathbb{R}^{(m-s)\times s}\end{array}\right.}\right\},$ (2.11)
$\mathrm{N}_{\mathcal{M}_{s}}(X)=\left\{{U_{\bot}JV_{\bot}^{\top}\mid J\in\mathbb{R}^{(m-s)\times(n-s)}}\right\}.$ (2.12)
Assembling the layers yields the bounded-rank set $\mathcal{M}_{\leq r}$, with its tangent and normal cones at $X\in\mathcal{M}_{s}$ formulated as follows,
$\mathrm{T}_{\mathcal{M}_{\leq r}}(X)=\mathrm{T}_{\mathcal{M}_{s}}(X)+\left\{{R\in\mathrm{N}_{\mathcal{M}_{s}}(X)\mid\operatorname{rank}(R)\leq r-s}\right\},$ (2.13)
$\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}}(X)=\begin{cases}\mathrm{N}_{\mathcal{M}_{s}}(X),\,\,\quad\ \mathrm{if}\,\,s=r,\\ \left\{0\right\},\,\,\quad\quad\quad\ \,\mathrm{if}\,\,s<r,\end{cases}$ (2.14)
$\mathrm{N}_{\mathcal{M}_{\leq r}}(X)=\left\{{R\in\mathrm{N}_{\mathcal{M}_{s}}(X)\mid\operatorname{rank}(R)\leq\min\{m,n\}-r}\right\}.$ (2.15)
Let $P_{W}:=WW^{\top}$ and $P_{W_{\bot}}:=I-P_{W}$ for any $W\in\mathrm{St}(n,s)$. The projection of $E\in\mathbb{R}^{m\times n}$ onto $\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ is given by
$\mathcal{P}_{\mathrm{T}_{\mathcal{M}_{\leq r}}(X)}(E)=P_{U}EP_{V}+P_{U}EP_{V_{\bot}}+P_{U_{\bot}}EP_{V}+\mathcal{P}_{\mathcal{M}_{r-s}}\left({P_{U_{\bot}}EP_{V_{\bot}}}\right).$
3 First- and second-order tangent sets
As introduced in section 1.1, the first- and second-order tangent sets play a key role in developing optimality conditions for constrained problems bonnans2000perturbationanalysis; chen2019exactSOC. Specifically, regarding the determinantal variety $\mathcal{M}_{\leq r}$, its first-order geometry is well understood luke2013Mordukhovichcone; vandereycken2013lowrankcompletion; schneider2015Lojaconvergence, while the second-order counterpart remains unclear. Additionally, we note that the rank function used to define $\mathcal{M}_{\leq r}$ in (1.1) is discontinuous, through which implementing the second-order variational analysis is not straightforward. To circumvent the challenges, we turn to the following perspective,
$\mathcal{M}_{\leq r}=\{X\in\mathbb{R}^{m\times n}\mid\sigma_{r+1}(X)=0\}.$ (3.1)
[…]
Although the mapping $\sigma_{r+1}$ still exhibits both nonsmoothness and nonconvexity, we notice that
  • 1)
    it is locally Lipschitz, since $|\sigma_{r+1}(X)-\sigma_{r+1}(X+\Delta)|\leq\left\|{\Delta}\right\|_{2}$ for any $\Delta\in\mathbb{R}^{m\times n}$ by Weyl’s inequality weyl1912inequality;
  • 2)
    and according to the truncated SVD, $\sigma_{r+1}$ can control the distance of points to the set $\mathcal{M}_{\leq r}$, i.e., $\operatorname{dist}(X,\mathcal{M}_{\leq r})\leq(\min\{m,n\}-r)^{1/2}\sigma_{r+1}(X)$ for any $X\in\mathbb{R}^{m\times n}$.
Motivated by the two observations, we distill the ideas into a rule for computing tangent sets to a general set, as presented in section 3.1. Subsequently, the developed results, together with several properties of singular value mappings reviewed in section 3.2, are applied to the determinantal variety in section 3.3.
3.1 Tangent sets to a general set
Given an Euclidean space $\mathbb{R}^{q}$ and two mappings $c_{1}:\mathbb{R}^{q}\to\mathbb{R}^{n_{1}}$ and $c_{2}:\mathbb{R}^{q}\to\mathbb{R}^{n_{2}}$, define the set $\mathcal{M}\subseteq\mathbb{R}^{q}$ by
$\mathcal{M}:=\{\tilde{X}\in\mathbb{R}^{q}\mid\ c_{1}(\tilde{X})=0,\ c_{2}(\tilde{X})\leq 0\},$
where the relations “==” and $\leq$” are understood component-wise. Around a point $X\in\mathcal{M}$, we introduce a regularity assumption as follows.
Assumption 1
There exists a neighborhood $\mathcal{B}$ of $X\in\mathcal{M}$ and a constant $\rho>0$ satisfying the following two conditions.
  • (i)
    Both $c_{1}$ and $c_{2}$ are Lipschitz in $\mathcal{B}$;
  • (ii)
    For all $\tilde{X}\in\mathcal{B}$, it holds that $\operatorname{dist}(\tilde{X},\mathcal{M})\leq\rho\|(c_{1}(\tilde{X}),\,[c_{2}(\tilde{X})]_{+})\|_{2}$, where $\cdot\,$$_{+}$ is a component-wise operation and maps each entry $e$ to $\max\{e,0\}$.
Under the above assumption, we can bridge the tangent sets to $\mathcal{M}$ with the directional derivatives of $c_{1}$ and $c_{2}$.
Theorem 3.1(Computation of tangent sets)
Suppose that $\mathcal{M}$ satisfies Assumption 1 at $X\in\mathcal{M}$, and define the index set $I_{0}(X):=\{j\in\{1,\dots,n_{2}\}\mid c_{2}(X)_{j}=0\}$.
  1. (i)
    (First-order) If $c_{1}$ and $c_{2}$ are directionally differentiable at $X$, then
    $\mathrm{T}_{\mathcal{M}}(X)=\big\{\eta\in\mathbb{R}^{q}\,|\ c_{1}^{\prime}(X;\eta)=0,\ c_{2}^{\prime}(X;\eta)_{j}\leq 0\ \ \text{for all}\ j\in I_{0}(X)\big\}.$
  2. (ii)
    (Second-order) If, in addition, $c_{1}$ and $c_{2}$ admit parabolic second-order directional derivatives at $X$ for every direction pair $(\eta,\zeta)$, then for any $\eta\in\mathrm{T}_{\mathcal{M}}(X)$,
    $\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)=\Big\{\zeta\in\mathbb{R}^{q}\,|\ c_{1}^{\prime\prime}(X;\eta,\zeta)=0,\ c_{2}^{\prime\prime}(X;\eta,\zeta)_{j}\leq 0\ \ \text{for all}\ j\in I_{1}(X;\eta)\Big\},$
    where $I_{1}(X;\eta):=\{j\in I_{0}(X)\,|\ c_{2}^{\prime}(X;\eta)_{j}=0\}$.
Proof
(i) If $\eta\in\mathrm{T}_{\mathcal{M}}(X)$, there exist $t_{i}\to 0,$ and $\eta_{i}\to\eta$ such that $X+t_{i}\eta_{i}\in\mathcal{M}$. Therefore, we have $0=c_{1}(X+t_{i}\eta_{i})-c_{1}(X)=t_{i}c_{1}^{\prime}(X;\eta)+o(t_{i})$ and $0\geq c_{2}(X+t_{i}\eta_{i})_{j}-c_{2}(X)_{j}=t_{i}c_{2}^{\prime}(X;\eta)_{j}+o(t_{i})$ for $j\in I_{0}(X)$, by the local Lipschitzness of $c_{1},c_{2}$. Dividing the (in)equalities by $t_{i}$ and letting $t_{i}\to 0$ yield $c_{1}^{\prime}(X;\eta)=0$ and $c_{2}^{\prime}(X;\eta)_{j}\leq 0$, respectively.
Conversely, given a direction $\eta$ with $c_{1}^{\prime}(X;\eta)=0$, which implies that $c_{1}(X+t\eta)-c_{1}(X)=o(t)$ and thus $c_{1}(X+t\eta)=o(t)$. For $j\in I_{0}(X)$, if $c_{2}^{\prime}(X;\eta)_{j}<0$, it holds that $c_{2}(X+t\eta)_{j}<0$ when $t$ is small enough. If $c_{2}^{\prime}(X;\eta)_{j}=0$, we have $c_{2}(X+t\eta)_{j}=c_{2}(X+t\eta)_{j}-c_{2}(X)_{j}=o(t)$. Consequently, we have $\operatorname{dist}(X+t\eta,\mathcal{M})=o(t)$ since $\operatorname{dist}(X+t\eta,\mathcal{M})\leq\rho\|(c_{1}(X+t\eta),\max\{c_{2}(X+t\eta),0\})\|$. Therefore, it is concluded that $\eta\in\mathrm{T}_{\mathcal{M}}(X)$ by definition of the Bouligand tangent cone.
(ii) If $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$, there exist $t_{i}\to 0$ and $\zeta_{i}\to\zeta$ such that $X+t_{i}\eta+\frac{1}{2}t_{i}^{2}\zeta_{i}\in\mathcal{M}$, which reveals that $0=c_{1}(X+t_{i}\eta+\frac{1}{2}t_{i}^{2}\zeta_{i})-c_{1}(X)=\frac{1}{2}t_{i}^{2}c_{1}^{\prime\prime}(X;\eta,\zeta)+o(t_{i}^{2})$, and $0\geq c_{2}(X+t_{i}\eta+\frac{1}{2}t_{i}^{2}\zeta_{i})_{j}-c_{2}(X)_{j}=\frac{1}{2}t_{i}^{2}c_{2}^{\prime\prime}(X;\eta,\zeta)_{j}+o(t_{i}^{2})$ for $j\in I_{1}(X;\eta)$, by the local Lipschitzness of $c_{1},c_{2}$. Hence, we divide the (in)equalities by $t_{i}^{2}$ and let $t_{i}\to 0$ to obtain $c_{1}^{\prime\prime}(X;\eta,\zeta)=0$ and $c_{2}^{\prime\prime}(X;\eta,\zeta)_{j}\leq 0$.
Conversely, if $c_{1}^{\prime\prime}(X;\eta,\zeta)=0$, it holds that $c_{1}(X+t\eta+\frac{1}{2}t^{2}\zeta)-c_{1}(X)=o(t^{2})$. Additionally, for $j\in I_{1}(X;\eta)$, if $c_{2}^{\prime\prime}(X;\eta,\zeta)_{j}<0$, it holds that $c_{2}(X+t\eta+\frac{1}{2}t^{2}\zeta)_{j}<0$ for small enough $t$; if $c_{2}^{\prime\prime}(X;\eta,\zeta)_{j}=0$, it holds that $c_{2}(X+t\eta+\frac{1}{2}t^{2}\zeta)_{j}=o(t^{2})$. Therefore, we derive that $\operatorname{dist}(X+t\eta+\frac{1}{2}t^{2}\zeta,\mathcal{M})=o(t^{2})$ since $\operatorname{dist}(X+t\eta+\frac{1}{2}t^{2}\zeta,\mathcal{M})\leq\rho\|(c_{1}(X+t\eta+\frac{1}{2}t^{2}\zeta),\max\{o(t^{2}),0\})\|$. By definition of the second-order tangent set, we have $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$.
Remark 1
In fact, the inequality $\operatorname{dist}(\tilde{X},\mathcal{M})\leq\rho\|(c_{1}(\tilde{X}),\,[c_{2}(\tilde{X})]_{+})\|_{2}$ in Assumption 1 is the so-called error bound property luo1993errorbound; luo1994errorboundapplication, which has garnered wide interest in optimization and variational analysis; see luo1993errorbound; luo1994errorboundapplication; aze2003errorboundsurvey and references therein for more details.
3.2 Directional derivatives of singular values
We aim at applying Theorem 3.1 to the determinantal variety $\mathcal{M}_{\leq r}=\{X\in\mathbb{R}^{m\times n}\mid\sigma_{r+1}(X)=0\}$. To this end, we review the first- and second-order directional derivatives of singular-value mappings, as developed in lewis2005nonsmoothPartI; zhang2013secondordersingular; ding2014introductiontomatrixCP. These will be instrumental in identifying the zeros of $\sigma_{r+1}^{\prime}$ and $\sigma_{r+1}^{\prime\prime}$, which in turn characterize the tangent sets to $\mathcal{M}_{\leq r}$.
We follow the notation introduced in zhang2013secondordersingular, which, although somewhat involved, is essential for the analysis. Without loss of generality, assume that $m\leq n$. Given $X\in\mathbb{R}^{m\times n}$ with the full SVD,
$X=\bar{U}[\bar{\varSigma}\ 0]\bar{V}^{\top},$ (3.2)
where $\operatorname{diag}(\bar{\varSigma})=(\sigma_{1}(X),\sigma_{2}(X),\ldots,\sigma_{m}(X))$. The set of such matrices $(\bar{U},\bar{V})$ is defined by
$\mathcal{O}^{m,n}(X):=\{(\bar{U}^{\prime},\bar{V}^{\prime})\in\mathcal{O}(m)\times\mathcal{O}(n)\mid X=\bar{U}^{\prime}[\bar{\varSigma}\ 0]\bar{V}^{\prime\top}\}.$
Similarly, if $X\in\mathcal{S}(m)$, we define the following set according to the spectral decomposition,
$\mathcal{O}^{n}(X):=\{\bar{U}^{\prime}\in\mathcal{O}(n)\mid X=\bar{U}^{\prime}\bar{\varLambda}\bar{U}^{\prime\top}\},$
where $\bar{\varLambda}=\operatorname{Diag}(\lambda_{1}(X),\lambda_{2}(X),\ldots,\lambda_{m}(X))$ collects the eigenvalues of $X$. Let $I\subseteq\{1,2,\ldots,m\}$ and $J\subseteq\{1,2,\ldots,n\}$ be index sets. For a matrix $Z$, we denote by $Z_{IJ}$ the submatrix obtained by selecting the rows indexed by $I$ and the columns indexed by $J$. Likewise, $Z_{I}$ denotes the submatrix of $Z$ formed by the columns indexed by $I$.
The index set $\{1,2,\ldots,n\}$ is divided into three subsets: $\alpha=\{i\mid\sigma_{i}(X)>0,1\leq i\leq m\}$, $\beta=\{i\mid\sigma_{i}(X)=0,1\leq i\leq m\}$, and $\beta_{0}=\{m+1,\ldots,n\}$. Suppose that $X$ admits $t+1$ distinct singular values, $\mu_{1}>\mu_{2}>\ldots>\mu_{t}>\mu_{t+1}=0$, based on which we categorize the index set $\alpha$ into $\alpha_{k}=\left\{i\mid\sigma_{i}(X)=\mu_{k},1\leq i\leq m\right\}$ for $k=1,\ldots,t$. Denote $\alpha_{t+1}=\beta$ and $\widehat{\beta}=\beta\cup\beta_{0}$. We then partition the matrices as $\bar{U}=[\bar{U}_{\alpha_{1}}\ \bar{U}_{\alpha_{2}}\ \cdots\ \bar{U}_{\alpha_{t+1}}]$ and $\bar{V}=[\bar{V}_{\alpha_{1}}\ \bar{V}_{\alpha_{2}}\ \cdots\ \bar{V}_{\alpha_{t}}\ \bar{V}_{\widehat{\beta}}$], where $\bar{U}_{\alpha_{k}}\in\mathbb{R}^{m\times|\alpha_{k}|}$ for $k=1,2,\ldots,t+1$, $\bar{V}_{\alpha_{k}}\in\mathbb{R}^{n\times|\alpha_{k}|}$ for $k=1,2,\ldots,t$, and $\bar{V}_{\widehat{\beta}}\in\mathbb{R}^{n\times|\widehat{\beta}|}$.
Given a direction $\eta\in\mathbb{R}^{m\times n}$, we denote $\bar{\eta}=\bar{U}^{\top}\eta\bar{V}$. For $k=1,2,\ldots,t$, let the eigenvalues of $\frac{1}{2}(\bar{\eta}_{\alpha_{k}\alpha_{k}}+\bar{\eta}_{\alpha_{k}\alpha_{k}}^{\top})$ be $\lambda_{i}^{k}$ ($i=1,2,\ldots,|\alpha_{k}|$) and the distinct eigenvalues be $\theta_{1}^{k}>\theta_{2}^{k}>\ldots>\theta_{N_{k}}^{k}$, which induce the partitions $\beta_{j}^{k}:=\{i\mid\lambda_{i}^{k}=\theta_{j}^{k},i=1,\ldots,|\alpha_{k}|\}$ for $j=1,2,\ldots,N_{k}$. In parallel, letting the distinct singular values of $\bar{\eta}_{\beta\widehat{\beta}}$ be $\theta_{1}^{t+1}>\theta_{2}^{t+1}>\ldots>\theta_{N_{t+1}+1}^{t+1}=0$, we define $\beta_{j}^{t+1}=\{i\mid\sigma_{i}(\bar{\eta}_{\beta\widehat{\beta}})=\theta_{j}^{t+1},i=1,\ldots,|\beta|\}$ for $j=1,2,\ldots,N_{t+1}+1$. Fig. 4 illustrates the partitions of the index sets. Moreover, we need the following mappings between index sets,
$q_{a}$ $:\{1,\ldots,m\}\rightarrow\{1,\ldots,t+1\},\ q_{a}(i)=k,\text{ if }i\in\alpha_{k},$ (3.3)
$l$ $:\{1,\ldots,m\}\rightarrow\mathbb{N},\ l(i)=i-\kappa_{q_{a}(i)-1},$
$q_{b}$ $:\{1,\ldots,m\}\rightarrow\mathbb{N},\ q_{b}(i)=e,\ \text{if}\ l(i)\in\beta_{e}^{q_{a}(i)},$
$l^{\prime}$ $:\{1,\ldots,m\}\rightarrow\mathbb{N},\ l^{\prime}(i)=l(i)-\kappa_{q_{b}(i)-1}^{(q_{a}(i))},$
where $\kappa_{i}:=\sum_{j=1}^{i}|\alpha_{j}|$ and $\kappa_{i}^{(k)}:=\sum_{j=1}^{i}|\beta_{j}^{k}|$.
[…]
Figure 4: Illustration of partitions for the index sets, where $s$ denotes the rank of $X$, $\alpha$ corresponds to the indices of nonzero singular values, $\beta$ corresponds to the zero singular values, and $\beta_{0}$ represents the indices of the remaining dimensions $m+1,\dots,n$.
Given an index $i\in\{1,2,\ldots,m\}$, the rules for computing the first- and second-order directional derivatives of $\sigma_{i}$ vary depending on the location of $i$, determined by the values $q_{a}(i)$, $l(i)$, $q_{b}(i)$, and $l^{\prime}(i)$. We organized the results of (zhang2013secondordersingular, Theorem 3.1) as the flowchart in Fig. 5, where the expression of $\widehat{V}_{k}$ for $k\in\{1,2,\ldots,t+1\}$ is detailed as follows,
$\widehat{V}_{k}(\eta,\zeta)\!=\!\begin{cases}\frac{\bar{\zeta}_{\alpha_{k}\alpha_{k}}+\bar{\zeta}^{\top}_{\alpha_{k}\alpha_{k}}}{2}+\frac{1}{\mu_{k}}(\frac{\bar{\eta}_{\alpha_{k}\alpha_{k}}-\bar{\eta}^{\top}_{\alpha_{k}\alpha_{k}}}{2})^{\top}(\frac{\bar{\eta}_{\alpha_{k}\alpha_{k}}-\bar{\eta}^{\top}_{\alpha_{k}\alpha_{k}}}{2})\\ +\frac{1}{\mu_{k}}(\bar{\eta}^{\top}_{\beta\alpha_{k}}\bar{\eta}_{\beta\alpha_{k}}+\bar{\eta}_{\alpha_{k}\beta}\bar{\eta}^{\top}_{\alpha_{k}\beta}+\bar{\eta}_{\alpha_{k}\beta_{0}}\bar{\eta}^{\top}_{\alpha_{k}\beta_{0}})\\ +\!\!\sum\limits_{\begin{subarray}{c}j\neq k\\ j\leq t\end{subarray}}\frac{\mu_{j}\bar{\eta}_{\alpha_{k}\alpha_{j}}\bar{\eta}_{\alpha_{j}\alpha_{k}}+\mu_{k}\bar{\eta}^{\top}_{\alpha_{j}\alpha_{k}}\bar{\eta}_{\alpha_{j}\alpha_{k}}+\mu_{k}\bar{\eta}_{\alpha_{k}\alpha_{j}}\bar{\eta}^{\top}_{\alpha_{k}\alpha_{j}}+\mu_{j}\bar{\eta}^{\top}_{\alpha_{j}\alpha_{k}}\bar{\eta}^{\top}_{\alpha_{k}\alpha_{j}}}{2(\mu_{k}^{2}-\mu_{j}^{2})},\hskip 2.84526pt\text{if}\ k\leq t,\\[2.84526pt] \bar{\zeta}_{\beta\widehat{\beta}}-2\bar{\eta}_{\beta\alpha}\bar{\varSigma}^{-1}_{\alpha\alpha}\bar{\eta}_{\alpha\widehat{\beta}},\hskip 162.1807pt\text{if}\ k=t+1,\end{cases}$
[…]
Figure 5: Flowchart of computing directional derivatives of singular values in directions $\eta,\zeta\in\mathbb{R}^{m\times n}$, where we abbreviate $\sigma_{i}^{\prime}(X;\eta)$ and $\sigma_{i}^{\prime\prime}(X;\eta,\zeta)$ as $\sigma_{i}^{\prime}$ and $\sigma_{i}^{\prime\prime}$, respectively.
Generally, to compute the directional derivatives, it first determines whether the singular value associated with $i$ is nonzero: if $1\leq i\leq\operatorname{rank}(X)$, i.e., $i\in\alpha_{k}$ for some $k\leq t$, the first row in the flowchart applies. Otherwise, we have $i\in\beta$, and a second stage further determines how the second-order derivative is evaluated, depending on whether the index $q_{b}(i)$ corresponds to a zero singular value of $\bar{\eta}_{\beta\widehat{\beta}}$.
3.3 Tangent sets to $\mathcal{M}_{\leq r}$
Given $X\in\mathcal{M}_{\leq r}$, we begin by identifying the mapping $\sigma_{r+1}^{\prime}(X;\cdot)$, and then find its zeros to explicitly give the tangent cone to the determinantal variety, as an application of Theorem 3.1. Section 3.2 reveals that computing directional derivatives of $\sigma_{r+1}$ is generally nontrivial. Nevertheless, we then show that the condition $\sigma_{r+1}(X)=0$ and the geometry of $\mathcal{M}_{\leq r}$ can facilitate the derivation.
[…]
$\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ $=\{\eta\in\mathbb{R}^{m\times n}\mid\sigma_{r+1}^{\prime}(X;\eta)=0\}$ (3.5)
$=\{\eta\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(U_{\bot}^{\top}\eta V_{\bot})\leq r-s\}.$
This recovers the existing result (2.13), which follows by parameterizing $\eta$ by $\eta=UW_{1}V^{\top}+UW_{2}V_{\bot}^{\top}+U_{\bot}W_{3}V^{\top}+R$ with $W_{i}$ ($i=1,2,3$) being matrices of appropriate dimensions and $R\in\mathrm{N}_{\mathcal{M}_{s}}(X)$. Additionally, a recent work olikier2025fourtangentproof focuses on various descriptions of the tangent cone to $\mathcal{M}_{\leq r}$; we remark that the characterization (3.5) serves as an alternative perspective.
More importantly, Theorem 3.1 enlightens the following proposition, which provides a closed-form expression for the second-order tangent set to $\mathcal{M}_{\leq r}$—a new development on the geometry of the determinantal variety.
Proposition 1
Given $X\in\mathcal{M}_{\leq r}$ and $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$, where $\operatorname{rank}{(X)}=s$ and $\operatorname{rank}(\mathcal{P}_{\mathrm{N}_{\mathcal{M}_{s}}(X)}(\eta))=\ell-s$ for some $s\leq\ell\leq r$. Let the compact SVDs be $X=U\varSigma V^{\top}$ and $\mathcal{P}_{\mathrm{N}_{\mathcal{M}_{s}}(X)}(\eta)=U_{\eta}\varSigma_{\eta}V_{\eta}^{\top}$, respectively. Take ${U}_{\eta\bot}$ and ${V}_{\eta\bot}$ such that $U\ U_{\eta}\ U_{\eta\bot}]\in\mathcal{O}(m)$ and $V\ V_{\eta}\ V_{\eta\bot}]\in\mathcal{O}(n)$, and denote $U^{+}=[U\ U_{\eta}]$, $V^{+}=[V\ V_{\eta}]$. It holds that
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)=\left\{{2\eta X^{\dagger}\eta+[U^{+}\ U_{\eta\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right][V^{+}\ V_{\eta\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{\ell\times\ell},\\ W_{2}\in\mathbb{R}^{\ell\times(n-\ell)},\\ W_{3}\in\mathbb{R}^{(m-\ell)\times\ell},\\ J\in\mathbb{R}^{(m-\ell)\times(n-\ell)},\\ \operatorname{rank}(J)\leq r-\ell\end{array}\right.}\right\}$. (3.6)
Proof
According to Theorem 3.1 and expression (3.5), given the $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$, the second-order tangent set to $\mathcal{M}_{\leq r}$ at $X$ in direction $\eta$ coincides with the zeros of the mapping $\sigma_{r+1}^{\prime\prime}(X;\eta,\cdot)$.
Without loss of generality, we can assume that the full SVD of $X$ is $X=\bar{U}[\bar{\varSigma}\ 0]\bar{V}^{\top}$, and it holds that $U=[\bar{U}_{\alpha_{1}}\ \bar{U}_{\alpha_{2}}\ \cdots\ \bar{U}_{\alpha_{t}}]\in\mathrm{St}(m,s)$, $U_{\eta}\ U_{\eta\bot}]=\bar{U}_{\beta}\in\mathrm{St}(m,m-s)$, $V=[\bar{V}_{\alpha_{1}}\ \bar{V}_{\alpha_{2}}\ \cdots\ \bar{V}_{\alpha_{t}}]\in\mathrm{St}(n,s)$, and $V_{\eta}\ V_{\eta\bot}]=\bar{V}_{\widehat{\beta}}\in\mathrm{St}(n,n-s)$. In this case,
$\bar{\eta}_{\beta\widehat{\beta}}=\bar{U}_{\beta}^{\top}\eta\bar{V}_{\widehat{\beta}}=[U_{\eta}\ U_{\eta\bot}]^{\top}\eta[V_{\eta}\ V_{\eta\bot}]=\left[\begin{matrix}\varSigma_{\eta}&0_{(\ell-s)\times(n-\ell)}\\ 0_{(m-\ell)\times(\ell-s)}&0_{(m-\ell)\times(n-\ell)}\\ \end{matrix}\right].$ (3.7)
Following the derivation in the first-order case (3.5), we have $q_{a}(r+1)=t+1$ and l(r+1)=rs+1l(r+1)=r-s+1. Through the second row of the flowchart in Fig. 5, the next step is to determine the indices $q_{b}(r+1)$ and $l^{\prime}(r+1)$.
We notice that $\operatorname{rank}(\bar{\eta}_{\beta\widehat{\beta}})=\operatorname{rank}(\varSigma_{\eta})=\ell-s\leq r-s$, and thus l(r+1)=rs+1l(r+1)=r-s+1 indexes a zero singular value of $\bar{\eta}_{\beta\widehat{\beta}}$. Hence, the definition of the mapping $q_{b}$ in (3.3) reveals that $q_{b}(r+1)=N_{t+1}+1$. Furthermore, substituting the values of $q_{a}(r+1)$, $l(r+1)$, and $q_{b}(r+1)$, we obtain
$l^{\prime}(r+1)=l(r+1)-\kappa^{(q_{a}(r+1))}_{q_{b}(r+1)-1}=r-s+1-\kappa^{(t+1)}_{N_{t+1}}=r-\ell+1,$
where the last equality holds since $\kappa^{(t+1)}_{N_{t+1}}=\operatorname{rank}(\bar{\eta}_{\beta\widehat{\beta}})=\ell-s$ counts the number of nonzero singular values of $\bar{\eta}_{\beta\widehat{\beta}}$.
Consequently, the indices $r+1\in\beta$ and $q_{b}(r+1)=N_{t+1}+1$ point the computation of $\sigma_{r+1}^{\prime\prime}$ to the branch at the bottom of the flowchart (see Fig. 5):
$\sigma_{r+1}^{\prime\prime}(X;\eta,\zeta)=$ $\ \sigma_{l^{\prime}(r+1)}(Q_{\beta\beta_{q_{b}(r+1)}^{t+1}}^{\top}\widehat{V}_{t+1}(\eta,\zeta)[\widehat{Q}_{\widehat{\beta}\beta_{q_{b}(r+1)}^{t+1}}\widehat{Q}_{\widehat{\beta}\beta_{0}}])$
$=$ $\ \sigma_{r-\ell+1}(Q_{\beta\beta_{N_{t+1}+1}^{t+1}}^{\top}\widehat{V}_{t+1}(\eta,\zeta)[\widehat{Q}_{\widehat{\beta}\beta_{N_{t+1}+1}^{t+1}}\widehat{Q}_{\widehat{\beta}\beta_{0}}]).$ (3.8)
Then we calculate $\widehat{V}_{t+1}(\eta,\zeta)$ according to the expressions provided in section 3.2:
$\widehat{V}_{t+1}(\eta,\zeta)$ $=[U_{\eta}\ U_{\eta\bot}]^{\top}(\zeta-2\eta V\varSigma^{-1}U^{\top}\eta)[V_{\eta}\ V_{\eta\bot}]$
$=[U_{\eta}\ U_{\eta\bot}]^{\top}(\zeta-2\eta X^{\dagger}\eta)[V_{\eta}\ V_{\eta\bot}].$ (3.9)
Moreover, noticing from (3.4) and (3.7) that $Q_{\beta\beta}$ and $\widehat{Q}_{\widehat{\beta}\widehat{\beta}}$ can be chosen as $Q_{\beta\beta}=I_{m-s}$ and $Q_{\widehat{\beta}\widehat{\beta}}=I_{n-s}$, which reveals that
$Q_{\beta\beta_{N_{t+1}+1}^{t+1}}=[0_{(m-\ell)\times(\ell-s)}\ I_{m-\ell}]^{\top}\ \text{and}\ \ [\widehat{Q}_{\widehat{\beta}\beta_{N_{t+1}+1}^{t+1}}\widehat{Q}_{\widehat{\beta}\beta_{0}}]=[0_{(n-\ell)\times(\ell-s)}\ I_{n-\ell}]^{\top}.$
Substituting the above equalities and the expression (3.9) into (3.8) yields
$\sigma_{r+1}^{\prime\prime}(X;\eta,\zeta)=\sigma_{r-\ell+1}(U_{\eta\bot}^{\top}(\zeta-2\eta X^{\dagger}\eta)V_{\eta\bot}).$
Therefore, $\sigma_{r+1}^{\prime\prime}(X;\eta,\zeta)=0$ is equivalent to $\operatorname{rank}(U_{\eta\bot}^{\top}(\zeta-2\eta X^{\dagger}\eta)V_{\eta\bot})\leq r$-$\ell$. Recall that $U^{+}=[U\ U_{\eta}]$, $V^{+}=[V\ V_{\eta}]$, and consider the decomposition $\zeta-2\eta X^{\dagger}\eta=U^{+}W_{1}V^{+\top}+U^{+}W_{2}V_{\eta\bot}^{\top}+U_{\eta\bot}W_{3}V^{+\top}+U_{\eta\bot}JV_{\eta\bot}^{\top}$ with $W_{i}$ ($i=1,2,3$) and $J$ being matrices of appropriate dimensions. In this view, $\sigma_{r+1}^{\prime\prime}(X;\eta,\zeta)=0$ is equivalent to $\operatorname{rank}(J)\leq r$-$\ell$. Applying Theorem 3.1 concludes the expression (3.6).
Note that $\mathcal{M}_{r}$ is relatively open in $\mathcal{M}_{\leq r}$, i.e., for any $X\in\mathcal{M}_{r}$, there exists a neighborhood $\mathcal{B}$ around $X$ such that $\mathcal{M}_{\leq r}\cap\mathcal{B}\subseteq\mathcal{M}_{r}$ olikier2022continuity. Therefore, the second-order tangent set to $\mathcal{M}_{r}$ at $X$ coincides with that to $\mathcal{M}_{\leq r}$. Specifically, we inherit the notation from Proposition 1 and find that $\ell=\operatorname{rank}(X)=r$, implying that $U_{\eta}$ and $V_{\eta}$ vanish, and thus $U_{\eta\bot}$ and $V_{\eta\bot}$ become $U_{\bot}$ and $V_{\bot}$, respectively. Consequently, the expression (3.6) reduces to the tangent set to the fixed-rank manifold, i.e., $\mathrm{T}^{2}_{\mathcal{M}_{r}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)$, and more concisely,
$\mathrm{T}^{2}_{\mathcal{M}_{r}}(X;\eta)=\left\{{2\eta X^{\dagger}\eta+[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&0\\ \end{matrix}\right][V\ V_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{r\times r},\\ W_{2}\in\mathbb{R}^{r\times(n-r)},\\ W_{3}\in\mathbb{R}^{(m-r)\times r}\end{array}\right.}\right\},$ (3.10)
for any $\eta\in\mathrm{T}_{\mathcal{M}_{r}}(X)$. This explicit formula for $\mathrm{T}^{2}_{\mathcal{M}_{r}}(X;\eta)$ is also a new result to our knowledge, shedding light on the (second-order) geometry of $\mathcal{M}_{r}$.
Remark 2
Given a smooth manifold $\mathcal{X}$ defined as a level set of a mapping $h$, where $\mathrm{D}h$ has full rank in $\mathcal{X}$, the second-order tangent set to $\mathcal{X}$ can, in theory, be computed via $\nabla h$ and $\nabla^{2}h$ (rockafellar2009variationalanalysis, Proposition 13.13). However, for the case of $\mathcal{M}_{r}$, the commonly adopted choice of $X\mapsto h(X)$ involves a specific partition of $X$ and the inverse of a submatrix (see (guillemin1974differentialtopology, §1.4) or (boumal2023introduction, §7.5)). As a result, incorporating such an $h$ directly into the computation makes it difficult to explicitly derive (3.10), which relies on the SVD of the full matrix XX—underlining the contribution of the approach developed in this section.
4 Extensions to structured low-rank sets
In this section, we show that the developed analysis can be extended to more scenarios. Specifically, Theorem 3.1 is applied to low-rank tensor varieties (4.1), symmetric matrices (4.12), and positive semidefinite matrices (4.15). Moreover, following the spirit of Theorem 3.1, we establish in Theorem 4.1 the intersection rules for the intersection of two general sets, which are then applied to obtain tangent sets to an array of structured low-rank sets; all the results are summarized in Table 1 and appendices.
4.1 Tangent sets to tensor varieties
Low-rank tensor decompositions compactly represent multi-dimensional data, capturing essential structure with far less storage. The Tucker format tucker1964extension, the tensor train (TT) format oseledets2011TTSVD, and more generally, the hierarchical Tucker (HT) format grasedyck2010hierarchicalSVD are among the most typical formats. Each format induces a kind of tensor variety when the low-rank structure is imposed, and in particular, the Tucker and TT varieties are special cases of the HT variety (see Appendix A.2). Therefore, in this section, we calculate the tangent sets to the HT variety, and then obtain the results for the Tucker and TT varieties as immediate reductions.
We call $\mathbf{X}$ a tensor if it is a $d$-dimensional array in the space $\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$, and introduce in Appendix A.1 the preliminaries for HT varieties, including the dimension tree $T$ with nodes denoted by $t\in T$ and the associated dimensions $\{n_{t},n_{t_{-}}\}_{t\in T}$, mode-$t$ matricization $X_{(t)}^{\mathrm{ht}}$ and tensorization $\mathrm{ten}^{\mathrm{ht}}_{(t)}(\cdot)$, and the HT rank $\mathrm{rank}_{\mathrm{ht}}(\mathbf{X})$. Given an array $\mathbf{r}=(r_{t})_{t\in T}$ of positive integers indexed by nodes of $T$, we define $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ as the set of tensors with an HT rank at most $\mathbf{r}$:
$\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}=\{\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\mathrm{rank}_{\mathrm{ht}}(\mathbf{X})\leq\mathbf{r}\},$ (4.1)
where the $\leq$” is understood component-wise. Through the lens of matricization, $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ coincides with the intersection of tensorized matrix varieties along different modes, i.e.,
$\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}=\bigcap_{t\in T}{\mathrm{ten}_{(t)}^{\mathrm{ht}}\left({\mathbb{R}^{n_{t}\times n_{t_{-}}}_{\leq r_{t}}}\right)},$ (4.2)
where we adopt $\mathbb{R}^{n_{1}\times n_{2}}_{\leq r}:=\{X\in\mathbb{R}^{n_{1}\times n_{2}}\mid\operatorname{rank}(X)\leq r\}$ to explicitly reveal the shape of the matrices. Therefore, $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ is also a real algebraic variety; we refer to it as the HT variety.
Motivated by (4.2), we extend the perspective (3.1) to the tensor scenario:
$\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}=\left\{{\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\sigma_{r_{t}+1}(X^{\mathrm{ht}}_{(t)})=0\ \text{for all}\ t\in T}\right\},$ (4.3)
which enlightens the application of Theorem 3.1. To this end, it suffices to verify that the $\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}$ given through (4.3) satisfies Assumption 1.
[…]
Consequently, we apply Theorem 3.1 to the variety $\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}$, unveiling the intersection rules for the associated tangent sets.
Proposition 2
The tangent sets to the tensor varieties $\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}$ equal the intersection of tensorized tangent sets to unfolding matrices along different modes, i.e.,
$\mathrm{T}_{\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}}(\mathbf{X})$ $=\bigcap_{t\in T}{\mathrm{ten}_{(t)}^{\mathrm{ht}}\left({\mathrm{T}_{\mathcal{R}_{t}}(X^{\mathrm{ht}}_{(t)})}\right)},$ (4.4)
$\mathrm{T}^{2}_{\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}}(\mathbf{X};\bm{\eta})$ $=\bigcap_{t\in T}{\mathrm{ten}_{(t)}^{\mathrm{ht}}\left({\mathrm{T}^{2}_{\mathcal{R}_{t}}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)})}\right)}\ \ \text{for any}\ \bm{\eta}\in\mathrm{T}_{\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}}(\mathbf{X}).$ (4.5)
where we denote $\mathcal{R}_{t}:=\mathbb{R}_{\leq r_{t}}^{n_{t}\times n_{t_{-}}}$.
Proof
See Appendix A.3.
To our knowledge, it is the first time that the intersection rules (4.4) and (4.5) are identified for low-rank HT varieties. In addition, with appropriate dimension trees, the results realize the tangent sets to TT varieties and to Tucker varieties as immediate reductions (see Appendix A.2). Specifically, the first-order rule (4.4) for TT and Tucker varieties have been reported in (kutschan2018tangentTT, Corollary 2.9) and (gao2025lowranktucker, Corollary 1), respectively; while the second-order counterparts are new results.
4.2 Tangent sets to intersection of sets
A natural question is whether the preceding analysis can be extended to the intersection of the low-rank set, e.g. $\mathcal{M}_{\leq r}$, with another set $\mathcal{K}$ defined as a level set of a mapping hh—a setting attracting growing interest in recent years cason2013iterative; li2020jotaspectral; li2023normalboundedaffine; yang2025spacedecouple. To this end, following the spirit of section 3.1, we analyze the tangent sets to a general intersection $\mathcal{M}\cap\mathcal{K}$, and then apply the results to several specific instances in sections 4.2.1-4.2.3; results are summarized in Table 1.
Given an Euclidean space $\mathbb{R}^{q}$, consider two sets as follows:
$\mathcal{M}=\{\tilde{X}\in\mathbb{R}^{q}\mid c_{1}(\tilde{X})=0\}\ \ \text{and}\ \ \mathcal{K}=\{\tilde{X}\in\mathbb{R}^{q}\mid h(\tilde{X})=0\},$ (4.6)
where $c_{1}:\mathbb{R}^{q}\to\mathbb{R}^{n_{1}}$ is a possibly nonsmooth mapping, and $h:\mathbb{R}^{q}\to\mathbb{R}^{n_{2}}$ is a smooth mapping. Studying the geometry of $\mathcal{M}\cap\mathcal{H}$ is obstructed in two respects: 1) the set $\mathcal{M}$ can be nonsmooth, and thus the intersection rule (2.6), which relies on transversality, becomes invalid; 2) application of the developed Theorem 3.1 is not straightforward, as it is uncertain whether the intersection $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1(ii)—partly because the metric projection onto the coupled set remains unclear in general.
To circumvent the nonsmooth geometry, we employ the idea of smooth parameterization, which introduces a smooth manifold $\overline{\mathcal{M}}\subseteq\mathbb{R}^{\bar{q}}$ and a smooth mapping $\phi:\mathbb{R}^{\bar{q}}\to\mathbb{R}^{q}$ with $\phi(\overline{\mathcal{M}})=\mathcal{M}$, originally proposed as a remedy for nonsmooth optimization problems levin2023remedy; levin2025effect. Subsequently, we pull back $\mathcal{K}$ through $\phi$ to obtain
$\overline{\mathcal{K}}:=\{\tilde{x}\in\mathbb{R}^{\bar{q}}\mid h(\phi(\tilde{x}))=0\}=\phi^{-1}(\mathcal{K}).$
Then $\mathbb{R}^{\bar{q}}$ is viewed as an auxiliary space, and it is hoped that the benign properties of $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ in the lift space $\mathbb{R}^{\bar{q}}$ can shed light on the analysis for $\mathcal{M}\cap\mathcal{K}$ in the original space $\mathbb{R}^{q}$. Consequently, we identify some mild regularity conditions as follows, which are illustrated in Fig. 6.
Assumption 2
Given $\mathcal{M}$ and $\mathcal{K}$ as in (4.6). At $X\in\mathcal{M}\cap\mathcal{K}$, $\mathcal{M}$ and $\mathcal{K}$ satisfy Assumption 1, respectively. Moreover, it admits a smooth parameterization $(\overline{\mathcal{M}}\subseteq\mathbb{R}^{\bar{q}},\phi)$ of $\mathcal{M}$ satisfying the following conditions:
  • (i)
    the differential $\mathrm{D}(h\circ\phi)$ has constant rank in a neighborhood of $\overline{\mathcal{K}}$, which implies that $\overline{\mathcal{K}}$ is a smooth manifold;
  • (ii)
    the manifolds $\overline{\mathcal{M}}$ and $\overline{\mathcal{K}}$ intersect transversally in the ambient space $\mathbb{R}^{\bar{q}}$;
  • (iii)
    there exists an $x\in\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ such that $\phi(x)=X$ and the restriction $\phi|_{\overline{\mathcal{M}}}:\overline{\mathcal{M}}\to\mathcal{M}$ is open [^2] The mapping $\phi_{\overline{\mathcal{M}}}:\overline{\mathcal{M}}\to\mathcal{M}$ is said to be open at $x\in\overline{\mathcal{M}}$ if it maps neighborhoods of xx—sets in $\overline{\mathcal{M}}$ containing $x$ in their interior—to neighborhoods of $\phi(x)\in\mathcal{M}$ endowed with the subspace topology inherited from the ambient space. at $x$.
[…]
Figure 6: Illustration of Assumption 2 with the three regularity conditions.
Remark 3
Assumption 2(i) ensures the regularity of the intersection $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$, while Assumption 2(iii) helps transfer local structure from the auxiliary space to the original space. In fact, these two conditions can be readily verified in certain scenarios. For example, when considering the LR parameterization for $\mathcal{M}_{\leq r}$, that is,
$\overline{\mathcal{M}}=\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\ \ \ \text{and}\ \ \ \phi:(L,R)\mapsto LR^{\top},$ (4.7)
the manifold $\overline{\mathcal{M}}$ coincides with the ambient Euclidean space, and thus the transversal property naturally holds for $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ provided $\overline{\mathcal{K}}$ is a manifold. In addition, the work levin2025effect showed that the openness of $\phi$ is commonly satisfied by smooth parameterizations of low-rank sets, which aligns with the spirit of our paper.
We prove in the following theorem that if $\mathcal{M}$ and $\mathcal{K}$ satisfy Assumption 2, the intersection $\mathcal{M}\cap\mathcal{H}$ satisfies Assumption 1, thereby validating the application of the developed Theorem 3.1 to $\mathcal{M}\cap\mathcal{K}$.
Theorem 4.1(Intersection rule)
Suppose that $\mathcal{M}$ and $\mathcal{K}$ satisfy Assumption 2 at $X\in\mathcal{M}\cap\mathcal{K}$. We have the following intersection rules for the tangent sets to $\mathcal{M}\cap\mathcal{K}$,
  1. (i)
    (First-order) If $c_{1}$ is directionally differentiable at $X$, then
    $\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)=\mathrm{T}_{\mathcal{M}}(X)\cap\mathrm{T}_{\mathcal{K}}(X);$ (4.8)
  2. (ii)
    (Second-order) If, in addition, $c_{1}$ admits parabolic second-order directional derivatives at $X$ for every direction pair $(\eta,\zeta)$, then for any $\eta\in\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)$,
    $\mathrm{T}^{2}_{\mathcal{M}\cap\mathcal{K}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{K}}(X;\eta).$ (4.9)
Proof
See Appendix B.
Remark 4
If Assumption 2 is relaxed to require Assumption 1 only for $\mathcal{M}$ rather than for both $\mathcal{M}$ and $\mathcal{K}$, the explicit calculation $\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)=\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime}(X;\eta)=0,\,h^{\prime}(X;\eta)=0\}$ still holds. The role of Assumption 1 for $\mathcal{K}$ is to guarantee the decoupling principle $\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime}(X;\eta)=0,\,h^{\prime}(X;\eta)=0\}=\mathrm{T}_{\mathcal{M}}(X)\cap\mathrm{T}_{\mathcal{K}}(X)$; an analogous rationale applies to the second-order counterpart.
Theorem 4.1 serves as an extension of Theorem 3.1, incorporating an additional set $\mathcal{K}$ and decoupling the computation of the tangent sets to $\mathcal{M}\cap\mathcal{K}$ into the intersection of those to each component. In essence, Theorem 4.1 provides a unified perspective to unveil the first-order rule (4.8), which were previously verified on a case-by-case basis for certain choices of $(\mathcal{M},\mathcal{K})$ cason2013iterative; li2023normalboundedaffine; yang2025spacedecouple; peng2025normalizedTT. More importantly, the second-order intersection rule (4.9) yields new theoretical insights into the geometry of the intersection $\mathcal{M}\cap\mathcal{K}$.
Next, we demonstrate the broad applicability of the developed Theorems 3.1 and 4.1 through several instances where low-rank sets intersect with additional structured sets. For clarity, we divide the discussion into three groups: first, low-rank rectangular matrices in $\mathbb{R}^{m\times n}$; second, low-rank symmetric matrices; and third, low-rank positive semidefinite matrices; see Table 1 for a summary.
4.2.1 Low-rank rectangular matrices
We begin with the intersection of the determinantal variety $\mathcal{M}_{\leq r}$ and another structured set $\mathcal{H}\subseteq\mathbb{R}^{m\times n}$. Four typical choices of $\mathcal{H}$ considered in existing literature are the affine manifold li2023normalboundedaffine, the Frobenius sphere cason2013iterative; yang2025spacedecouple; peng2025normalizedTT, the oblique manifold yang2025spacedecouple, and the hyperbolic manifold jawanpuria2019lowrankhyperbolic:
$\mathcal{H}_{1}$ $=\mathrm{Aff}(m,n)=\{X\in\mathbb{R}^{m\times n}\,|\ \mathcal{A}(X)-b={0}\},$ (4.10)
$\mathcal{H}_{2}$ $=\mathrm{S_{F}}(m,n)=\{X\in\mathbb{R}^{m\times n}\,|\ \left\|{X}\right\|^{2}_{\mathrm{F}}-1=0\},$
$\mathcal{H}_{3}$ $=\mathrm{Ob}(m,n)=\{X\in\mathbb{R}^{m\times n}\,|\ \operatorname{diag}(XX^{\top})-{\bf 1}={0}\},$
$\mathcal{H}_{4}$ $=\mathbb{H}^{n}_{m-1}=\{X\in\mathbb{R}^{m\times n}\mid X_{i}\in\mathbb{H}_{m-1}\ \text{for}\ i=1,2,\ldots,n\},$
where ${\bf 1}\in\mathbb{R}^{m}$ denotes an all-ones vector, $\mathbb{H}_{m-1}:=\{x\in\mathbb{R}^{m}\mid-x_{1}y_{1}+\sum_{i=2}^{m}x_{i}y_{i}=-1,\ x_{1}>0\}$, and $X_{i}$ extracts the $i$-th column of $X$. Checking that all the $\mathcal{M}_{\leq r}\cap\mathcal{H}_{j}$ ($j=1,2,3,4$) satisfy Assumption 2 by choosing $(\overline{\mathcal{M}},\phi)$ as the LR parameterization (4.7), we then apply Theorem 4.1 to obtain the following intersection rules,
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}_{j}}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathcal{H}_{j}}(X),$
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}\cap\mathcal{H}_{j}}(X;\eta)$ $=\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{H}_{j}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}_{j}}(X).$
The above results enable us to derive the closed-form expressions for the tangent sets to $\mathcal{M}_{\leq r}\cap\mathcal{H}_{j}$; more details are given in Appendix C.
4.2.2 Low-rank symmetric matrices
Subsequently, we turn to the symmetric scenario when $m=n$ and $\mathcal{H}=\mathcal{S}(n)\cap\mathcal{U}$ (resp. $\mathcal{S}^{+}(n)\cap\mathcal{U}$) for some $\mathcal{U}\subseteq\mathcal{S}(n)$—this is a topic of independent interest pataki1998lowranksolution; li2020jotaspectral, and thus we rewrite the intersection $\mathcal{M}_{\leq r}\cap\mathcal{H}$ in the more specific form $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}$ (resp. $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$), where it is recalled that
$\mathcal{S}_{\leq r}(n)=\left\{{X\in\mathcal{S}(n)\mid\operatorname{rank}(X)\leq r}\right\}.$ (4.11)
Inspired by the perspective (3.1), we consider the eigenvalue mappings arranged in a non-increasing order, $\lambda_{1}(X)\geq\lambda_{2}(X)\geq\cdots\geq\lambda_{n}(X)$, and then draw on the following characterization of $\mathcal{S}_{\leq r}(n)$,
$\mathcal{S}_{\leq r}(n)=\bigcup_{j=1}^{r+1}\mathcal{S}_{j},\ \text{with}\ \mathcal{S}_{j}:=\{X\in\mathcal{S}(n)\mid\lambda_{j}(X)=0,\,\lambda_{j+n-r-1}(X)=0\}.$ (4.12)
We briefly explain the above decomposition. In fact, any $X\in\mathcal{S}(n)$ with $\operatorname{rank}(X)\leq r$ must have at least nrn-r consecutive eigenvalues $\lambda_{i}$ equal zero, and by the non-increasing ordering of the $\lambda_{i}$, we conclude that $X\in\mathcal{S}_{j}$ for some $1\leq j\leq r+1$.
Combining the rules (2.4) and (4.12), we can identify the tangent sets to $\mathcal{S}_{\leq r}(n)$ by taking the union of those to each $\mathcal{S}_{j}$, which shifts our focus to $\mathcal{S}_{j}$. In fact, it can be verified that for $1\leq j\leq r+1$, the set $\mathcal{S}_{j}$ satisfies Assumption 1. Subsequently, we apply Theorem 3.1 to $\mathcal{S}_{j}$, obtaining the associated tangent sets as the zeros of directional derivatives of eigenvalue mappings, which is achievable since explicit expressions for directional derivatives of $\lambda_{i}$ are given in torki2001secondtoeigen; zhang2013secondordersingular; the tangent sets to $\mathcal{S}_{j}$ are derived in Proposition 15 of Appendix D.1. Collecting the results produces tangent sets to $\mathcal{S}_{\leq r}(n)$, as presented in the following Proposition.
Proposition 3
Given $X\in\mathcal{S}_{\leq r}(n)$ with $\operatorname{rank}{(X)}=s$ and the spectral decomposition $X=U\varLambda U^{\top}$ with $U\in\mathrm{St}(n,s)$. The tangent cone to $\mathcal{S}_{\leq r}(n)$ can be characterized by
$\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X)=\left\{{[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&J\\ \end{matrix}\right][U\ U_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(s),\\ W_{2}\in\mathbb{R}^{s\times(n-s)},\\ J\in\mathcal{S}_{\leq r-s}(n-s)\end{array}\right.}\right\}$. (4.13)
Additionally, given a direction $\eta\in\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X)$ parameterized in the above manner with $\operatorname{rank}(J)=\ell-s$ for some $s\leq\ell\leq r$. Let the spectral decomposition of $U_{\bot}JU^{\top}_{\bot}$ be $U_{\bot}JU^{\top}_{\bot}=U_{\eta}\varSigma_{\eta}U_{\eta}^{\top}$ with $U_{\eta}\in\mathrm{St}(n,\ell-s)$. Take ${U}_{\eta\bot}$ such that $U\ U_{\eta}\ U_{\eta\bot}]\in\mathcal{O}(n)$, and denote $U^{+}=[U\ U_{\eta}]$. It holds that
$\mathrm{T}^{2}_{\mathcal{S}_{\leq r}(n)}(X;\eta)=\left\{{2\eta X^{\dagger}\eta+[U^{+}\ U_{\eta\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&L\\ \end{matrix}\right][U^{+}\ U_{\eta\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(\ell),\\ W_{2}\in\mathbb{R}^{\ell\times(n-\ell)},\\ L\in\mathcal{S}_{\leq r-\ell}(n-\ell)\end{array}\right.}\right\}$. (4.14)
Proof
See Appendix D.2.
Furthermore, imposing an additional constraint $\mathcal{U}$ on $\mathcal{S}_{\leq r}(n)$ has recently received increasing attention. A representative example is $\mathcal{U}=\{X\in\mathcal{S}(n)\mid\|X\|_{\mathrm{F}}^{2}-1=0\}$ cason2013iterative; li2020jotaspectral. Extending the spirit of the rule (2.4) and the decomposition (4.12), we have $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}=\bigcup_{j=1}^{r+1}(\mathcal{S}_{j}\cap\mathcal{U})$, and thus it suffices to compute the tangent sets to each $\mathcal{S}_{j}\cap\mathcal{U}$, followed by assembling them to obtain those to the union $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}$; the results are provided in Appendix D.2.
4.2.3 Low-rank positive semidefinite matrices
As shown by Pataki pataki1998lowranksolution, semidefinite programs (SDPs) often admit low-rank solutions, underlining the importance of the geometry of low-rank positive semidefinite matrices:
$\mathcal{S}^{+}_{\leq r}(n)=\left\{{X\in\mathcal{S}(n)\ |\ X\succeq 0,\,\operatorname{rank}(X)\leq r}\right\}.$ (4.15)
It is noteworthy to observe that $\mathcal{S}^{+}_{\leq r}(n)$ coincides with $S_{r+1}$ defined in (4.12), i.e.,
$\mathcal{S}^{+}_{\leq r}(n)=\mathcal{S}_{r+1}=\{X\in\mathcal{S}(n)\mid\lambda_{r+1}(X)=0,\lambda_{n}(X)=0\}.$ (4.16)
The equality holds from the non-increasing ordering of eigenvalues, $\lambda_{1}(X)\geq\cdots\geq\lambda_{r}(X)\geq 0=\lambda_{r+1}(X)=\cdots=\lambda_{n}(X)$. Since the tangent sets to the $\mathcal{S}_{j}$ in (4.12) have been derived in Proposition 15, we specify the computation in the following proposition by taking $j=r+1$.
Proposition 4
Given $X\in\mathcal{S}^{+}_{\leq r}(n)$ with $\operatorname{rank}{(X)}=s$ and the spectral decomposition $X=U\varLambda U^{\top}$, where $U\in\mathrm{St}(n,s)$. The tangent cone to $\mathcal{S}^{+}_{\leq r}(n)$ can be characterized by
$\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)}(X)=\left\{{[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&J\\ \end{matrix}\right][U\ U_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(s),\\ W_{2}\in\mathbb{R}^{s\times(n-s)},\\ J\in\mathcal{S}^{+}_{\leq r-s}(n-s)\end{array}\right.}\right\}$.
Additionally, given a direction $\eta\in\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)}(X)$ parameterized in the above manner with $\operatorname{rank}(J)=\ell-s$ for some $s\leq\ell\leq r$. Let the spectral decomposition of $U_{\bot}JU^{\top}_{\bot}$ be $U_{\bot}JU^{\top}_{\bot}=U_{\eta}\varSigma_{\eta}U_{\eta}^{\top}$ with $U_{\eta}\in\mathrm{St}(n,\ell-s)$. Take ${U}_{\eta\bot}$ such that $U\ U_{\eta}\ U_{\eta\bot}]\in\mathcal{O}(n)$, and denote $U^{+}=[U\ U_{\eta}]$. It holds that
$\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)}(X;\eta)=\left\{{2\eta X^{\dagger}\eta+[U^{+}\ U_{\eta\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&L\\ \end{matrix}\right][U^{+}\ U_{\eta\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(\ell),\\ W_{2}\in\mathbb{R}^{\ell\times(n-\ell)},\\ L\in\mathcal{S}^{+}_{\leq r-\ell}(n-\ell)\end{array}\right.}\right\}$.
Proof
Setting $j=r+1$ in Proposition 15, the associated quantities $s_{+}$ and $\ell_{+}$ reduce to $s_{+}=s$ and $\ell_{+}=\ell-s$, respectively, and thus the tangent sets of $\mathcal{S}_{j}$ in (D.10) and (D.11) reduce directly to those of $\mathcal{S}_{r+1}=\mathcal{S}^{+}_{\leq r}(n)$.
Subsequently, taking into account an additional structured set $\mathcal{U}\subseteq\mathcal{S}(n)$, the geometry of the coupled set $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ becomes more complicated. Typically, low-rank SDPs with linear equality constraints have found a broad range of applications, which motivates the study of $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ with $\mathcal{U}(n)=\{X\in\mathcal{S}(n)\mid\mathcal{A}(X)-b=0\}$ boumal2020deterministic; levin2025effect. In this case, the tangent cone to $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ was first derived in (levin2025effect, Corollary 4.12), and we note that applying the established Theorem 4.1 not only recovers the first-order result directly, but also identifies the second-order tangent set to $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$; see Appendix D.3 for details.
5 Tangent sets bridge optimization landscapes
Consider a general constrained optimization problem as follows,
$\min_{X\in\mathcal{M}}\ f(X),$
where the feasible region $\mathcal{M}$ may be nonsmooth and nonconvex, with a typical example being $\mathcal{M}=\mathcal{M}_{\leq r}$. To circumvent the irregularity of $\mathcal{M}$, a common approach is to employ the technique of smooth parameterization levin2023remedy; rebjock2024boundedrank; gao2024desingularizationtensor; levin2025effect; yang2025spacedecouple, which introduces a smooth manifold $\overline{\mathcal{M}}$ to (over)parameterize $\mathcal{M}$, thereby inducing a Riemannian optimization problem; see Fig. 2 for an illustration and see absil2008optimization; boumal2023introduction for more background of Riemannian optimization.
Then, the central question is: what is the relationship between the landscapes—or more precisely, the stationary points—of the reformulated problem (P-M) and the original problem (P)? To answer this, we unveil that the first- and second-order tangent sets to $\mathcal{M}$ and $\overline{\mathcal{M}}$ play a crucial role, indeed bridging the landscapes of the two optimization problems. Specifically, section 5.1 reviews existing results on smooth parameterization. More essentially, section 5.2 presents the key finding: we precisely identify a sufficient and necessary condition under which the second-order stationary points of (P-M) map to those of (P).
5.1 Smooth parameterization for optimization problems
We begin by revisiting the first- and second-order stationarity for a constrained optimization problem; see ruszczynski2006nonlinearopt. Specifically, consider the problem $\min_{X\in\mathcal{M}}f(X)$. We say a point $X^{*}\in\mathcal{M}$ is first-order stationary if $\mathcal{P}_{\mathrm{T}_{\mathcal{M}}(X^{*})}(-\nabla f(X^{*}))=0$, or equivalently, $\nabla f(X^{*})\in\hat{\mathrm{N}}_{\mathcal{M}}(X^{*})$; and $X^{*}$ is second-order stationary if, in addition, it satisfies
$\langle{\nabla f(X^{*}),\zeta}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle\geq 0,\quad\text{for all}\ \zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X^{*};\eta),$ (5.1)
for every $\eta\in\mathrm{T}_{\mathcal{M}}(X^{*})$ such that $\langle{\nabla f(X^{*}),\eta}\rangle=0$. The definitions are general, and apply analogously when the region $\mathcal{M}$ is replaced by other sets, such as $\overline{\mathcal{M}}$.
Implementing algorithms directly on $\mathcal{M}$ may suffer from the nonsmoothness. To address this, smooth parameterization is introduced levin2023remedy; rebjock2024boundedrank; levin2025effect; yang2025spacedecouple. Specifically, as demonstrated in Fig. 2, let $\mathcal{M}$ and $\overline{\mathcal{M}}$ denote a (possibly nonsmooth) set in $\mathcal{E}$ and a smooth manifold embedded in $\mathcal{W}$, respectively, and let $\phi:\mathcal{W}\to\mathcal{E}$ be a smooth mapping between the two Euclidean spaces such that $\phi(\overline{\mathcal{M}})=\mathcal{M}$. Through the parameterization $(\overline{\mathcal{M}},\phi)$, the original nonsmooth problem (P) can thus be cast as a smooth Riemannian optimization problem (P-M).
However, nonlinear parameterizations may distort the optimization landscape, underscoring the importance of studying the relationship between the stationary points of the two problems. More formally, we say that the parameterization $(\overline{\mathcal{M}},\phi)$ satisfies $p\Rightarrow q$}” ($p,q=1$ or $2$) at $Y$, if for any objective function $f$, $Y$ being a $p$-th-order stationary point for problem (P-M) implies that $\phi(Y)$ is a $q$-th-order stationary point for problem (P).
Levin et al. levin2025effect established a comprehensive analysis to answer when the implications $1\!\Rightarrow\!1$” or $2\!\Rightarrow\!1$” hold. Specifically, let $X=\phi(Y)$, and it is proved in (levin2025effect, Theorem 2.4) that $1\!\Rightarrow\!1$” holds at $Y$ if and only if the image $\operatorname{im}(\mathrm{D}\phi_{Y}(\mathrm{T}_{\overline{\mathcal{M}}}(Y)))=\mathrm{T}_{\mathcal{M}}(X)$—that is, the differential $\mathrm{D}\phi$ fully preserves the information contained in the Bouligand tangent cone. Moreover, (levin2025effect, Theorem 3.23) provides several conditions to guarantee the property $2\!\Rightarrow\!1$”. These results collectively reveal that the parameterization approach is effective in finding first-order stationary points on the nonsmooth $\mathcal{M}$.
However, exactly identifying the conditions under which $2\!\Rightarrow\!2$” holds remains an open problem, since second-order stationarity on the (possibly nonsmooth) set $\mathcal{M}$ is more involved than its first-order counterpart, as remarked in (levin2025effect, §6).
5.2 Equivalence between second-order stationary points
We provide in this section a sufficient and necessary condition to characterize when second-order stationary points of (P-M) map to those of (P), i.e., $2\!\Rightarrow\!2$” holds. We define the following mappings, which borrow the idea from levin2025effect,
$\mathrm{\bf L}_{Y}:\mathrm{T}_{\overline{\mathcal{M}}}(Y)\to\mathcal{E}:\ v\mapsto\mathrm{D}\phi_{Y}[v],\ \ \text{for}\ Y\in\overline{\mathcal{M}},$
$\mathrm{\bf Q}_{Y,v}:\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)\to\mathcal{E}:\ u_{v}\mapsto\mathrm{D}\phi_{Y}[u_{v}]+\mathrm{D}^{2}\phi_{Y}[v,v],\ \ \text{for}\ Y\in\overline{\mathcal{M}}\ \text{and}\ v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y).$
Let $X=\phi(Y)$. In fact, the mappings $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ convey the geometric information encoded in the tangent sets of $\overline{\mathcal{M}}$ to those of $\mathcal{M}$. To see this, we note that for the manifold $\overline{\mathcal{M}}$, any $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ admits a curve $\gamma(t)$ such that $\gamma(0)=Y,\gamma^{\prime}(0)=v$, and thus $\mathrm{\bf L}_{Y}(v)=(\phi\circ\gamma)^{\prime}(0)$, which implies $\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathrm{T}_{\mathcal{M}}(X)$. Moreover, it is revealed from (rockafellar2009variationalanalysis, Proposition 13.13) that given any $u_{v}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)$ associated with $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, there exists a curve $\beta(t)$ satisfying $\beta(0)=Y$, $\beta^{\prime}(0)=v$, and $\beta^{\prime\prime}(0)=u_{v}$. Hence we have $\mathrm{\bf Q}_{Y,v}(u_{v})=(\phi\circ\beta)^{\prime\prime}(0)$, indicating that $\operatorname{im}(\mathrm{\bf Q}_{Y,v})\subseteq\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))$.
It is concluded in (levin2025effect, Theorem 2.4) that $1\!\Rightarrow\!1$” fails at $Y$ only when $\mathrm{\bf L}_{Y}$ loses information, i.e., $\operatorname{im}(\mathrm{\bf L}_{Y})\subsetneq\mathrm{T}_{\mathcal{M}}(X)$, which suggests that a “comparison” between the second-order tangent sets to $\overline{\mathcal{M}}$ and to $\mathcal{M}$ through $\phi$ would facilitate characterizing the $2\!\Rightarrow\!2$” property. This, however, appears more intricate than the first-order counterpart, suffering from two pains: 1) a direction $\eta=\mathrm{\bf L}_{Y}(v)$ may admit multiple preimages $v^{\prime}$ under $\mathrm{\bf L}_{Y}$; 2) the asymptotic behavior of $v_{i}$ approaching such $v^{\prime}$ (in the sense of $\mathrm{\bf L}_{Y}(v_{i})\to\mathrm{\bf L}_{Y}(v^{\prime})=\eta$) also plays a role, which necessitates aggregating the images of $\mathrm{\bf Q}_{Y,v_{i}}$ to capture $\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$. Motivated by the two considerations, we formalize the idea in Theorem 5.1.
Before delving into the analysis, we present some basic computations. The derivatives of $\bar{f}=f\circ\phi$ can be computed as follows,
$\nabla\bar{f}(Y)=\mathrm{\bf L}^{*}_{Y}(\nabla f(X))\ \ \text{and}\ \ \nabla^{2}\bar{f}(Y)=\mathrm{\bf L}^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{\bf L}_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y),$ (5.2)
where $\mathrm{\bf L}^{*}_{Y}:\mathcal{E}\to\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ is the adjoint of $\mathrm{\bf L}_{Y}$, and the mapping $\phi_{\eta}:\mathcal{W}\to\mathbb{R}$ is given by $\phi_{\eta}(W)=\langle{\eta,\phi(W)}\rangle$ for any $(\eta,W)\in(\mathcal{E},\mathcal{W})$. Then, given a pair $(v,u_{v})$, we can specify the following computation,
$\ \langle{\nabla\bar{f}(Y),u_{v}}\rangle+\langle{v,\nabla^{2}\bar{f}(Y)[v]}\rangle$
$=$ $\ \langle{\mathrm{\bf L}_{Y}^{*}(\nabla f(X)),u_{v}}\rangle+\left\langle v,\left({\mathrm{\bf L}^{*}_{Y}\circ\nabla^{2}f(X)\circ\mathrm{\bf L}_{Y}+\nabla^{2}\phi_{\nabla f(X)}(Y)}\right)[v]\right\rangle$
$=$ $\ \langle{\nabla f(X),\mathrm{\bf L}_{Y}(u_{v})}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle+\langle{\nabla f(X),\mathrm{D}^{2}\phi_{Y}[v,v]}\rangle$
$=$ $\ \langle{\nabla f(X),\mathrm{\bf Q}_{Y,v}(u_{v})}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle,$ (5.3)
where the first equality comes by substituting the derivatives of $\bar{f}$ (5.2). Moreover, we say that a sequence $\{\zeta_{i}+Z\}_{i\in\mathbb{N}}$ of translates of a subspace $Z\subseteq\mathcal{E}$ converges (necessarily to a translate of $Z$) if there exists a sequence $\{z_{i}\}_{i\in\mathbb{N}}\subseteq Z$ such that $\{\zeta_{i}+z_{i}\}_{i\in\mathbb{N}}$ converges.
In the following theorem, we use $\overline{\mathrm{conv}}(\cdot)$ to denote the closed convex hull of a set, and propose a sufficient and necessary condition for the property $2\!\Rightarrow\!2$”. More specifically, when the condition is violated, the failure of $2\!\Rightarrow\!2$” can be witnessed by an explicitly constructed objective $f$.
Theorem 5.1
The parameterization $(\overline{\mathcal{M}},\phi)$ satisfies “ $2\!\Rightarrow\!2$” at $Y\in\overline{\mathcal{M}}$ if and only if $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ where $X=\phi(Y)$, and for all $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$,
$\mathrm{T}^{2}_{\mathcal{M}}(X;\mathrm{\bf L}_{Y}(v))\subseteq\overline{\mathrm{conv}}\left({\bigcup_{\{v_{i}\}_{i\in\mathbb{N}}:\,\mathrm{\bf L}_{Y}(v_{i})\to\mathrm{\bf L}_{Y}(v)}\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,v_{i}}(u_{v_{i}})+\operatorname{im}(\mathrm{\bf L}_{Y})}\right)}\right).$ (5.4)
If “ $2\!\Rightarrow\!2$” does not hold, there always exists a smooth function $f$ such that $Y$ is second-order stationary for (P-M) while $\phi(Y)$ is not second-order stationary for (P).
Proof
We first prove the “if ” part by assuming that $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ and the inclusion (5.4) holds for all $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Suppose that $Y$ is second-order stationary for problem (P-M). Then, the first-order condition $\nabla\bar{f}(Y)\in\mathrm{N}_{\overline{\mathcal{M}}}(Y)$ implies that
$0=\langle{v,\nabla\bar{f}(Y)}\rangle=\langle{v,\mathrm{\bf L}_{Y}^{*}(\nabla f(X))}\rangle=\langle{\mathrm{\bf L}_{Y}(v),\nabla f(X)}\rangle,\ \ \text{for all}\ v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y),$
which, together with $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ confirms the first-order stationarity of $X$.
Subsequently, we turn to the second-order condition. Given an arbitrary $\eta\in\mathrm{T}_{\mathcal{M}}(X)$ and an associated $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$. By $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$, pick $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\eta=\mathrm{\bf L}_{Y}(v)$. According to the assumed inclusion (5.4), for any $\varepsilon>0$, there exist $v_{i}^{(j)}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\mathrm{\bf L}_{Y}(v_{i}^{(j)})\to\eta$, $u_{i}^{(j)}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v_{i}^{(j)})$, and
$s^{(j)}\in\lim_{i\to\infty}\big(\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+\operatorname{im}(\mathrm{\bf L}_{Y})\big)\ \ \text{such that}\ \ \|\,\zeta-\sum_{j=1}^{N}\mu_{j}s^{(j)}\,\|<\varepsilon,$ (5.5)
where the coefficients $\{\mu_{j}\}_{j=1}^{N}$ satisfy $\mu_{j}\geq 0$ and $\sum_{j=1}^{N}\mu_{j}=1$. For each $j$, choose a sequence $\{z_{i}^{(j)}\}_{i\in\mathbb{N}}\subseteq\operatorname{im}(\mathrm{\bf L}_{Y})$ realizing $s^{(j)}$ with $s^{(j)}=\lim_{i\to\infty}\big(\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+z_{i}^{(j)}\big)$. Since $Y$ is second-order stationary for (P-M), the condition (5.1) indicates that for any $(v_{i}^{(j)},u_{i}^{(j)})$, it holds that $\langle{\nabla\bar{f}(Y),u_{i}^{(j)}}\rangle+\langle{v_{i}^{(j)},\nabla^{2}\bar{f}(Y)[v_{i}^{(j)}]}\rangle\geq 0$. Applying the equality (5.3), and substituting $z^{(j)}_{i}\in\operatorname{im}(\mathrm{\bf L}_{Y})$ together with the first-order condition $\nabla f(X)\bot\operatorname{im}(\mathrm{\bf L}_{Y})$ yield that
$\langle{\nabla f(X),\mathrm{\bf Q}_{Y,v_{i}^{(j)}}(u_{i}^{(j)})+z^{(j)}_{i}}\rangle+\left\langle\mathrm{\bf L}_{Y}(v_{i}^{(j)}),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v_{i}^{(j)})]\right\rangle\geq 0.$
Letting $i\to\infty$ in the inequality, we obtain $\langle{\nabla f(X),s^{(j)}}\rangle+\left\langle\eta,\nabla^{2}f(X)[\eta]\right\rangle\geq 0$. Taking the convex combination with the coefficients $\{\mu_{j}\}_{j=1}^{N}$ gives
$\Big\langle\nabla f(X),\,\sum_{j=1}^{N}\lambda_{j}s^{(j)}\Big\rangle+\langle\eta,\nabla^{2}f(X)[\eta]\rangle\geq 0.$
Finally, letting the parameter $\varepsilon$ in (5.5) tend to 0, and using the continuity of $s\mapsto\langle\nabla f(X),s\rangle$, we have $\langle\nabla f(X),\zeta\rangle\ +\ \langle\eta,\nabla^{2}f(X)[\eta]\rangle\ \geq\ 0$. By the arbitrariness of $\eta\in\mathrm{T}_{\mathcal{M}}(X)$ and $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$, we conclude the second-order stationarity of $X=\phi(Y)\in\mathcal{M}$ for problem (P).
We then turn to the “only if ” part. Firstly, suppose that $\operatorname{im}(\mathrm{\bf L}_{Y})\subsetneq\mathrm{T}_{\mathcal{M}}(X)$. Taking polars reverses the inclusion $(\mathrm{T}_{\mathcal{M}}(X))^{\circ}\subsetneq\operatorname{im}(\mathrm{\bf L}_{Y}))^{\circ}$, where we note that the strict inclusion still holds since $\operatorname{im}(\mathrm{\bf L}_{Y})$ is a linear space. Pick $w\in\mathrm{T}_{\mathcal{M}}(X)\setminus\operatorname{im}(\mathrm{\bf L}_{Y})$. Let $H=\mathcal{P}_{\operatorname{span}\{w\}}$ and $f(\tilde{X})=\tfrac{1}{2}\langle{\tilde{X}-X,H(\tilde{X}-X)}\rangle$ with $\mathcal{P}_{\operatorname{span}\{w\}}$ representing the projection matrices onto $\operatorname{span}\{w\}$. Then $\nabla f(X)=0$ and $\nabla^{2}f(X)=H$. By the computation (5.2), $\nabla\bar{f}(Y)=0$ and $\nabla^{2}\bar{f}(Y)=\mathrm{\bf L}_{Y}^{*}\circ H\circ\mathrm{\bf L}_{Y}\succeq 0$. Hence $Y$ is second-order stationary for (P-M). However, at $X$, $\langle\nabla f(X),w\rangle=0$ and $\langle{w,\nabla^{2}f(X)[w]}\rangle=-\|w\|^{2}<0$, violating the second-order condition for (P). Therefore, to guarantee the property $2\!\Rightarrow\!2$”, $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ must hold.
Denote by $\mathcal{C}$ the set on the right side of (5.4). Then, we suppose that the inclusion (5.4) fails: there exist a $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ (accordingly, $\eta:=\mathrm{\bf L}_{Y}(v)$) and a $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$ such that $\zeta\notin\mathcal{C}$. Since $\mathcal{C}$ is both closed and convex, we apply the separation theorem to obtain $w\in\mathcal{E}\setminus\{0\}$ such that
$\langle w,\zeta\rangle\ <\ 0\ \ \text{and}\ \ \langle w,s\rangle\ \geq\ 0\ \ \text{for all}\ s\in\mathcal{C}.$
Noticing that the subspace $\operatorname{im}(\mathrm{\bf L}_{Y})$ belongs to $\mathcal{C}$, we have $w\perp\operatorname{im}(\mathrm{\bf L}_{Y})$, and thus $\mathrm{\bf L}_{Y}^{*}(w)=0$. Let $H=\mathcal{P}_{\operatorname{im}(\mathrm{\bf L}_{Y})\cap\operatorname{span}\{\eta\}^{\perp}}$, and then, for all $v^{\prime}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, it holds that
$\langle w,\,\mathrm{\bf Q}_{Y,v^{\prime}}(u_{v^{\prime}})\rangle+\langle\mathrm{\bf L}_{Y}(v^{\prime}),H\circ\mathrm{\bf L}_{Y}(v^{\prime})\rangle\geq 0,\ \ \text{for all}\ u_{v^{\prime}}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v^{\prime}),$ (5.6)
since $\langle w,\cdot\rangle\geq 0$ on $\mathcal{C}$ and $\mathrm{\bf Q}_{Y,v^{\prime}}(u_{v^{\prime}})\in\mathcal{C}$ by definition of $\mathcal{C}$.
Let $f(\tilde{X})=\langle w,\tilde{X}-X\rangle+\tfrac{1}{2}\langle{\tilde{X}-X,H(\tilde{X}-X)}\rangle$, which reveals that $\nabla f(X)=w$ and $\nabla^{2}f(X)=H$. By the computations (5.2), (5.3), and (5.6), we have
$\nabla\bar{f}(Y)=\mathrm{\bf L}_{Y}^{*}(w)=0,\ \ \text{and}\ \ \langle{\nabla\bar{f}(Y),u_{v^{\prime}}}\rangle+\langle{v^{\prime},\nabla^{2}\bar{f}(Y)[v^{\prime}]}\rangle\geq\ 0$
for all $v^{\prime}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ and $u_{v^{\prime}}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v^{\prime})$. Therefore, $Y$ is second-order stationary for (P-M). On the other hand, at $X$, we have $\langle\nabla f(X),\eta\rangle=\langle w,\mathrm{\bf L}_{Y}(v)\rangle=\langle{\mathrm{\bf L}_{Y}^{*}(w),\eta}\rangle=0$, and for the associated $\zeta\in\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)$,
$\langle\nabla f(X),\zeta\rangle+\langle{\eta,\nabla^{2}f(X)[\eta]}\rangle=\underbrace{\langle w,\zeta\rangle}_{<\,0}+\underbrace{\langle{\eta,H(\eta)}\rangle}_{=0}\ <\ 0,$
which violates the second-order optimality condition on $\mathcal{M}$. This contradicts the property $2\!\Rightarrow\!2$”, thereby showing the necessity of the inclusion (5.4).
To broaden the applicability, we extend the result to composition of parameterizations, which is inspired by (levin2025effect, §3.3)
Proposition 5
Let $(\overline{\mathcal{M}},\phi)$ be a smooth parameterization of $\mathcal{M}$. Given another smooth manifold $\mathcal{Z}$, let $\varphi:\mathcal{Z}\to\overline{\mathcal{M}}$ be a smooth mapping such that $\psi:=\phi\circ\varphi$ is surjective. Then $(\mathcal{Z},\psi)$ is a smooth parameterization of $\mathcal{M}$. Moreover, for $Z\in\mathcal{Z}$ and $Y:=\varphi(Z)\in\overline{\mathcal{M}}$, the following properties hold.
  • (i)
    If $(\mathcal{Z},\psi)$ satisfies “ $2\!\Rightarrow\!2$” at $Z$, then $(\overline{\mathcal{M}},\phi)$ satisfies “ $2\!\Rightarrow\!2$” at $Y$.
  • (ii)
    If $\varphi$ is a submersion at $Z$ and $(\overline{\mathcal{M}},\phi)$ satisfies “ $2\!\Rightarrow\!2$” at $Y$, then $(\mathcal{Z},\psi)$ satisfies “ $2\!\Rightarrow\!2$” at $Z$.
Proof
(i) Given any objective function $f$, suppose that $Y=\varphi(Z)$ is a second-order stationary point on $\overline{\mathcal{M}}$ (with respect to $f\circ\phi$). We note that any curve $\gamma(t)$ on $\mathcal{Z}$ passing through $Z$ at $t=0$ satisfies that
$(f\circ\phi\circ\varphi\circ\gamma)^{\prime}(0)=0\ \ \ \text{and}\ \ \ (f\circ\phi\circ\varphi\circ\gamma)^{\prime\prime}(0)\geq 0,$ (5.7)
which holds by viewing $\varphi\circ\gamma(t)$ as a curve on $\overline{\mathcal{M}}$ and considering the second-order stationarity of $Y$. Therefore, by (5.7), the point $Z$ is indeed second-order stationary on $\mathcal{Z}$, which, together with the $2\!\Rightarrow\!2$”property at $Z$, reveals the second-order stationarity of $\phi(Y)=\psi(Z)$.
(ii) Suppose that $Z$ is a second-order stationary point on $\mathcal{Z}$ (with respect to $f\circ\phi\circ\varphi$). Given any curve $\beta(t)$ on $\overline{\mathcal{M}}$ with $\beta(0)=Y$, since $\varphi$ is a submersion at $Z$, we can apply the local section theorem (lee2012manifolds, Theorem 4.26) to obtain a (locally defined) mapping $l:\overline{\mathcal{M}}\to\mathcal{Z}$ such that $(\varphi\circ l\circ\beta)(t)=\beta(t)$. Letting $\gamma(t):=(l\circ\beta)(t)$, the second-order stationarity of $Z$ indicates (5.7). Substituting $\gamma=l\circ\beta$ and $\varphi\circ l\circ\beta=\beta$ leads to $(f\circ\phi\circ\beta)^{\prime}(0)=0\ \text{and}\ (f\circ\phi\circ\beta)^{\prime\prime}(0)\geq 0$, the arbitrariness of $\beta(t)$ yielding the second-order stationarity of $Y$. Hence, the property $2\!\Rightarrow\!2$” at $Y$ concludes that $\psi(Z)=\phi(Y)$ is a second-order stationary point at $\mathcal{M}$, validating the $2\!\Rightarrow\!2$” propety at $Z$.
In summary, Theorem 5.1 establishes that the parameterization $(\overline{\mathcal{M}},\phi)$ avoids introducing spurious second-order stationary points only when the mappings $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ preserve the tangent information, that is, $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ and condition (5.4) holds. Moreover, Proposition 5 offers an additional perspective for verifying the $2\!\Rightarrow\!2$” property through the lens of compositions. These results will later be applied to two well-known parameterizations of $\mathcal{M}_{\leq r}$, the LR parameterization and the desingularization khrulkov2018desingularization; rebjock2024boundedrank; levin2025effect, enabling us to exactly determine the points on $\mathcal{M}_{\leq r}$ where the $2\!\Rightarrow\!2$” property holds.
6 Second-order optimality on bounded-rank matrices
This section applies the framework developed in sections 3-5 to low-rank optimization problems. We begin with problem (1.2), where the feasible region is the matrix variety $\mathcal{M}_{\leq r}$ coupled with an additional structured constraint $\mathcal{H}$. By substituting the characterizations of the first- and second-order tangent sets, we derive the first- and second-order optimality conditions in section 6.1. Then, the focus is shifted to the scenario when $\mathcal{H}=\mathbb{R}^{m\times n}$, leading to the following formulation,
$\min_{X\in\mathbb{R}^{m\times n}}$ $f(X)$ (6.1)
$\mathrm{s.\,t.}$ $X\in\mathcal{M}_{\leq r},$
which is of independent interest schneider2015Lojaconvergence; levin2023remedy. It is shown in section 6.2 that checking second-order optimality over $\mathcal{M}_{\leq r}$ is NP-hard in general. Nevertheless, in section 6.3, we clarify that second-order stationarity remains attainable in certain special cases, through the lens of smooth parameterizations.
6.1 Second-order optimality conditions
Regarding the feasible region $\mathcal{M}_{\leq r}\cap\mathcal{H}$ of problem (1.2), we recall the intersection rules developed in section 4.2.1,
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathcal{H}}(X),$ (6.2)
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X;\eta)$ $=\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{H}}(X;\eta),\ \ \text{for all}\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X),$
where $\mathcal{H}$ accommodates the ambient space $\mathbb{R}^{m\times n}$, the affine manifold li2023normalboundedaffine, the Frobenius sphere cason2013iterative; yang2025spacedecouple, the oblique manifold yang2025spacedecouple, and the hyperbolic manifold jawanpuria2019lowrankhyperbolic, as outlined in Table 1 and (4.10). Then the first- and second-order optimality conditions are derived in the following proposition for $\mathcal{H}$ in (4.10).
Proposition 6
Given $X^{*}\in\mathcal{M}_{\leq r}\cap\mathcal{H}$, it is a first-order stationary point for problem (1.2) if for all $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X^{*})\cap\mathrm{T}_{\mathcal{H}}(X^{*})$, it holds that $\langle{\nabla f(X^{*}),\eta}\rangle=0$; and $X^{*}$ is second-order stationary if it additionally satisfies that for all $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X^{*})\cap\mathrm{T}_{\mathcal{H}}(X^{*})$,
$\langle{\nabla f(X^{*}),\zeta}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle\geq 0,\quad\text{for all}\ \zeta\in\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X^{*};\eta)\cap\mathrm{T}^{2}_{\mathcal{H}}(X^{*};\eta).$
Proof
Notice that each $\mathcal{H}$ in (4.10) is a manifold, which implies that $\mathrm{T}_{\mathcal{H}}X$ is always a linear space. Combining this observation with the expression (2.13) and the rule (6.2) reveals that a direction $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X^{*})$ if and only if all $\rho\in\mathbb{R}$ yield $\rho\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X^{*})$. Hence we have $\nabla f(X^{*})\in\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X^{*})$ if and only if $\langle{\nabla f(X^{*}),\eta}\rangle=0$ for all $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X^{*})$, which, together with (6.2), can be substituted into (5.1) to produce the desired conclusion.
Taking $\mathcal{H}=\mathbb{R}^{m\times n}$ in (1.2), the problem reduces to (6.1)—in this case, specifying the computation in Proposition 6 is of independent interest.
Proposition 7
Given $X^{*}\in\mathcal{M}_{\leq r}$ with $\operatorname{rank}(X^{*})=s$, it is a second-order stationary point for (6.1) if it satisfies
[…]
where $\nabla_{\mathcal{M}_{r}}$ and $\nabla^{2}_{\mathcal{M}_{r}}$ denote the Riemannian gradient and Riemannian Hessian on $\mathcal{M}_{r}$, respectively.
Proof
The computation of Riemannian derivatives on $\mathcal{M}_{r}$ can be found in (boumal2023introduction, §7.5). Let the SVD of $X^{*}$ be $X^{*}=U\varSigma V^{\top}$. When $s=r$, the condition $\nabla f(X^{*})\in\mathrm{N}_{\mathcal{M}_{\leq r}}(X^{*})$ is equivalent, by (2.11), to $\nabla f(X^{*})-P_{U_{\bot}}\nabla f(X^{*})P_{V_{\bot}}=0$, which, in turn, is equivalent to the Riemannian first-order optimality condition $\nabla_{\mathcal{M}_{r}}f(X^{*})=\mathcal{P}_{\mathrm{T}_{\mathcal{M}_{r}}X^{*}}(\nabla f(X^{*}))=0$. For the second-order condition, substituting the explicit form of $\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X^{*};\eta)$ in (3.10) into (5.1) yields
$0\leq$ $\ \langle{\nabla f(X^{*}),\zeta}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle$
$=$ $\ \langle{\nabla f(X^{*}),2\eta X^{*\dagger}\eta+[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&0\\ \end{matrix}\right][V\ V_{\bot}]^{\top}}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle$
$=$ $\ \langle{\nabla f(X^{*}),2\eta X^{*\dagger}\eta}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle,$
where the last equality holds from $\nabla f(X^{*})=P_{U_{\bot}}\nabla f(X^{*})P_{V_{\bot}}$. The condition $\langle{\nabla f(X^{*}),2\eta X^{*\dagger}\eta}\rangle+\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle\geq 0$ for all $\eta\in\mathrm{T}_{\mathcal{M}_{r}}(X^{*})$ is equivalent to $\nabla^{2}_{\mathcal{M}_{r}}f(X^{*})\succeq 0$, as stated in (boumal2023introduction, §7.5).
When $s<r$, considering the expression (2.14), the first-order optimality requires that $\nabla f(X^{*})\in\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}}(X^{*})=\{0\}$. Taking $\nabla f(X^{*})=0$ in (5.1) reveals that $\langle{\eta,\nabla^{2}f(X^{*})[\eta]}\rangle\geq 0\ \text{for all}\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X^{*})$.
In fact, for the optimization problem over the determinantal variety (6.1), Proposition 7 shows that the second-order optimality condition coincides with the Riemannian one at points of rank $r$, whereas, at rank-deficient points, i.e., $\operatorname{rank}(X^{*})<r$, the second-order condition requires that the Euclidean Hessian admit no negative curvature directions along the tangent cone.
6.2 NP-hardness of verifying second-order optimality
Building on the optimality condition characterized in (6.3), one might expect that identifying a second-order stationary point would be a tractable task. However, we find that this problem is NP-hard in general. Worse still, unless P=NP, there does not exist a fully polynomial-time algorithm for deciding whether a given point is second-order stationary in an approximate sense. This section aims to establish the hardness of verifying second-order optimality for (6.1) through a reduction from a combinatorial problem that is known to be NP-complete.
Given a point $X\in\mathcal{M}_{\leq r}$ with $\operatorname{rank}(X)=s<r$ and the SVD $X=U\varSigma V^{\top}$, checking the first-order optimality can be accomplished in polynomial time cason2013iterative; schneider2015Lojaconvergence. Hence, according to (6.3), the main bottleneck lies in deciding the existence of a negative curvature direction $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$, which can be formulated as the following problem
$\min_{\eta\in\mathbb{R}^{m\times n}}$ $\langle{\eta,\mathcal{A}(\eta)}\rangle$ (6.4)
$\mathrm{s.\,t.}$ $\|\eta\|_{\mathrm{F}}=1,$
$\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X).$
Note that we abstract the Hessian as a symmetric operator $\mathcal{A}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}$. Let $\lambda^{\star}$ denote the optimal value, and then the task of verifying second-order optimality conditions, which is abbreviated as VERSOC, is equivalent to checking whether $\lambda^{\star}<0$.
Problem: VERSOC Input: Parameters $m,n,r$; point $X\in\mathcal{M}_{\leq r}$; symmetric operator $\mathcal{A}$. Question: Does the optimal value of (6.4) $\lambda^{\star}<0$?
Next, we introduce a combinatorial problem. Consider an undirected graph $G=(\mathcal{V},E)$ where $\mathcal{V}=\{1,2,\ldots,n\}$ is the vertex set and $E$ is the edge set with $(i,j)\in E$ indicating the existence of an edge between vertices $i$ and $j$. Additionally, a subset $S\subseteq\mathcal{V}$ is called a clique if every pair of vertices in $S$ is connected by an edge in $E$, and accordingly, the clique number of the graph is defined by $\omega(G):=\max\{|S|\mid S\subseteq\mathcal{V}\ \text{is a clique}\}$. The clique decision problem, abbreviated as CLIQUE, asks whether there exists a clique of size $K$, or equivalently, whether $\omega(G)\geq K$.
Problem: CLIQUE Input: Undirected graph $G=(\mathcal{V},E)$; clique size $K$. Question: Does there exist a clique of size $K$ in $G$?
In fact, CLIQUE is one of the original $21$ NP-complete problems karp1972reducibility. More importantly, Motzkin and Straus motzkin1965maxima bridged the quantity $\omega(G)$ with an optimization problem via the following formula,
$1-\frac{1}{\omega(G)}=\max_{x\in\mathrm{\Delta}_{n}}\sum_{(i,j)\in E}x_{i}x_{j},$ (6.5)
where $\mathrm{\Delta}_{n}:=\{x\in\mathbb{R}^{n}:0\leq x_{i}\leq 1,i=1,2,\ldots,n,\,x_{1}+x_{2}+\cdots+x_{n}=1\}$ is the unit simplex. Let $e_{1},e_{2},\ldots,e_{n}$ be the standard basis in $\mathbb{R}^{n}$.
Now, we are ready to show the reduction from CLIQUE to VERSOC, with the main idea outlined as follows—given an input $(G,K)$ for CLIQUE, we can construct an input $(m,n,r,X,\mathcal{A})$ for VERSOC with $m=n$, $r=1$, $X=0$ and the symmetric operator $\mathcal{A}$ satisfying
$\mathcal{A}:\mathbb{R}^{m\times n}\to\mathbb{R}^{m\times n}:\,\eta\mapsto(1-\frac{1}{K})\eta-\frac{1}{4}\sum_{(i,j)\in E}(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top})\eta(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top}).$
In this manner, $\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ coincides with $\{\eta\in\mathbb{R}^{m\times n}\mid\operatorname{rank}(\eta)\leq 1\}$, and thus (6.4) can be specified as follows,
$\min_{u,v\in\mathbb{R}^{n}}$ $1-\frac{1}{K-1}-\sum_{(i,j)\in E}\langle{uv^{\top},A_{ij}uv^{\top}A_{ij}}\rangle$ (6.6)
$\mathrm{s.\,t.}$ $\|u\|_{\mathrm{F}}=\|v\|_{\mathrm{F}}=1.$
where we denote $A_{ij}=\frac{1}{2}(e_{i}e_{j}^{\top}+e_{j}e_{i}^{\top})$ for $(i,j)\in E$. Then it can be shown that $G$ admits a $K$-clique if and only if the optimal value of (6.6) $\lambda^{\star}<0$.
Theorem 6.1(NP-hardness)
The problem CLIQUE is polynomially reducible to VERSOC, and thus verifying second-order optimality is NP-hard.
Proof
Notice that the matrices $A_{ij}$ are symmetric matrices. Therefore, we can derive
$\lambda^{\star}=$ $\min_{\|u\|_{\mathrm{F}}=\|v\|_{\mathrm{F}}=1}\ 1-\frac{1}{K-1}-\sum_{(i,j)\in E}\langle{uv^{\top},A_{ij}uv^{\top}A_{ij}}\rangle$
$=$ $\min_{\|u\|_{\mathrm{F}}=1}\ 1-\frac{1}{K-1}-\sum_{(i,j)\in E}(u^{\top}A_{ij}u)^{2}$ (6.7)
$=$ $\min_{\|u\|_{\mathrm{F}}=1}\ 1-\frac{1}{K-1}-\sum_{(i,j)\in E}u_{i}^{2}u_{j}^{2}$
$=$ $\ \frac{1}{\omega(G)}-\frac{1}{K-1},$ (6.8)
where (6.7) is obtained from (he2010approximationhomogeneous, Proposition 2) and (6.8) comes by considering $\left\|{u}\right\|_{\mathrm{F}}=1$ and parameterizing $x_{i}=u_{i}^{2}$ in (6.5).
If the graph $G$ admits a clique $S$ of size $K$, i.e., $\omega(G)>K-1$. Then, the derivation (6.8) indicates that $\lambda^{\star}=\frac{1}{\omega(G)}-\frac{1}{K-1}<0$. Conversely, if there does not exist a clique of size $K$ in $G$, i.e., $\omega(G)\leq K-1$, we have $\lambda^{\star}\geq 0$. Therefore, the result for VERSOC answers CLIQUE, which implies that VERSOC is NP-hard.
The above theorem concludes the NP-hardness of verifying second-order optimality for the problem (1.2). Furthermore, as we shall show, even finding an approximate solution for (6.4)—which one might hope could alleviate the hardness—still does not admit a polynomial-time strategy, unless P=NP.
Specifically, we call an algorithm a fully polynomial-time approximation scheme (or FPTAS), if given any $\varepsilon\in(0,1)$, it can return an $\varepsilon$-approximate solution $\tilde{\lambda}$ for (6.4) in the sense that $\tilde{\lambda}-\lambda^{\star}\leq\varepsilon$, and the complexity is upper bounded by a polynomial function of the problem size and $\varepsilon^{-1}$.
Theorem 6.2(No FPTAS)
Unless P=NP, there is no FPTAS for verifying whether a point is second-order stationary for (1.2).
Proof
Suppose that there exists a polynomial-time scheme, and we take $\varepsilon=\frac{1}{2K(K-1)}$. Therefore, if the graph admits a $K$-size clique, the FPTAS will yield $\tilde{\lambda}\leq\lambda^{\star}+\varepsilon=\frac{1}{\omega(G)}-\frac{1}{2}(\frac{1}{K-1}+\frac{1}{K})<0$. Conversely, if there is no $K$-size clique, we have $\tilde{\lambda}\geq\lambda^{\star}\geq 0$. In conclusion, we can answer CLIQUE if there were an FPTAS for VERSOC, which is possible only when P=NP.
6.3 Two parameterizations for matrix varieties
Although section 6.2 establishes the NP-hardness of identifying second-order stationary points in general, it remains natural to ask whether finding such points is still possible in certain special cases. To this end, we resort to the technique of smooth parameterization levin2025effect, since the conditions to guarantee $2\!\Rightarrow\!2$” have been developed in section 5.2.
Specifically, two parameterizations for $\mathcal{M}_{\leq r}$ are considered: the LR parameterization mishra2014fixedLR,
$\mathcal{M}_{\mathrm{LR}}=\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r},\ \ \ \ \ \phi_{\mathrm{LR}}(L,R)=LR^{\top},$ (LR)
and the desingularization khrulkov2018desingularization; rebjock2024boundedrank,
$\mathcal{M}_{\mathrm{desing}}=\{(X,G)\in\mathbb{R}^{m\times n}\times\mathrm{Gr}(n,n-r)\mid XG=0\},\ \ \phi_{\mathrm{desing}}(X,G)=X,$ (Desing)
where the Grassmann manifold bendokat2024grassmann is viewed as an embedded submanifold in $\mathcal{S}(n)$, i.e., $\mathrm{Gr}(n,s):=\{G\in\mathcal{S}(n)\mid G^{2}=G,\,\operatorname{rank}(G)=s\}$. More background on the two parameterizations can be found in khrulkov2018desingularization; rebjock2024boundedrank; levin2025effect; yang2025spacedecouple.
We now apply Theorem 5.1 to the two parameterizations in turn. The main principle is to compute explicitly the images of $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ defined in section 5.2, and substitute them into the conditions identified in Theorem 5.1, thereby determining when the parameterization produces second-order stationary points on $\mathcal{M}_{\leq r}$.
Proposition 8
The LR parameterization of $\mathcal{M}_{\leq r}$ given by (LR) satisfies “ $2\!\Rightarrow\!2$” at $(L,R)$ if and only if $\operatorname{rank}(LR^{\top})=r$.
Proof
If $\operatorname{rank}(LR^{\top})<r$, the expression (2.13) reveals that $\mathrm{T}_{\mathcal{M}_{\leq r}}(LR^{\top})$ is not a linear space, and thus $\operatorname{im}(\mathrm{\bf L}_{(L,R)})\neq\mathrm{T}_{\mathcal{M}_{\leq r}}(LR^{\top})$. Applying Theorem 5.1 shows the necessity of $\operatorname{rank}(L)=\operatorname{rank}(R)=r$.
Then we turn to the “if ” part. In preparation, we present the computations of $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ in this context:
$\mathrm{\bf L}_{(L,R)}:(\dot{L},\dot{R})\mapsto\dot{L}R^{\top}+L\dot{R}^{\top},\ \mathrm{\bf Q}_{(L,R),(\dot{L},\dot{R})}:(L^{\prime},R^{\prime})\mapsto 2\dot{L}\dot{R}^{\top}+L{R}^{\prime\top}+L^{\prime}R^{\top}.$ (6.9)
Let $X=LR^{\top}$ and the SVD of $X$ be $X=U\varSigma V^{\top}$. Then $\operatorname{rank}(X)=r$ indicates that $(L,R)$ can be expressed as $(L,R)=(UB,VC)$ for some invertible $B,C\in\mathbb{R}^{r\times r}$ satisfying $BC^{\top}=\varSigma$. Given any $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ written as $\eta=U\dot{A}V^{\top}+U\dot{B}V_{\bot}^{\top}+U_{\bot}\dot{C}V^{\top}$, we can construct $\dot{L}=U\dot{A}C^{-\top}+U_{\bot}\dot{C}C^{-\top}$ and $\dot{R}=V_{\bot}\dot{B}^{\top}B^{-\top}$ to obtain $\mathrm{\bf L}_{(L,R)}(\dot{L},\dot{R})=\eta$. Hence, the arbitrariness of $\eta$ concludes that $\operatorname{im}(\mathrm{\bf L}_{(L,R)})=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$. Subsequently, we aim to show that
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)=\mathrm{\bf Q}_{(L,R),(\dot{L},\dot{R})}(L^{\prime},R^{\prime})+\operatorname{im}(\mathrm{\bf L}_{(L,R)})$ (6.10)
for some $(L^{\prime},R^{\prime})$, thereby implying the inclusion (5.4). On the one hand, according to the closed-form expression (3.10), we have
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)=\eta X^{\dagger}\eta+\mathrm{T}_{\mathcal{M}_{r}}(X)=U_{\bot}\dot{C}\varSigma^{-1}\dot{B}V_{\bot}^{\top}+\mathrm{T}_{\mathcal{M}_{r}}(X),$ (6.11)
where the second equality follows by substituting the expressions of $\eta$ and $X^{\dagger}=V\varSigma^{-1}U^{\top}$. On the other hand, taking the expressions of $(\dot{L},\dot{R})$ into (6.9) leads to
$\mathrm{\bf Q}_{(L,R),(\dot{L},\dot{R})}(L^{\prime},R^{\prime})\equiv U_{\bot}\dot{C}C^{-\top}B^{-1}\dot{B}V_{\bot}^{\top}\ (\bmod\ \mathrm{T}_{\mathcal{M}_{r}}(X)),$
where $\equiv$” indicates equality modulo the tangent space. Substituting $BC^{\top}=\varSigma$ and combining the result with (6.11) yields (6.10). Therefore, applying Theorem 5.1 verifies the $2\!\Rightarrow\!2$” property at $(L,R)$ whenever $\operatorname{rank}(LR^{\top})=r$.
Proposition 9
The desingularization of $\mathcal{M}_{\leq r}$ given by (Desing) satisfies “ $2\!\Rightarrow\!2$” at $(X,G)$ if and only if $\operatorname{rank}(X)=r$.
Proof
It has been proved in (levin2025effect, Proposition 2.9) that $\operatorname{im}(\mathrm{\bf L}_{(X,G)})=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ if and only if $\operatorname{rank}(X)=r$, which validates the “only if ” part of our proposition.
[…]
Let $\psi(L,R)=X$, i.e., $X=LR^{\top}$, and the SVD of $X$ be $X=U\varSigma V^{\top}$ (which indicates that $(L,R)=(UB,VC)$ with $B=\varSigma V^{\top}R$ and $C=\varSigma U^{\top}L(L^{\top}L)^{-1}$). The computations of $\mathrm{\bf L}$ and $\mathrm{\bf Q}$ directly follows (6.9). Given any $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ written as $\eta=U\dot{A}V^{\top}+U\dot{B}V_{\bot}^{\top}+U_{\bot}\dot{C}V^{\top}$, we can construct $\dot{L}=U\dot{A}C^{-\top}+U_{\bot}\dot{C}C^{-\top}$ and $\dot{R}=V_{\bot}\dot{B}^{\top}B^{-\top}\in\mathrm{T}_{\mathrm{St}(n,r)}(R)$ to obtain $\mathrm{\bf L}_{(L,R)}(\dot{L},\dot{R})=\eta$. Hence, the arbitrariness of $\eta$ concludes that $\operatorname{im}(\mathrm{\bf L}_{(L,R)})=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$.
Then, we move on to prove (6.10), thereby verifying condition (5.4). The analysis parallels that in Proposition 8. Therefore, applying Theorem 5.1 confirms the $2\!\Rightarrow\!2$” property of $(\mathbb{R}^{m\times r}\times\mathrm{St}(n,r),\psi)$ at $(L,R)$, while Proposition 5 further implies that $(\mathcal{M}_{\mathrm{desing}},\phi_{\mathrm{desing}})$ also satisfies the $2\!\Rightarrow\!2$” property at $(X,G)$.
Riemannian trust-region algorithms are guaranteed to accumulate at second-order stationary points on smooth manifolds absil2007trustRTR, and thus Propositions 8 and 9 reveal that smooth parameterizations may find second-order stationary points on $\mathcal{M}_{\leq r}$, provided that the returned point happens to have rank $r$. From this perspective, the NP-hardness characterized in section 6.2 can be essentially attributed to the singularities of $\mathcal{M}_{\leq r}$, i.e., points with rank strictly lower than $r$.
7 Geometry of the graph of the normal cone mapping
Viewing the Mordukhovich normal cone induced by the determinantal variety as a set-valued mapping, that is,
$\mathrm{N}_{\mathcal{M}_{\leq r}}:\mathbb{R}^{m\times n}\rightrightarrows\mathbb{R}^{m\times n}:\,X\mapsto\mathrm{N}_{\mathcal{M}_{\leq r}}(X),$
we aim to give an explicit formula for the Mordukhovich normal cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$. Specifically, according to (2.15), the graph of $\mathrm{N}_{\mathcal{M}_{\leq r}}$ can be characterized by
$\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}=\left\{{(X,Y)\in\mathcal{M}_{\leq r}\times\mathbb{R}^{m\times n}\mid Y\in\mathrm{N}_{\mathcal{M}_{s}}(X),\ \operatorname{rank}(Y)\leq\min\{m,n\}-r}\right\},$
where we denote $s=\operatorname{rank}(X)$.
In this section, we investigate the variational geometry of $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$, with the derivation illustrated below.
$\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}$$\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}$$\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}$$\mathrm{polar}$$\lim$
In detail, we first characterize the Bouligand tangent cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ in Theorem 7.1, and then take the polar operation to obtain the Fréchet normal cone in Corollaries 1-2. Consequently, in Theorem 7.2, the Mordukhovich normal cone to $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ is identified as the outer limit of the developed Fréchet normal cone.
In preparation, we introduce some notation used throughout this section. We denote $k:=\min\{m,n\}$. Given $(X,Y)$, the ranks of $X$ and YY are represented by $s$ and $k$-$\ell$, respectively, and specifically, the condition $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ implies that $0\leq s\leq r\leq\ell\leq k$.
7.1 Bouligand tangent cone to the graph
As a preview, we note that the derived tangent cone (7.4) is characterized via a parameterization built upon the SVD of the reference point $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$. It is admitted that the coupling relationships among parameters are slightly involved, and thus we extract part of them in the following lemma, which appears technical but forms the basis of Theorem 7.1.
Lemma 1
Given $0\leq s\leq r\leq k=\min\{m,n\}$ and $R\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$ with the compact SVD, $R=U_{R}\varSigma_{R}V_{R}^{\top}$. Suppose that $K\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s}$ satisfies $K^{\top}R=0$ and $RK^{\top}=0$. Then, for any
$\tilde{D}\in\{D\in\mathbb{R}^{(m-s)\times(n-s)}\mid K^{\top}DV_{R\bot}=0,\,U_{R\bot}^{\top}DK^{\top}=0\}\cap\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R),$ (7.1)
there exist smooth curves $(R(t),K(t))\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}\times\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s}$ such that $R^{\prime}(0)=\tilde{D}$, $(R(0),K(0))=(R,K)$, $K(t)^{\top}R(t)=0$ and $R(t)K(t)^{\top}=0$.
Proof
Let $\operatorname{rank}(R)=k$-$\ell$ with $r\leq\ell\leq k$. Denote the set on the right side of (7.1) by $\mathcal{N}(R,K)$, and we can give an explicit characterization for it, that is,
$\mathcal{N}(R,K)=\left\{{\zeta_{1}+\zeta_{2}\ \left|\,\begin{array}[]{l}\zeta_{1}\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{k-\ell}}(R),\ \zeta_{2}\in\mathrm{N}_{\mathbb{R}^{(m-s)\times(n-s)}_{k-\ell}}(R),\\[5.69054pt] \operatorname{rank}(\zeta_{2})\leq\ell-r,\ K^{\top}\zeta_{2}=0,\ \zeta_{2}K^{\top}=0\end{array}\right.}\right\},$ (7.2)
which can be verified by considering the expressions from (2.11) to (2.13).
Then, for $\tilde{D}=\zeta_{1}+\zeta_{2}\in\mathcal{N}(R,K)$, we prove the lemma by construction. In view of (7.2), given any tangent vector $\zeta_{1}$ to the analytic manifold $\mathbb{R}^{(m-s)\times(n-s)}_{k-\ell}$ at $R$, there exists an analytic curve $R_{1}(t)$ on the manifold with $R_{1}(0)=R$ and $R_{1}^{\prime}(0)=\zeta_{1}$. Subsequently, (bunse1991analyticSVD, Theorem 1) reveals that $R_{1}(t)$ admits an analytic singular value decomposition, i.e.,
$R_{1}(t)=\left[{U_{R}(t)\ U_{R\bot}(t)}\right]\left[\begin{matrix}\varSigma_{R}\left(t\right)&0\\ 0&\varSigma_{R\bot}\left(t\right)\\ \end{matrix}\right]\left[{V_{R}(t)\ V_{R\bot}(t)}\right]^{\top}.$ (7.3)
Without loss of generality, suppose $U_{R}(0)\ U_{R\bot}(0)]=[U_{R}\ U_{R\bot}]$, $\varSigma_{R}(0)=\varSigma_{R}$, and $V_{R}(0)\ V_{R\bot}(0)]=[V_{R}\ V_{R\bot}]$. Since $\operatorname{rank}(R_{1}(t))\equiv k$-$\ell$ and $\operatorname{rank}(\varSigma_{R}(0))=k$-$\ell$, we can find an interval $\left({-\varepsilon,+\varepsilon}\right)$ such that $\varSigma_{R\bot}(t)\equiv 0$, which means $R_{1}(t)=U_{R}(t)\varSigma_{R}(t)V^{\top}_{R}(t)$.
The next step is to additionally introduce the direction $\zeta_{2}$ in (7.2). Let $\operatorname{rank}(K)=r-c\leq r-s$. The conditions $K^{\top}R=0$ and $RK^{\top}=0$ indicate that $K=U_{R\bot}U_{R\bot}^{\top}KV_{R\bot}V_{R\bot}^{\top}$, and thus it admits the decomposition $K=U_{R\bot}U_{K}\varSigma_{K}V_{K}^{\top}V_{R\bot}^{\top}$ for some $U_{K}\in\mathrm{St}(m-k+\ell-s,r-c)$ and $V_{K}\in\mathrm{St}(n-k+\ell-s,r-c)$, the complements of which are denoted by $U_{K\bot}$ and $V_{K\bot}$, respectively. For $\zeta_{2}\in\mathrm{N}_{\mathbb{R}^{(m-s)\times(n-s)}_{k-\ell}}(R)$ of $\operatorname{rank}(\zeta_{2})\leq\ell-r$, the conditions $K^{\top}\zeta_{2}=0$ and $\zeta_{2}K^{\top}=0$ imply that $\zeta_{2}$ can be parameterized by $\zeta_{2}=U_{R\bot}U_{K\bot}{Z}V_{K\bot}^{\top}V_{R\bot}^{\top}$ for some $Z$ with $\operatorname{rank}(Z)=\operatorname{rank}(\zeta_{2})$.
Collecting $R_{1}(t)$ given in (7.3) and $R_{2}(t):=tU_{R\bot}(t)U_{K\bot}ZV_{K\bot}^{\top}V_{R\bot}(t)^{\top}$, we obtain $R(t):=R_{1}(t)+R_{2}(t)\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$ and $K(t):=U_{R\bot}(t)U_{K}\varSigma_{K}V_{K}^{\top}V_{R\bot}(t)^{\top}\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s}$ satisfying $R^{\prime}(0)=\tilde{D}=\zeta_{1}+\zeta_{2}$ and $(R(0),K(0))=(R,K)$. Moreover, $K(t)^{\top}R(t)=0$ and $R(t)K(t)^{\top}=0$ hold in the interval $(-\varepsilon,+\varepsilon)$.
We now proceed to derive the tangent cone to the graph.
Theorem 7.1(Bouligand tangent cone)
Given $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ with $\operatorname{rank}(X)=s$, the SVD $X=U\varSigma V^{\top}$, and $Y=U_{\bot}RV_{\bot}^{\top}$ for some $R\in\mathbb{R}^{(m-s)\times(n-s)}$. Suppose that $R$ admits the compact SVD $R=U_{R}\varSigma_{R}V_{R}^{\top}$. Then, the Bouligand tangent cone at $(X,Y)$ can be expressed as follows,
$\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\left\{{(\eta,\xi)\ \left|\,\begin{array}[]{l}\eta=UAV^{\top}+U_{\bot}B\varSigma V^{\top}+U\varSigma C^{\top}V_{\bot}^{\top}+U_{\bot}KV^{\top}_{\bot}\\ \xi=U_{\bot}DV_{\bot}^{\top}-UB^{\top}RV_{\bot}^{\top}-U_{\bot}RCV^{\top},\\ A\in\mathbb{R}^{s\times s},\,B\in\mathbb{R}^{(m-s)\times s},\,C\in\mathbb{R}^{(n-s)\times s},\\ D\in\mathrm{T}_{\mathbb{R}_{\leq k-r}^{(m-s)\times(n-s)}}(R),\,K\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s},\\ K^{\top}R=0,\,RK^{\top}=0,\\ K^{\top}DV_{R\bot}=0,\ U_{R\bot}^{\top}DK^{\top}=0\end{array}\right.}\right\}\!$.\!\!\!\!\! (7.4)
Proof
We begin by proving the $\supseteq$” part of (7.4). Given $K\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s}$ satisfying $K^{\top}R=0$ and $RK^{\top}=0$, we can construct curves $R(t)\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$ and $K(t)\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq r-s}$ as stated in Lemma 1. Consider, in addition, curves $U(t)\ U_{\bot}(t)]\in\mathcal{O}(m)$ and $V(t)\ V_{\bot}(t)]\in\mathcal{O}(n)$ passing through $U\ U_{\bot}]$ and $V\ V_{\bot}]$ at $t=0$, respectively. Then, we assemble the curves in the following manner:
$\gamma(t):=U(t)\varSigma(t)V(t)^{\top}+U_{\bot}(t)\cdot tK(t)\cdot V_{\bot}(t)^{\top}\ \ \text{and}\ \ \beta(t):=U_{\bot}(t)R(t)V_{\bot}(t)^{\top},$
where $\varSigma(t)\in\mathbb{R}^{s\times s}$ with $\varSigma(0)=\varSigma$.
Note that $\beta(t)^{\top}\gamma(t)=0$, $\gamma(t)\beta(t)^{\top}=0$, $\operatorname{rank}(\gamma(t))\leq r$, and $\operatorname{rank}(\beta(t))\leq k-r$. Hence, we obtain a smooth curve $\alpha(t):=(\gamma(t),\beta(t))$ in $\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$. Differentiating $\alpha(t)$ at $t=0$ yields $\alpha^{\prime}(0)=(\gamma^{\prime}(0),\beta^{\prime}(0))\in\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ with
$\gamma^{\prime}(0)$ $=U^{\prime}(0)\varSigma V^{\top}+U\varSigma^{\prime}(0)V^{\top}+U\varSigma V^{\prime}(0)^{\top}+U_{\bot}KV_{\bot}^{\top},$ (7.5)
$\beta^{\prime}(0)$ $=U_{\bot}^{\prime}(0)RV_{\bot}^{\top}+U_{\bot}R^{\prime}(0)V_{\bot}^{\top}+U_{\bot}RV_{\bot}^{\prime}(0)^{\top}.$
Taking into account the tangent space $\mathrm{T}_{\mathcal{O}(m)}([U\ U_{\bot}])$ (see (absil2008optimization, §3.5)), for any $\Omega_{u}\in\mathrm{Skew}(s)$, $\Omega_{u\bot}\in\mathrm{Skew}(m-s)$, and $B\in\mathbb{R}^{(m-s)\times s}$, it is reasonable to construct the curve such that $U^{\prime}(0)=U\Omega_{u}+U_{\bot}B$ and $U_{\bot}^{\prime}(0)=U_{\bot}\Omega_{u\bot}-UB^{\top}$. Similarly, we can arrange to have $V^{\prime}(0)=V\Omega_{v}+V_{\bot}C$ and $V_{\bot}^{\prime}(0)=V_{\bot}\Omega_{v\bot}-VC^{\top}$ for any $\Omega_{v}\in\mathrm{Skew}(s)$, $\Omega_{v\bot}\in\mathrm{Skew}(n-s)$, and $C\in\mathbb{R}^{(n-s)\times s}$. Moreover, notice that $\varSigma(t)\in\mathbb{R}^{s\times s}$ is unconstrained, and thus $\varSigma^{\prime}(0)$ is allowed to be any $L\in\mathbb{R}^{s\times s}$. Substituting the discussed quantities into (7.5) leads to
$\gamma^{\prime}(0)$ $=U(\Omega_{u}\varSigma+L-\varSigma\Omega_{v})V^{\top}+U_{\bot}B\varSigma V^{\top}+U\varSigma C^{\top}V_{\bot}^{\top}+U_{\bot}KV_{\bot}^{\top},$ (7.6)
$\beta^{\prime}(0)$ $=U_{\bot}(\Omega_{u\bot}R+R^{\prime}(0)-R\Omega_{v\bot})V_{\bot}^{\top}-UB^{\top}RV_{\bot}^{\top}-U_{\bot}RCV^{\top}.$
We denote $A=\Omega_{u}\varSigma+L-\varSigma\Omega_{v}$ and $D=\Omega_{u\bot}R+R^{\prime}(0)-R\Omega_{v\bot}$. The arbitrariness of $L\in\mathbb{R}^{s\times s}$ reveals that of $A\in\mathbb{R}^{s\times s}$. Additionally, the conditions $K^{\top}R=0$ and $RK^{\top}=0$, together with the freedom in choosing $R^{\prime}(0)\in\mathcal{N}(R,K)$ (cf. Lemma 1), reveal the arbitrariness of $D\in\mathcal{N}(R,K)$, which concludes the $\supseteq$” part.
Next, we turn to show the $\subseteq$” in (7.4). Given any $(\eta,\xi)\in\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$, by definition of the Bouligand tangent cone, it admits sequences $\{t_{k}\}\subseteq\mathbb{R}_{+}$ and $\{(\eta_{k},\xi_{k})\}\subseteq\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times n}$ such that $t_{k}\downarrow 0$, $(\eta_{k},\xi_{k})\to(\eta,\xi)$, and $(X,Y)+t_{k}(\eta_{k},\xi_{k})\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$, or equivalently,
$X+t_{k}\eta_{k}\in\mathcal{M}_{\leq r}\ \ \text{and}\ \ Y+t_{k}\xi_{k}\in\mathrm{N}_{\mathcal{M}_{\leq r}}(X+t_{k}\eta_{k}).$ (7.7)
We note that $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$, and according to the expression (2.13), $(\eta,\xi)$ has the following form,
$\eta=$ $\ UA_{\eta}V^{\top}+U_{\bot}B_{\eta}V^{\top}+UC_{\eta}V^{\top}_{\bot}+U_{\bot}KV^{\top}_{\bot},$ (7.8)
$\xi=$ $\ UA_{\xi}V^{\top}+U_{\bot}B_{\xi}V^{\top}+UC_{\xi}V^{\top}_{\bot}+U_{\bot}DV^{\top}_{\bot},$
where $\operatorname{rank}(K)\leq r-s$. Then we are going to determine the relationships between the involved parameters (e.g., $A_{\eta},A_{\xi}$).
Firstly, the requirement (7.7) reveals that
$(X+t_{k}\eta_{k})^{\top}(Y+t_{k}\xi_{k})=t_{k}\eta_{k}^{\top}Y+t_{k}X^{\top}\xi_{k}+t_{k}^{2}\eta_{k}^{\top}\xi_{k}=0.$ (7.9)
Dividing the equation by $t_{k}$ and letting $k\to+\infty$ yield $\eta^{\top}Y+X^{\top}\xi=0$, i.e., $VB_{\eta}^{\top}RV_{\bot}+V_{\bot}^{\top}K^{\top}RV_{\bot}^{\top}+V\varSigma A_{\xi}V^{\top}+V\varSigma C_{\xi}V_{\bot}^{\top}=0$, which implies $C_{\xi}=-\varSigma^{-1}B_{\eta}^{\top}R$, $A_{\xi}=0$, and $K^{\top}R=0$. In a similar way, we can derive $B_{\xi}=-RC^{\top}_{\eta}\varSigma^{-1}$ and $RK^{\top}=0$. Substituting these equalities into (7.8) and comparing the expression with (7.4) reduce the task to verifying that $D\in\mathcal{N}(R,K)$ as defined in Lemma 1. To see this, noticing that $U_{\bot}^{\top}\xi_{k}V_{\bot}\to U_{\bot}^{\top}\xi V_{\bot}=D$, and
$U_{\bot}^{\top}(Y+t_{k}\xi_{k})V_{\bot}=R+t_{k}U_{\bot}^{\top}\xi_{k}V_{\bot}\in\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}$,
we have $D\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)$. Moreover, pre- and postmultiplying (7.9) by $V_{\bot}V_{\bot}^{\top}$ and $V_{\bot}V_{R\bot}$, respectively, we obtain $V_{\bot}V_{\bot}^{\top}\eta_{k}^{\top}\xi_{k}V_{\bot}V_{R\bot}=0$. Letting $k\to+\infty$ and incorporating the expression of $(\eta,\xi)$ (7.8) reveal that $V_{\bot}(K^{\top}U_{\bot}^{\top}+C^{\top}_{\eta}U^{\top})(UC_{\xi}+U_{\bot}D)V_{R\bot}=0$, which is simplified to $V_{\bot}(K^{\top}D+C_{\eta}^{\top}C_{\xi})V_{R\bot}=0$. Consequently, we use $C_{\xi}V_{R\bot}=-\varSigma^{-1}B_{\eta}^{\top}RV_{R\bot}=0$ to derive $K^{\top}DV_{R\bot}=0$, and similarly, we can also find that $U_{R\bot}^{\top}DK^{\top}=0$, indicating that $D\in\mathcal{N}(R,K)$. Therefore, we have identified all the relationships to conclude that $(\eta,\xi)$ belongs to the set on the right side of (7.4).
7.2 Fréchet normal cone to the graph
As shown in (7.4), the characterization of the tangent cone at $(X,Y)$ is relevant to the rank of $X$. This observation leads to different treatments of $\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ in two cases, depending on whether $X$ attains rank $r$, which mirrors the first-order result in (2.14). Accordingly, Corollaries 1 and 2 address the cases $\operatorname{rank}(X)<r$ and $\operatorname{rank}(X)=r$, respectively.
Corollary 1
Given $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ with $\operatorname{rank}(X)=s<r$, the SVD $X=U\varSigma V^{\top}$, and $Y=U_{\bot}RV_{\bot}^{\top}$ for some $R\in\mathbb{R}^{(m-s)\times(n-s)}$. Suppose that $\operatorname{rank}(R)=k$-$\ell$ and it admits the compact SVD $R=U_{R}\varSigma_{R}V_{R}^{\top}$. Then, the Fréchet normal cone at $(X,Y)$ can be expressed as follows,
$\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\left\{{(\hat{\upsilon},\hat{\omega})\ \left|\,\begin{array}[]{l}\hat{\upsilon}=U_{\bot}U_{R}\varSigma_{R}\hat{C}_{1}^{\top}V^{\top}+U\hat{B}_{1}^{\top}\varSigma_{R}V_{R}^{\top}V_{\bot}^{\top}\\ \hskip 19.91692pt+U_{\bot}ZV^{\top}_{\bot},\\ \hat{\omega}=U\hat{A}V^{\top}+U_{\bot}U_{R}\hat{B}_{1}\varSigma V^{\top}+U_{\bot}U_{R\bot}\hat{B}_{2}V^{\top}\\ \hskip 19.91692pt+U\varSigma\hat{C}_{1}V_{R}^{\top}V_{\bot}^{\top}+U\hat{C}_{2}V_{R\bot}^{\top}V_{\bot}^{\top}+U_{\bot}\hat{Z}V_{\bot}^{\top}\\ \hat{A}\in\mathbb{R}^{s\times s},\,\hat{B}_{1}\in\mathbb{R}^{(k-\ell)\times s},\,\hat{B}_{2}\in\mathbb{R}^{(m-k+\ell-s)\times s}\\ \hat{C}_{1}\in\mathbb{R}^{s\times(k-\ell)},\,\hat{C}_{2}\in\mathbb{R}^{s\times(n-k+\ell-s)},\\ {Z}\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{k-\ell}}(R),\,\hat{Z}\in\hat{\mathrm{N}}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)\end{array}\right.}\right\}$. (7.10)
Proof
Taking the polar operation on both sides of (7.4) will yield the desired Fréchet normal cone. Specifically, notice that $(\eta,\xi)$ belongs to $\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ if and only if $(-\eta,-\xi)$ is in it, and thus $(\hat{\upsilon},\hat{\omega})\in\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ is equivalent to $\langle{(\hat{\upsilon},\hat{\omega}),(\eta,\xi)}\rangle=0$ for all $(\eta,\xi)\in\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$.
Given $(\hat{\upsilon},\hat{\omega})\in\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times n}$, we can parameterize them as follows,
$\hat{\upsilon}$ $=U\hat{A}_{\upsilon}V^{\top}+U_{\bot}\hat{B}_{\upsilon}V^{\top}+U\hat{C}_{\upsilon}V_{\bot}^{\top}+U_{\bot}ZV_{\bot}^{\top},$ (7.11)
$\hat{\omega}$ $=U\hat{A}_{\omega}V^{\top}+U_{\bot}\hat{B}_{\omega}V^{\top}+U\hat{C}_{\omega}V_{\bot}^{\top}+U_{\bot}\hat{Z}V_{\bot}^{\top}.$
Considering $(\eta,\xi)$ as expressed in (7.4), we have
$\langle{(\hat{\upsilon},\hat{\omega}),(\eta,\xi)}\rangle=$ $\ \langle{\hat{A}_{\upsilon},A}\rangle+\langle{\hat{A}_{\omega},0}\rangle+\langle{\hat{B}_{\upsilon}\varSigma-R\hat{C}_{\omega}^{\top},B}\rangle$ (7.12)
$+\langle{\hat{C}_{\upsilon}^{\top}\varSigma-R^{\top}\hat{B}_{\omega},C}\rangle+\langle{Z,K}\rangle+\langle{\hat{Z},D}\rangle$
Letting $A,B,C,K$ be zero matrices of corresponding shapes, the orthogonal requirement $\langle(\hat{\upsilon},\hat{\omega}),(\eta,\xi)\rangle=0$ implies that $\langle{\hat{Z},D}\rangle=0$ for all $D\in\mathrm{T}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)$; similar processes conclude that $\hat{A}_{\upsilon}=0$, $\hat{A}_{\omega}$ is freely chosen from $\mathbb{R}^{s\times s}$, and $\langle{Z,K}\rangle=0$ for all $K$ described in (7.4). Moreover, we let $A,K,D$ be zero matrices to find that
$\langle{\hat{B}_{\upsilon}\varSigma-R\hat{C}^{\top}_{\omega},B}\rangle+\langle{\hat{C}_{\upsilon}^{\top}\varSigma-R^{\top}\hat{B}_{\omega},C}\rangle=0,\ \ \text{for all}\ (B,C)\in\mathbb{R}^{(m-s)\times s}\times\mathbb{R}^{(n-s)\times s},$
which, therefore, enforces that $\hat{B}_{\upsilon}=R\hat{C}^{\top}_{\omega}\varSigma^{-1}$ and $\hat{C}_{\upsilon}=\varSigma^{-1}\hat{B}_{\omega}^{\top}R$. Taking the relationships, and substituting $\hat{C}_{\omega}=\varSigma\hat{C}_{1}V^{\top}_{R}+\hat{C}_{2}V_{R\bot}^{\top}$ and $\hat{B}_{\omega}=U_{R}\hat{B}_{1}\varSigma+U_{R\bot}\hat{B}_{2}$ into (7.11) lead to the formulation (7.10).
We then turn to the case $\operatorname{rank}(X)=r$. The main difference is that when $X$ attains rank $r$, the parameter $K$ in (7.4) is forced to vanish, implying that the component of $\eta$ spanned by $U_{\bot}$ and $V_{\bot}$ disappears.
Corollary 2
Given $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ with $\operatorname{rank}(X)=r$, the SVD $X=U\varSigma V^{\top}$, and $Y=U_{\bot}RV_{\bot}^{\top}$ for some $R\in\mathbb{R}^{(m-s)\times(n-s)}$. Suppose that $\operatorname{rank}(R)=k$-$\ell$ and it admits the compact SVD $R=U_{R}\varSigma_{R}V_{R}^{\top}$. Then, the Fréchet normal cone at $(X,Y)$ can be expressed as follows,
$\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\left\{{(\hat{\upsilon},\hat{\omega})\ \left|\,\begin{array}[]{l}\hat{\upsilon}=U_{\bot}U_{R}\varSigma_{R}\hat{C}_{1}^{\top}V^{\top}+U\hat{B}_{1}^{\top}\varSigma_{R}V_{R}^{\top}V_{\bot}^{\top}\\ \hskip 19.91692pt+U_{\bot}ZV^{\top}_{\bot},\\ \hat{\omega}=U\hat{A}V^{\top}+U_{\bot}U_{R}\hat{B}_{1}\varSigma V^{\top}+U_{\bot}U_{R\bot}\hat{B}_{2}V^{\top}\\ \hskip 19.91692pt+U\varSigma\hat{C}_{1}V_{R}^{\top}V_{\bot}^{\top}+U\hat{C}_{2}V_{R\bot}^{\top}V_{\bot}^{\top}+U_{\bot}\hat{Z}V_{\bot}^{\top}\\ \hat{A}\in\mathbb{R}^{s\times s},\,\hat{B}_{1}\in\mathbb{R}^{(k-\ell)\times s},\,\hat{B}_{2}\in\mathbb{R}^{(m-k+\ell-s)\times s}\\ \hat{C}_{1}\in\mathbb{R}^{s\times(k-\ell)},\,\hat{C}_{2}\in\mathbb{R}^{s\times(n-k+\ell-s)},\\ Z\in\mathbb{R}^{(m-r)\times(n-r)},\,\hat{Z}\in\hat{\mathrm{N}}_{\mathbb{R}^{(m-s)\times(n-s)}_{\leq k-r}}(R)\end{array}\right.}\right\}$. (7.13)
Proof
The derivation follows the same reasoning as in the proof of Corollary 1, except that $\operatorname{rank}(X)=r$ enforces $K=0$ in (7.12). Consequently, $Z$ becomes unconstrained in $\mathbb{R}^{(m-r)\times(n-r)}$.
It is worth noting that $\mathcal{M}_{r}$ is relatively open in $\mathcal{M}_{\leq r}$ olikier2022continuity. That is, in a neighborhood of a point $X$ with $\operatorname{rank}(X)=r$, the determinantal variety $\mathcal{M}_{\leq r}$ coincides with the smooth manifold $\mathcal{M}_{r}$. Since $\mathcal{M}_{r}$ is an embedded submanifold of $\mathbb{R}^{m\times n}$, Theorem 6.23 in lee2012manifolds implies that the normal bundle $\mathrm{N}\mathcal{M}_{r}$ is an embedded submanifold of $\mathbb{R}^{m\times n}\times\mathbb{R}^{m\times n}$ of dimension $mn$. Consequently, when $\operatorname{rank}(X)=r$, we have $\mathrm{T}_{\mathrm{N}\mathcal{M}_{r}}(X,Y)=\mathrm{T}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$, and $\mathrm{N}_{\mathrm{N}\mathcal{M}_{r}}(X,Y)=\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)=\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$, as characterized by Corollary 2.
7.3 Mordukhovich normal cone to the graph
We are now in a position to develop the Mordukhovich normal cone, for which two auxiliary lemmas are required.
Let $\mathbb{R}_{\geq}^{d}$ be the set of vectors in $\mathbb{R}^{d}$ with elements being positive and arranged in a non-increasing order, i.e.,
$\mathbb{R}_{\geq}^{d}:=\{x\in\mathbb{R}^{d}\mid x_{1}\geq x_{2}\geq\cdots\geq x_{d}>0\}.$
We define the mapping $\mathfrak{D}:\mathbb{R}_{\geq}^{d_{1}}\times\mathbb{R}_{\geq}^{d_{2}}\rightarrow\mathbb{R}^{d_{1}\times d_{2}}$ as follows,
$\mathfrak{D}(x,y)_{jt}:=\frac{x_{j}}{x_{j}+y_{t}},\ \ \text{for}\ \ j=1,2,\dots,d_{1}\ \text{and}\ t=1,2,\ldots,d_{2}.$
Then, the mapping $\mathfrak{D}$ introduces the following set,
$\mathrm{\Theta}(d_{1},d_{2}):=\left\{{\lim_{i\rightarrow\infty}\mathfrak{D}(z_{1}^{i},z_{2}^{i})\in\mathbb{R}^{d_{1}\times d_{2}}\mid z_{1}^{i}\rightarrow 0,z_{2}^{i}\rightarrow 0,z_{1}^{i}\in\mathbb{R}_{\geq}^{d_{1}},z_{2}^{i}\in\mathbb{R}_{\geq}^{d_{2}}}\right\}.$ (7.14)
In fact, the set $\mathrm{\Theta}$ helps characterize the asymptotic behavior of two families of positive singular values (see the proof of Theorem 7.2). Additionally, we remark that the definitions of $\mathbb{R}_{\geq}^{d}$, $\mathfrak{D}$, and $\mathrm{\Theta}$ are inspired by (ding2014SDCMPCC, §3.2) and (wu2014SDCMPCC, §3.2), where related concepts were introduced for studying normal cones to $\mathrm{gph}\mathrm{N}_{\mathcal{S}^{+}(n)}$.
The subsequent lemma identifies a basic relation in matrix computation, which was proved in (wu2014SDCMPCC, Lemma 2.2).
Lemma 2
Given vectors $b\in\mathbb{R}^{d_{1}}_{\geq}$, $q\in\mathbb{R}^{d_{2}}_{\geq}$ and matrices $B,Q\in\mathbb{R}^{d_{1}\times d_{2}}$, it holds that
$\operatorname{Diag}\left(b\right)B=Q\operatorname{Diag}\left(q\right)\Longleftrightarrow\mathfrak{D}(b,q)\odot B+\left(\mathfrak{D}(b,q)-{\bm{1}}_{d_{1}\times d_{2}}\right)\odot Q=0,$
where $\odot$ denotes the Hadamard product and ${\bm{1}}_{d_{1}\times d_{2}}$ is the all-ones matrix of size $d_{1}\times d_{2}$.
Note that the Mordukhovich normal cone is defined through the outer limit (2.3), and thus we introduce the next lemma to capture the behavior of sequences convergent to $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$. Specifically, it is shown that the convergence of $(X_{i},Y_{i})$} to $(X,Y)$ induces the convergence of the subspaces spanned by $X_{i}$ and $Y_{i}$ as well as of their orthogonal complements (possibly after taking a subsequence), which indeed extends the idea of (olikier2022continuity, Lemma 4.2).
Lemma 3
[…]
$\{(U_{i_{j}},\tilde{U}_{i_{j}\bot},\breve{U}_{i_{j}\bot},\tilde{U}_{i_{j}Y\bot},U_{i_{j}Y})\}_{j\in\mathbb{N}}$ $\subseteq\ \mathcal{O}(m),$
$\{(V_{i_{j}},\tilde{V}_{i_{j}\bot},\breve{V}_{i_{j}\bot},\tilde{V}_{i_{j}Y\bot},V_{i_{j}Y})\}_{j\in\mathbb{N}}$ $\subseteq\ \mathcal{O}(n),$
satisfying the following properties.
  • For all $i_{j}$, $\operatorname{im}[U_{i_{j}}\ \tilde{U}_{i_{j}\bot}]=\operatorname{im}X_{i_{j}}$, $\operatorname{im}[V_{i_{j}}\ \tilde{V}_{i_{j}\bot}]=\operatorname{im}X^{\top}_{i_{j}}$, $\operatorname{im}[U_{i_{j}Y\bot}\ \tilde{U}_{i_{j}Y}]=\operatorname{im}Y_{i_{j}}$, and $\operatorname{im}[V_{i_{j}Y\bot}\ \tilde{V}_{i_{j}Y}]=\operatorname{im}Y^{\top}_{i_{j}}$.
  • It holds that
    $\lim_{j\to\infty}[U_{i_{j}}\ \tilde{U}_{i_{j}\bot}\ \breve{U}_{i_{j}\bot}\ \tilde{U}_{i_{j}Y\bot}\ U_{i_{j}Y}]$ $=[U\ \tilde{U}_{\bot}\ \breve{U}_{\bot}\ \tilde{U}_{Y\bot}\ U_{Y}],$ (7.15)
    $\lim_{j\to\infty}[V_{i_{j}}\ \tilde{V}_{i_{j}\bot}\ \breve{V}_{i_{j}\bot}\ \tilde{V}_{i_{j}Y\bot}\ V_{i_{j}Y}]$ $=[V\ \tilde{V}_{\bot}\ \breve{V}_{\bot}\ \tilde{V}_{Y\bot}\ V_{Y}].$
Proof
See Appendix E.
Recalling from the definition of the Mordukhovich normal cone (2.3), the direction $(\upsilon,\omega)\in\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ if and only if there exist a sequence $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ and the associated $(\hat{\upsilon}_{i},\hat{\omega}_{i})\in\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X_{i},Y_{i})$ such that
$\lim_{i\to\infty}(X_{i},Y_{i})=(X,Y)\ \ \text{and}\ \ \lim_{i\to\infty}(\hat{\upsilon}_{i},\hat{\omega}_{i})=(\upsilon,\omega).$ (7.16)
Then, we arrive at the characterization of the Mordukhovich normal cone.
Theorem 7.2(Mordukhovich normal cone)
[…]
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.17)
$\omega=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}A&C&C^{\omega}\\ B&0&0\\ B^{\omega}&0&\hat{Z}\\ \end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$
where $U^{+}:=[U\ \tilde{U}_{\bot}$], $U^{+}_{Y}:=[U_{Y}\ \tilde{U}_{Y\bot}]$, $V^{+}:=[V\ \tilde{V}_{\bot}$], $V^{+}_{Y}:=[V_{Y}\ \tilde{V}_{Y\bot}]$; $Z_{1}=0$ if ${\underline{r}}<r$ otherwise $Z_{1}$ is not restricted; $\hat{Z}=0$ if $\bar{r}>r$ otherwise $\hat{Z}$ is not restricted; and $B^{\upsilon}$, $B^{\omega}$, $C^{\upsilon}$, $C^{\omega}$ have the following forms,
$B^{\upsilon}=\left[\begin{matrix}\varSigma^{-1}G_{1}^{\top}\varSigma_{Y}&0\\ E_{1}^{\upsilon}&F_{1}^{\upsilon}\\ \end{matrix}\right],\,B^{\omega}=\left[\begin{matrix}G_{1}&0\\ E_{1}^{\omega}&F_{1}^{\omega}\\ \end{matrix}\right],\,C^{\upsilon}=\left[\begin{matrix}\varSigma_{Y}G_{2}^{\top}\varSigma^{-1}&E_{2}^{\upsilon}\\ 0&F_{2}^{\upsilon}\\ \end{matrix}\right],\,C^{\omega}=\left[\begin{matrix}G_{2}&E_{2}^{\omega}\\ 0&F_{2}^{\omega}\\ \end{matrix}\right],$ (7.18)
satisfying
$D\odot F_{1}^{\omega}+(D-{\bm{1}}_{(\ell-\bar{r})\times({\underline{r}}-s)})\odot(F^{\upsilon}_{1})^{\top}=0,$ (7.19)
$D^{\top}\odot F_{2}^{\omega}+(D^{\top}-{\bm{1}}_{({\underline{r}}-s)\times(\ell-\bar{r})})\odot(F^{\upsilon}_{2})^{\top}=0,$
for some $D\in\mathrm{\Theta}(\ell-\bar{r},{\underline{r}}-s)$. The dimensions of the matrix parameters are summarized in Table 2.
Table 2: Dimensions of matrix parameters in Theorem 7.2, where $\bar{r}-{\underline{r}}$ is abbreviated as $\Delta r$.
Param. AA $B$ $C$
Dim. ${\underline{r}}\times{\underline{r}}$} $(m-k+\Delta r)\times{\underline{r}}$} ${\underline{r}}\times(n-k+\Delta r)$}
Param. $D$ $G_{1}$, $G_{2}^{\top}$ $E_{1}^{\omega}$, $(E_{2}^{\omega})^{\top}$
Dim. $(\ell-\bar{r})\times({\underline{r}}-s)$} $(k-\ell)\times s$} $(\ell-\bar{r})\times s$}
Param. $E_{1}^{\upsilon}$, $(E_{2}^{\upsilon})^{\top}$ $F^{\upsilon}_{1}$, $(F_{1}^{\omega})^{\top}$, $(F^{\upsilon}_{2})^{\top}$, $F_{2}^{\omega}$ $Z_{1}$
Dim. $({\underline{r}}-s)\times(k-\ell)$} $({\underline{r}}-s)\times(\ell-\bar{r})$} $(m-k+\Delta r)\times(n-k+\Delta r)$
Param. $Z_{2}$ $Z_{3}$ $_{4}, Z^\hat{Z}$
Dim. $(m-k+\Delta r)\times(k-\bar{r})$ $(k-\bar{r})\times(n-k+\Delta r)$ $(k-\bar{r})\times(k-\bar{r})$
Proof
We remark that Corollaries 12 provide the parameterization in terms of $Y=U_{\bot}RV_{\bot}^{\top}$ and $R=U_{R}\varSigma_{R}V_{R}^{\top}$. To align with the notation used in this theorem, we can therefore identify $(U_{Y},\varSigma_{Y},V_{Y})=(U_{\bot}U_{R},\varSigma_{R},V_{\bot}V_{R})$; a similar argument applies when considering normal cones at $(X_{i},Y_{i})$.
To prove the “if” part of the theorem, we will construct a sequence $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ with $\operatorname{rank}(X_{i})={\underline{r}}$, $\operatorname{rank}(Y_{i})=k-\bar{r}$ and the associated $(\hat{\upsilon}_{i},\hat{\omega}_{i})$, which produces $(\upsilon,\omega)$ expressed by (7.17) in the manner of (7.16). To this end, taking into account the matrix $D\in\mathrm{\Theta}(\ell-\bar{r},{\underline{r}}-s)$ in (7.19), we assume that $D$ is generated by the sequences $\{z^{i}_{1}\}\subseteq\mathbb{R}_{\geq}^{\ell-\bar{r}}$ and $\{z^{i}_{2}\}\subseteq\mathbb{R}_{\geq}^{{\underline{r}}-s}$ as in (7.14). Subsequently, the $(X_{i},Y_{i})$ can be given by
$X_{i}=\left[U\,\,\tilde{U}_{\bot}\right]\left[\begin{matrix}\varSigma&0\\ 0&\operatorname{Diag}(z^{i}_{2})\\ \end{matrix}\right]\left[V\,\,\tilde{V}_{\bot}\right]^{\top}\!\!,\ Y_{i}=\left[U_{Y}\,\,\tilde{U}_{Y\bot}\right]\left[\begin{matrix}\varSigma_{Y}&0\\ 0&\operatorname{Diag}(z^{i}_{1})\\ \end{matrix}\right]\left[V_{Y}\,\,\tilde{V}_{Y\bot}\right]^{\top}.$
According to (7.10) and (7.13), we can design $(\hat{\upsilon}_{i},\hat{\omega}_{i})\in\hat{\mathrm{N}}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X_{i},Y_{i})$ in the following form,
$\hat{\upsilon}_{i}=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{array}[]{ccc}0&0&B_{i}^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}_{i}&Z_{3}&Z_{4}\end{array}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.20)
$\hat{\omega}_{i}=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}A&C&C_{i}^{\omega}\\ B&0&0\\ B_{i}^{\omega}&0&\hat{Z}\\ \end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$
where $Z_{1}=0$ if ${\underline{r}}<r$ otherwise $Z_{1}$ is not restricted, and $\hat{Z}=0$ if $\bar{r}>r$ otherwise $\hat{Z}$ is not restricted. Next, we detail the construction of $(B^{\omega}_{i},B^{\upsilon}_{i})$, which is required to satisfy the following coupling relation, as revealed by the expression (7.13),
$\left[\begin{matrix}\varSigma_{Y}&0\\ 0&\operatorname{Diag}(z^{i}_{1})\\ \end{matrix}\right]B_{i}^{\omega}=\left(B_{i}^{\upsilon}\right)^{\top}\left[\begin{matrix}\varSigma&0\\ 0&\operatorname{Diag}(z^{i}_{2})\\ \end{matrix}\right].$ (7.21)
We denote $D^{i}=\mathfrak{D}(z^{i}_{1},z^{i}_{2})$, $D^{i}_{jt}=D^{i}(j,t)$, and $D_{jt}=\lim_{i\to\infty}D^{i}_{jt}$, where “(j,t)(j,t)” refers to the entry in the $j$-th row and $t$-th column of a matrix. Then, two sequences of matrices, $\{F^{\omega}_{1i}\}_{i\in\mathbb{N}}$ and $\{F^{\upsilon}_{1i}\}_{i\in\mathbb{N}}$ are defined as follows,
[…]
for $i\in\mathbb{N}$, $j=1,2,\ldots,\ell-\bar{r}$, and $t=1,2,\ldots,{\underline{r}}-s$. Finally, the construction of $(B^{\omega}_{i},B^{\upsilon}_{i})$ is divided into four blocks, which is aligned with (7.18),
$B_{i}^{\omega}:=\left[\begin{matrix}G_{1}&\ \varSigma^{-1}_{Y}(E^{\upsilon}_{1})^{\top}\operatorname{Diag}(z_{2}^{i})\\ E_{1}^{\omega}&F_{1i}^{\omega}\\ \end{matrix}\right],\ \ \ B_{i}^{\upsilon}:=\,\,\left[\begin{matrix}\varSigma^{-1}G_{1}^{\top}\varSigma_{Y}&\ \varSigma^{-1}(E^{\omega}_{1})^{\top}\operatorname{Diag}(z^{i}_{1})\\ E_{1}^{\upsilon}&F_{1i}^{\upsilon}\\ \end{matrix}\right].$
It can be verified that $(B^{\omega}_{i},B^{\upsilon}_{i})$ satisfies (7.21) according to (7.22). Moreover, by the relationship (7.19), taking $i\to\infty$ in (7.22) shows that $\left(F_{1i}^{\omega},F_{1i}^{\upsilon}\right)\to(F_{1}^{\omega},F_{1}^{\upsilon})$. Additionally, the matrices $(C_{i}^{\omega},C^{\upsilon}_{i})$ can be constructed in a similar manner. Consequently, taking $i\to\infty$, the limit of $(\hat{\upsilon}_{i},\hat{\omega}_{i})$ in (7.20) yields the desirable $(\upsilon,\omega)$ in (7.17).
We then turn to the “only if” part of the theorem. Suppose that $(\upsilon,\omega)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}(X,Y)$ is generated by the sequence $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ in the sense of  (7.16), and then we will identify that the structure of $(\upsilon,\omega)$ coincides with (7.17). To this end, note that there exists $({\underline{r}},\bar{r})$ such that $s\leq{\underline{r}}\leq r\leq\bar{r}\leq\ell\leq k$ and a subsequence $\{(X_{i_{j}},Y_{i_{j}})\}_{j\in\mathbb{N}}$ satisfying $\operatorname{rank}(X_{i_{j}})={\underline{r}}$ and $\operatorname{rank}(Y_{i_{j}})=k-\bar{r}$ for every $j\in\mathbb{N}$, and the associated subspace sequences (7.15) as constructed in Lemma 3.
Subsequently, we concentrate on the sequence $\{(X_{i_{j}},Y_{i_{j}})\}_{j\in\mathbb{N}}$ and re-assign the index as $i\in\mathbb{N}$, e.g., $\{(X_{i},Y_{i})\}_{i\in\mathbb{N}}$ for simplicity. Inheriting the notation in the statement of Lemma 3 and following the expression (7.10), we have
$\hat{\upsilon}_{i}$ $=\left[U_{i}^{+}\,\,\breve{U}_{i}\,\,U_{iY}^{+}\right]\left[\begin{matrix}0&0&B_{i}^{\upsilon}\\ 0&Z_{1i}&Z_{2i}\\ C^{v}_{i}&Z_{3i}&Z_{4i}\end{matrix}\right]\left[V_{i}^{+}\,\,\breve{V}_{i}\,\,V_{iY}^{+}\right]^{\top},$
$\hat{\omega}_{i}$ $=\left[U_{i}^{+}\,\,\breve{U}_{i}\,\,U_{iY}^{+}\right]\left[\begin{matrix}A_{i}^{\omega}&C_{i2}^{\omega}&C_{i1}^{\omega}\\ B_{i2}^{\omega}&0&0\\ B_{i1}^{\omega}&0&\hat{Z}_{i}\\ \end{matrix}\right]\left[V_{i}^{+}\,\,\breve{V}_{i}\,\,V_{iY}^{+}\right]^{\top},$
where $U^{+}_{i}:=[U_{i}\ \tilde{U}_{i\bot}]$, $U^{+}_{iY}:=[\tilde{U}_{iY\bot}\ U_{iY}]$, $V^{+}_{i}:=[V_{i}\ \tilde{V}_{i\bot}]$, $V^{+}_{iY}:=[\tilde{V}_{iY\bot}\ V_{iY}]$; $Z_{1i}=0$ if ${\underline{r}}<r$ otherwise $Z_{1i}$ is not restricted; and $\hat{Z}_{i}=0$ if $\bar{r}>r$ otherwise $\hat{Z}_{i}$ is not restricted. Moreover, we note that
$B^{\upsilon}_{i}=(U^{+}_{i})^{\top}\hat{\upsilon}_{i}V^{+}_{iY},\ C^{\upsilon}_{i}=(U^{+}_{iY})^{\top}\hat{v}_{i}V_{i}^{+},\ Z_{1i}=\breve{U}_{i}^{\top}\hat{\upsilon}_{i}\breve{V}_{i},$
$Z_{2i}=\breve{U}_{i}^{\top}\hat{\upsilon}_{i}{V}_{iY}^{+},\ Z_{3i}=({U}_{iY}^{+})^{\top}\hat{\upsilon}_{i}\breve{V}_{i},\ Z_{4i}=({U}_{iY}^{+})^{\top}\hat{\upsilon}_{i}{V}_{iY}^{+},$
$A_{i}^{\omega}=(U_{i}^{+})^{\top}\hat{\omega}_{i}V_{i}^{+},\ B_{i1}^{\omega}=(U_{iY}^{+})^{\top}\hat{\omega}_{i}V_{i}^{+},\ B_{i2}^{\omega}=(\breve{U}_{i})^{\top}\hat{\omega}_{i}V_{i}^{+},$
$C_{i1}^{\omega}=(U_{i}^{+})^{\top}\hat{\omega}_{i}V_{iY}^{+},\ C_{i2}^{\omega}=(U_{i}^{+})^{\top}\hat{\omega}_{i}\breve{V}_{i},\ \hat{Z}_{i}=({U}_{iY}^{+})^{\top}\hat{\omega}_{i}{V}_{iY}^{+}.$
All the quantities listed above converge (taking a subsequence if necessary), and thus, letting $i\to+\infty$ leads to the following equalities,
$\upsilon=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}0&0&B^{\upsilon}\\ 0&Z_{1}&Z_{2}\\ C^{v}&Z_{3}&Z_{4}\end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$ (7.23)
$\omega=\left[U^{+}\,\,\breve{U}_{\bot}\,\,U_{Y}^{+}\right]\left[\begin{matrix}A^{\omega}&C_{2}^{\omega}&C_{1}^{\omega}\\ B_{2}^{\omega}&0&0\\ B_{1}^{\omega}&0&\hat{Z}\\ \end{matrix}\right]\left[V^{+}\,\,\breve{V}_{\bot}\,\,V_{Y}^{+}\right]^{\top},$
Next, we examine the relation between the blocks $B_{1}^{\omega}$ and $B^{\upsilon}$. Through (7.13), we have $\varSigma_{R_{i}}B^{\omega}_{i1}=(B_{i}^{\upsilon})^{\top}\varSigma_{i}$, which, by Lemma 2, is equivalent to
${\varLambda}_{i}\odot B_{i1}^{\omega}+\left({\varLambda}_{i}-{\bm{1}}_{\left(k-\ell\right)\times{\underline{r}}}\right)\odot\left(B_{i}^{\upsilon}\right)^{\top}=0,$ (7.24)
where $\varLambda_{i}=\mathfrak{D}(\operatorname{diag}(\varSigma_{R_{i}}),\operatorname{diag}(\varSigma_{i}))$. Notice that the last $(\ell-\bar{r})$ diagonal entries of $\varSigma_{R_{i}}$ vanish in the limit, and similarly, the last $({\underline{r}}-s)$ diagonal entries of $\varSigma_{i}$ vanish. Accordingly, we partition the matrices into four blocks $B_{1}^{\omega}=\left[\begin{matrix}G^{\omega}&J^{\omega}\\ E^{\omega}&F^{\omega}\\ \end{matrix}\right]\ \text{and}\ B^{\upsilon}=\left[\begin{matrix}G^{\upsilon}&J^{\upsilon}\\ E^{\upsilon}&F^{\upsilon}\\ \end{matrix}\right]$, and then take $i\to+\infty$ in (7.24) to obtain
$\left[\begin{matrix}\varXi&{\bm{1}}_{(k-\ell)\times({\underline{r}}-s)}\\ {\bm{0}}_{(\ell-\bar{r})\times s}&D\\ \end{matrix}\right]\odot\left[\begin{matrix}G^{\omega}&J^{\omega}\\ E^{\omega}&F^{\omega}\\ \end{matrix}\right]+\left[\begin{matrix}\varXi-{\bm{1}}&{\bm{0}}_{(k-\ell)\times({\underline{r}}-s)}\\ -{\bm{1}}_{(\ell-\bar{r})\times s}&D-{\bm{1}}\\ \end{matrix}\right]\odot\left[\begin{matrix}G^{\upsilon}&J^{\upsilon}\\ E^{\upsilon}&F^{\upsilon}\\ \end{matrix}\right]^{\top}\!\!\!=0,$
where $\varXi=\mathfrak{D}(\operatorname{diag}(\varSigma_{Y}),\operatorname{diag}(\varSigma))$, and $D$ is an element in $\mathrm{\Theta}(\ell-\bar{r},{\underline{r}}-s)$. Therefore, it is concluded that $J^{\omega}=0$, $J^{\upsilon}=0$, $\varSigma_{Y}G^{\omega}=(G^{\upsilon})^{\top}\varSigma$, and $(F^{\omega},F^{\upsilon})$ satisfies $D\odot F^{\omega}+(D-{\bm{1}}_{(\ell-\bar{r})\times({\underline{r}}-s)})\odot(F^{\upsilon})^{\top}=0$. A parallel analysis for the pair $(C_{1}^{\omega},C^{\upsilon})$ in (7.23) can be conducted, thereby completing the proof.
Recalling the definition from mordukhovich2006variationalI, the (Mordukhovich) coderivative to $\mathrm{N}_{\mathcal{M}_{\leq r}}(\cdot)$ at $(X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}$ is a set-valued mapping $\mathrm{D}^{*}\mathrm{N}_{\mathcal{M}_{\leq r}}(X,Y):\mathbb{R}^{m\times n}\rightrightarrows\mathbb{R}^{m\times n}$ given as follows,
$\mathrm{D}^{*}\mathrm{N}_{\mathcal{M}_{\leq r}}(X,Y)[\omega^{*}]=\left\{{\upsilon^{*}\in\mathbb{R}^{m\times n}\mid(\upsilon^{*},-\omega^{*})\in\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)}\right\},$
for all $\omega^{*}\in\mathbb{R}^{m\times n}$. Consequently, the explicit formula of $\mathrm{N}_{\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}}(X,Y)$ identified in Theorem 7.2 allows for the direct computation of the coderivative.
8 Bilevel programming problems with low-rank structure
Bilevel optimization, in which upper- and lower-level problems are nested with each other, has witnessed various applications yang2025sobiRL; yang2025lancbio and theoretical developments lin2014solving. When the lower-level problem possesses specific structures, e.g., the semidefinite constraints, a more tailored treatment is required ding2014SDCMPCC; wu2014SDCMPCC; dempe2018optimality.
In this section, we consider the bilevel programming problem (LRBP), where the lower level seeks a solution constrained to the set of bounded-rank matrices. Our goal is to derive an optimality condition, as a direct application of the results developed in Section 7.
8.1 Motivating applications
We now present two representative applications falling into the scope of the formulation (LRBP).
Bilevel optimization with low-rank adaptation.
Natural language processing has increasingly adopted bilevel optimization to address various tasks grangier2023LLMshift; shen2025seal; zangrando2025debora. Specifically, the upper level introduces a task-oriented variable $x\in\mathbb{R}^{q}$ while the lower level trains a large language model (LLM), which resorts to the popular parameter-efficient fine-tuning approach hu2022lora, i.e., freezing the pretrained model weight $\bar{X}\in\mathbb{R}^{m\times n}$ and optimizing an additive low-rank trainable matrix $X\in\mathcal{M}_{\leq r}$. Therefore, the following bilevel formulation summarizes the discussed applications,
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{q},X^{*}\in\mathbb{R}^{m\times n}}&\ \frac{1}{\left|\mathcal{D}_{1}\right|}\sum_{y_{i}\in\mathcal{D}_{1}}\mathcal{L}(x,\bar{X}+X^{*};y_{i})\\[8.53581pt] \mathrm{s.\,t.}\!\!\!\!&\ G(x)\leq 0,\\ &\ \begin{aligned} X^{*}\in\operatorname*{arg\,min}_{X\in\mathbb{R}^{m\times n}}&\ \ \frac{1}{\left|\mathcal{D}_{2}\right|}\sum_{y_{i}\in\mathcal{D}_{2}}\mathcal{L}(x,\bar{X}+X;y_{i}),\\ \mathrm{s.\,t.}&\ \ X\in\mathcal{M}_{\leq r},\end{aligned}\end{array}$
where $\mathcal{L}$ denotes the loss function, $\mathcal{D}_{j}$ ($j=1,2$) are different datasets, and $\{y_{i}\}$ label the associated data points.
Data hyper-cleaning with low-rank model.
A line of applications in machine learning community only receives corrupted or noisy data while is required to train a reliable model. To this end, the approach, data hyper-cleaning shaban2019truncated formulates the task as a bilevel problem,
$\begin{array}[]{cl}\min\limits_{w\in\mathbb{R}^{q},X^{*}\in\mathbb{R}^{m\times n}}&\ \frac{1}{\left|\mathcal{D}_{1}\right|}\sum_{y_{i}\in\mathcal{D}_{1}}\mathcal{L}(w,X^{*};y_{i})\\[5.69054pt] \mathrm{s.\,t.}&\ \begin{aligned} X^{*}&\in\operatorname*{arg\,min}_{X\in\mathbb{R}^{m\times n}}\ \frac{1}{\left|\mathcal{D}_{2}\right|}\sum_{y_{i}\in\mathcal{D}_{2}}c(w_{i})\mathcal{L}(w,X;y_{i}),\end{aligned}\end{array}$
where the upper level searches for a weight $w$ deciding the confidence of each data through a mapping $c:\mathbb{R}^{q}\to\mathbb{R}_{+}$, and the lower level trains a model according to the weighted data. When the lower-level model $X\in\mathbb{R}^{m\times n}$ possesses a low-rank structure—typical examples including image recovery zhang2013hyperspectral; wang2017reweighted and network training idelbayev2020lowrankcompress; yaras2024compressible—it is advantageous to impose the constraint $\operatorname{rank}(X)\leq r$, which will effectively circumvent parameter redundancy while preserving a decent performance.
8.2 Optimality conditions via a relaxation
Note that finding a global minimizer of a function subject to the bounded-rank constraint is NP-hard in general gillis2011NPlowrank. Nevertheless, existing literature schneider2015Lojaconvergence; levin2023remedy; olikier2023RFDR is able to find a first-order point in the sense that the antigradient belongs to the Mordukhovich normal cone of the determinantal variety. Therefore, we turn to the formulation (1.3), which serves as a relaxation for (LRBP) by replacing the lower-level global optimality with the first-order stationarity. Subsequently, introducing a slack variable $Y\in\mathbb{R}^{m\times n}$, we obtain the following formulation equivalent to (1.3),
$\min_{x\in\mathbb{R}^{q},X\in\mathbb{R}^{m\times n}}$ $\ \mathcal{L}(x,X)$ (M-LRBP)
$\mathrm{s.\,t.}$ $\ G(x)\leq 0,$
$\ \nabla_{X}F(x,X)+Y=0,$
$\ (X,Y)\in\mathrm{gph}\mathrm{N}_{\mathcal{M}_{\leq r}}.$
Since the relaxation is based on the Mordukhovich normal cone, we prefix the name of (LRBP) with an additional “M-”. Moreover, we assume that the mappings $\mathcal{L}:\mathbb{R}^{q}\times\mathbb{R}^{m\times n}\to\mathbb{R}$ and $G:\mathbb{R}^{q}\to\mathbb{R}^{p}$ are continuously differentiable, while $F:\mathbb{R}^{q}\times\mathbb{R}^{m\times n}\to\mathbb{R}$ is twice continuously differentiable.
We then investigate the relationship between (LRBP) and (M-LRBP) in terms of local optimal solutions, namely, points that minimize the objective $\mathcal{L}$ over a neighborhood in the feasible region.
Proposition 10
[…]
Proof
[…]
Finally, taking into account the coderivative of the Mordukhovich normal cone mapping developed in section 7, we can give a Fritz John type M-stationary condition for (M-LRBP), where the notation follows from Theorem 7.2, e.g., $k:=\min\{m,n\}$ and $0\leq s\leq r\leq\ell\leq k$, and we denote by $\mathcal{J}_{X}$ the partial Jacobian of a mapping with respect to $X$.
Proposition 11
[…]
$\mu\nabla_{x}\mathcal{L}(\tilde{x},\tilde{X})+\nabla G(\tilde{x})\lambda+\mathcal{J}_{X}(\nabla_{x}F)(\tilde{x},\tilde{X})[\delta]=0,$ (8.1)
$\mu\nabla_{X}\mathcal{L}(\tilde{x},\tilde{X})+\mathcal{J}_{X}(\nabla_{X}F)(\tilde{x},\tilde{X})[\delta]+\omega=0,$ (8.2)
$\delta+\upsilon=0,$ (8.3)
$\langle{G(\tilde{x}),\lambda}\rangle_{\mathbb{R}^{p}}=0,\ \ \lambda\geq 0,$ (8.4)
[…]
Proof
[…]
9 Conclusions and perspectives
In this paper, we conduct the variational analysis of determinantal varieties. Specifically, we provide a unified framework for analyzing first- and second-order tangent sets to various low-rank set, recovering existing results, and revealing a range of new ones. Drawing on the tangent sets, we establish a sufficient and necessary condition to characterize the second-order equivalence between a general nonsmooth problem and its smooth parameterization. The developed framework is applied to low-rank optimization. In another thread of analysis, we investigate the geometry of the graph of the Mordukhovich normal cone to the matrix variety, which plays a role in low-rank bilevel programs. We conclude with several remarks and outline potential directions for future research inspired by this work.
Extension to sparsity constraints. In fact, Theorem 3.1 also finds potential applications in sparse scenarios. Specifically, consider the sparse set $\mathcal{C}_{s}:=\{x\in\mathbb{R}^{q}\mid\|x\|_{0}\leq s\}$, where $\|\cdot\|_{0}$ counts the number of nonzero entries of a vector. Let $|x|^{\downarrow}$ denote the vector obtained by sorting $(|x_{1}|,\dots,|x_{q}|)$ in a non-increasing order. Then, analogous to (3.1), the sparse set admits the characterization $\mathcal{C}_{s}=\{x\in\mathbb{R}^{q}\mid(|x|^{\downarrow})_{s+1}=0\}$. It can be verified that $\mathcal{C}_{s}$ satisfies the error bound condition: $\operatorname{dist}(y,\mathcal{C}_{s})\leq\sqrt{q-s}\,(|y|^{\downarrow})_{s+1}$ for any $y\in\mathbb{R}^{q}$. Additionally, we note that $(|x|^{\downarrow})_{s+1}=\lambda_{s+1}(\operatorname{Diag}(|x|))$, where both the mappings $x\mapsto\operatorname{Diag}(|x|)$ and $\lambda_{s+1}(\cdot)$ are locally Lipschitz and admit first- and second-order directional derivatives; hence, the composite mapping $x\mapsto(|x|^{\downarrow})_{s+1}$ inherits the same properties. Therefore, Theorem 3.1 can be invoked to derive the first- and second-order tangent sets to $\mathcal{C}_{s}$. Extending the spirit, Theorem 4.1 may further be employed to develop the intersection rules, when an additional constraint is imposed to the sparse set.
Low-rank sets intersecting with inequality constraints. An immediate extension of Theorem 4.1 involves cases where $\mathcal{K}$ is defined by a system including inequalities, i.e., $\mathcal{K}=\{X\in\mathbb{R}^{q}\mid h(X)=0,\;g(X)\leq 0\}$, mirroring the setup in Theorem 3.1. In such a scenario, the parameterization $(\overline{\mathcal{M}},\phi)$ may pull $\mathcal{K}$ back to a manifold with boundary lee2012manifolds. An appropriate generalization of Theorem 4.1 holds promise for applications to the intersection of $\mathcal{M}_{\leq r}$ with constraints such as the closed unit Frobenius ball, the symmetric box, or the spectrahedron, as discussed in li2020jotaspectral.
Algorithms for low-rank bilevel programming problems. While Proposition 11 gives an optimality condition for problem (M-LRBP), designing an algorithm provable to accumulate at such stationary points remains a challenge. We envision that progress can be made by initially restricting attention to scenarios where the lower-level low-rank problem minimizes a strongly convex objective park2018findingefficiently, a direction inspired by the avenue developed in the existing bilevel optimization literature ghadimi2018approximation.
Appendix A Tangent sets to tensor varieties
In this section, we introduce the hierarchical Tucker (HT) variety hackbusch2009newHT; grasedyck2010hierarchicalSVD, and then show that the Tucker and tensor train (TT) varieties tucker1964extension; oseledets2011TTSVD arise as specific cases of the HT variety. Finally, a proof for Proposition 2 is provided.
A.1 Hierarchical Tucker varieties
Given the order $d$, a dimension tree $T$ on $\{1,2,\ldots,d\}$ is a binary tree whose nodes are nonempty subsets $t\subseteq\{1,\ldots,d\}$ such that: 1) the root is $t_{r}=\{1,\ldots,d\}$; 2) the leaves are the singletons $\{k\}$, $k=1,2,\ldots,d$; 3) if $t$ is an internal node with children $t_{1}$ and $t_{2}$, then $t=t_{1}\cup t_{2}$ and $k_{1}<k_{2}$ for all $k_{1}\in t_{1}$ and $k_{2}\in t_{2}$. For each subset $t\subseteq\{1,2,\ldots,d\}$, we denote the associated dimension by $n_{t}:=\prod_{k\in t}n_{k}$, and the set complementary to $t$ by $\tt_{-}=\{1,2,\ldots,d\}\backslash t$.
Based on a fixed dimension tree $T$, the HT mode-$t$ unfolding (or matricization) of $\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ is formed by arranging the modes in $t=\{\mu_{1},\ldots,\mu_{p}\}$ along the row dimension and those in $t_{-}=\{v_{1},\ldots,v_{d-p}\}$ along the column dimension:
$X^{\mathrm{ht}}_{(t)}\in\mathbb{R}^{n_{t}\times n_{t_{-}}}\ \ \text{with}\ \ X^{\mathrm{ht}}_{(t)}(i_{\mu_{1}},i_{\mu_{2}},\ldots,i_{\mu_{p}};i_{\nu_{1}},i_{\nu_{2}},\ldots,i_{\nu_{d-p}})=\mathbf{X}_{i_{1},\ldots,i_{d}},$
Each matricization operator is invertible, as it is a one-to-one rearrangement of the tensor entries. Hence the mapping $\mathbf{X}\mapsto X^{{\mathrm{ht}}}_{(t)}$ admits a unique inverse $\mathrm{ten}^{\mathrm{ht}}_{(t)}:\mathbb{R}^{n_{t}\times n_{t_{-}}}\to\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$, called the mode-$t$ tensorization.
The HT rank of a tensor $\mathbf{X}$ is defined as the tuple
$\mathrm{rank}_{\mathrm{ht}}(\mathbf{X})=\left(r_{t}\right)_{t\in T}\ \text{ with }\ r_{t}=\operatorname{rank}(X^{\mathrm{ht}}_{(t)})\ \text{for}\ t\in T.$
Consequently, we can introduce the set of hierarchical tensors with an HT rank at most $\mathbf{r}\in\mathbb{N}^{|T|}$:
$\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}=\{\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\mathrm{rank}_{\mathrm{ht}}(\mathbf{X})\leq\mathbf{r}\}.$
Note that $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ is the common zero set of all $(r_{t}+1)$-minors of the mode-$t$ unfolding matrices $\{X^{{\mathrm{ht}}}_{(t)}\}_{t\in T}$, i.e., $\operatorname{rank}(X^{\mathrm{ht}}_{(t)})\leq r_{t}$ for all $t\in T$. Hence $\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}$ is a real algebraic variety; we refer to it as the HT variety.
A.2 Reduction to Tucker varieties and tenser train varieties
Different choices of the dimension tree $T$ with the associated rank parameter $\mathbf{r}=(r_{t})_{t\in T}$ will yield varieties with different structures. We then specify the constructions of TT and $(r_{t})_{t\in T}$ to reduce the HT variety to two typical instances—the Tucker variety tucker1964extension; gao2025lowranktucker and the tenser train variety oseledets2011TTSVD; kutschan2018tangentTT, respectively.
To derive the concept of the Tucker variety, let $T$ be the dimension tree on $\{1,2,\ldots,d\}$ that, at each level, divides the first spatial index to the left child and assigns the rest to the right child. Formally, set $t_{r}=\{1,2,\ldots,d\}$, and recursively, for the internal node $t=\{k,\ldots,d\}$ with $2\leq k<d$, define its two children by $t_{1}=\{k\}$ and $t_{2}=\{k+1,\ldots,d\}$. Moreover, letting $r_{\{k\}}$ denote the rank of the mode-$\{k\}$ unfolding of a given $\mathbf{X}$ for $k=1,2,\ldots,d$, we can define the mapping $\operatorname{rank}_{\mathrm{tc}}:\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\to\mathbb{N}^{d}:\,\mathbf{X}\mapsto(r_{\{1\}},r_{\{2\}},\ldots,r_{\{d\}})$. Consequently, the Tucker variety can be defined as
$\mathcal{M}^{\mathrm{tc}}_{\leq\mathbf{r}}=\{\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\mathrm{rank}_{\mathrm{tc}}(\mathbf{X})\leq\mathbf{r}^{\mathrm{tc}}\},$ (A.1)
where $\mathbf{r}^{\mathrm{tc}}=(r^{\mathrm{tc}}_{1},r^{\mathrm{tc}}_{2},\ldots,r^{\mathrm{tc}}_{d})\in\mathbb{N}^{d}$ is a given Tucker rank parameter such that $r^{\mathrm{tc}}_{k}\leq\min\{n_{\{k\}},n_{\{k\}_{-}}\}$ for $k\in\{1,2,\ldots,d\}$.
In parallel with the derivation of (A.1), we can also view the TT variety as a special case of the HT variety, which has been discussed in (uschmajew2013geometryofHT, §5); for completeness, we briefly recall the construction. We adopt the same partition of the tree $T$ as in the Tucker case, namely, setting the root $t_{r}=\{1,2,\ldots,d\}$, and recursively, for the internal node $t=\{k,\ldots,d\}$ with $2\leq k<d$, defining the children by $t_{1}=\{k\}$ and $t_{2}=\{k+1,\ldots,d\}$. Then, letting $r_{\{k,\ldots,d\}}$ denote the rank of the mode-$\{k,\ldots,d\}$ unfolding of a given $\mathbf{X}$ for $k=2,\ldots,d$, we can define the mapping $\operatorname{rank}_{\mathrm{tt}}:\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\to\mathbb{N}^{d-1}:\,\mathbf{X}\mapsto(r_{\{2,\ldots,d\}},r_{\{3,\ldots,d\}},\ldots,r_{\{d\}})$, which introduces the definition of the TT variety as follows,
$\mathcal{M}^{\mathrm{tt}}_{\leq\mathbf{r}}=\{\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\mathrm{rank}_{\mathrm{tt}}(\mathbf{X})\leq\mathbf{r}^{\mathrm{tt}}\},$ (A.2)
where $\mathbf{r}^{\mathrm{tt}}=(r^{\mathrm{tt}}_{1},r^{\mathrm{tt}}_{2},\ldots,r^{\mathrm{tt}}_{d-1})\in\mathbb{N}^{d-1}$ is a given TT rank parameter such that $r^{\mathrm{tt}}_{k}\leq\min\{n_{\{1,\ldots,k\}},n_{\{1,\ldots,k\}_{-}}\}$ for $k\in\{1,2,\ldots,d-1\}$.
A.3 Proof of Proposition 2
Proof
[…]
$\operatorname{dist}(\mathbf{Y},\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}})\leq\|\mathbf{Y}-\mathcal{P}_{\leq\mathbf{r}}(\mathbf{Y})\|_{\mathrm{F}}\leq\sqrt{\sum_{t\in T}\sum_{i>r_{t}}\sigma^{2}_{i}(Y^{{\mathrm{ht}}}_{(t)})}\leq\rho\sqrt{\sum_{t\in T}\sigma^{2}_{r_{t}+1}(Y^{{\mathrm{ht}}}_{(t)})}\,$,
where $\rho=\sqrt{\max_{t\in T}\{\min\{n_{t},n_{t_{-}}\}-r_{t}\}}$. Therefore, we can apply Theorem 3.1 to $\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}$ given through (4.3).
Regarding the tangent cone, we have
$\mathrm{T}_{\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}}(\mathbf{X})$ $=\left\{{\bm{\eta}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\sigma^{\prime}_{r_{t}+1}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)})=0\ \text{for}\ t\in T}\right\}$
$=\bigcap_{t\in T}\left\{{\bm{\eta}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\sigma^{\prime}_{r_{t}+1}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)})=0}\right\}$
$=\bigcap_{t\in T}{\mathrm{ten}_{(t)}^{\mathrm{ht}}\left({\mathrm{T}_{\mathcal{R}_{t}}(X^{\mathrm{ht}}_{(t)})}\right)}$
where $\mathcal{R}_{t}:=\mathbb{R}_{\leq r_{t}}^{n_{t}\times n_{t_{-}}}$ and the last equality comes from (3.5). Similarly, let $\bm{\eta}\in\mathrm{T}_{\mathcal{M}_{\leq\mathbf{r}}^{\mathrm{ht}}}(\mathbf{X})$, and it holds that
$\mathrm{T}^{2}_{\mathcal{M}^{\mathrm{ht}}_{\leq\mathbf{r}}}(\mathbf{X};\bm{\eta})$ $=\left\{{\bm{\zeta}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\sigma^{\prime\prime}_{r_{t}+1}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)},\zeta^{\mathrm{ht}}_{(t)})=0\ \text{for}\ t\in T}\right\}$
$=\bigcap_{t\in T}\left\{{\bm{\zeta}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}\mid\sigma^{\prime\prime}_{r_{t}+1}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)},\zeta^{\mathrm{ht}}_{(t)})=0}\right\}$
$=\bigcap_{t\in T}{\mathrm{ten}_{(t)}^{\mathrm{ht}}\left({\mathrm{T}^{2}_{\mathcal{R}_{t}}(X^{\mathrm{ht}}_{(t)};\eta^{\mathrm{ht}}_{(t)})}\right)}.$
Appendix B Proof of Theorem 4.1
In this section, we provide the proof for Theorem 4.1. The main idea is to show that if $(\mathcal{M},\mathcal{K})$ satisfies Assumption 2, their intersection $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1, thereby applying the developed Theorem 3.1 gives the conclusion.
As a preliminary, we show that any set $\mathcal{K}$ realized as a level set of a smooth mapping with constant rank satisfies the error bound property, which is closely related to Assumption 2(i).
Lemma 4
Let $h:\mathbb{R}^{q}\to\mathbb{R}^{k}$ be smooth and let $\mathcal{K}=\{\tilde{X}\in\mathbb{R}^{q}\mid h(\tilde{X})=0\}$. Given $X\in\mathcal{K}$, suppose that there exists a neighborhood $\mathcal{B}_{1}\subseteq\mathbb{R}^{q}$ around $X$ where the Jacobian $\mathrm{D}h$ has constant rank $r$. Then it admits a neighborhood $\mathcal{B}_{2}\subseteq\mathbb{R}^{q}$ around $X$ and a constant $\rho>0$ such that $\operatorname{dist}(Y,\mathcal{K})\leq\rho\|h(Y)\|_{2}$ for all $Y\in\mathcal{B}_{2}$.
Proof
By the constant rank theorem (lee2012manifolds, Theorem 4.12), there exist open neighborhoods $\mathcal{B}\subseteq\mathbb{R}^{q}$ of $X$ and $\mathcal{W}\subseteq\mathbb{R}^{k}$ of $h(X)=0\in\mathbb{R}^{k}$, a diffeomorphism $\phi:\mathcal{B}\to\mathcal{B}^{\prime}\subseteq\mathbb{R}^{q}$ with $\phi(X)=0$, and a diffeomorphism $\psi:\mathcal{W}\to\mathcal{W}^{\prime}\subseteq\mathbb{R}^{k}$ with $\psi(0)=0$, such that in the coordinates $(u,v)\in\mathbb{R}^{r}\times\mathbb{R}^{q-r}$ one has
$(\psi\circ h\circ\phi^{-1})(u,v)=(u,0)\in\mathbb{R}^{r}\times\mathbb{R}^{\,k-r}.$
In these coordinates, we have $\mathcal{K}\cap\mathcal{B}=\{\tilde{X}\in\mathcal{B}\mid\phi(\tilde{X})=(0,v)\ \text{for some}\ v\in\mathbb{R}^{q-r}\}$.
Shrink $\mathcal{B}$ if necessary such that $\phi$ and $\phi^{-1}$ are Lipschitz on $\mathcal{B}$ and $\mathcal{B}^{\prime}$ with constants $L_{\phi}$ and $L_{\phi^{-1}}$, respectively; and similarly, shrink $\mathcal{W}$ such that $\psi$ is Lipschitz on $\mathcal{W}$ with a constant $L_{\psi}$. Moreover, take the neighborhood $\mathcal{B}_{2}\subseteq\mathcal{B}$ and $h(\mathcal{B}_{2})\subseteq\mathcal{W}$. Given any $Y\in\mathcal{B}_{2}$, let $\phi(Y)=(u,v)$ and $Y_{p}\in\mathcal{K}$ be the point with the coordinate $\phi(Y_{p})=(0,v)$, which leads to
$\operatorname{dist}(Y,\mathcal{K})\leq\|Y-Y_{p}\|_{2}\leq L_{\phi^{-1}}\|(u,v)-(0,v)\|_{2}=L_{\phi^{-1}}\|u\|_{2}.$ (B.1)
On the other hand, it holds that $\|u\|_{2}=\|(u,0)\|_{2}=\|\psi\circ h\circ\phi^{-1}(u,v)\|_{2}=\|\psi(h(Y))-\psi(0)\|_{2}\leq L_{\psi}\|h(Y)\|_{2}$. Combining it with (B.1) gives $\operatorname{dist}(Y,\mathcal{K})\leq L_{\phi^{-1}}L_{\psi}\|h(Y)\|_{2}$. Setting $\rho:=L_{\phi^{-1}}L_{\psi}$ completes the proof.
Subsequently, we show that when two manifolds intersect transversally, the distance to their intersection can be bounded by the distances to each manifold. This property resonates with Assumption 2(ii).
Lemma 5
Let $\mathcal{K}_{1},\mathcal{K}_{2}\subseteq\mathbb{R}^{q}$ be smooth embedded manifolds, and let $X\in\mathcal{K}_{1}\cap\mathcal{K}_{2}$. Suppose that $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ intersect transversally at $X$, i.e., $\mathrm{T}_{\mathcal{K}_{1}}(X)+\mathrm{T}_{\mathcal{K}_{2}}(X)=\mathbb{R}^{q}$. Then there exist a neighborhood $\mathcal{B}\subseteq\mathbb{R}^{q}$ of $X$ and a constant $C>0$ such that
$\operatorname{dist}(Y,\mathcal{K}_{1}\cap\mathcal{K}_{2})\ \leq\ C\Bigl(\operatorname{dist}(Y,\mathcal{K}_{1})+\operatorname{dist}(Y,\mathcal{K}_{2})\Bigr)\ \ \text{for all}\ Y\in\mathcal{B}.$
Proof
Since $\mathcal{K}_{i}$ are embedded, according to (boumal2023introduction, Theorem 8.75), there exist neighborhoods $\mathcal{B}_{i}$ of $X$ and smooth $h_{i}:\mathcal{B}_{i}\to\mathbb{R}^{k_{i}}$ with full-rank Jacobians at $X$, such that $\mathcal{K}_{i}\cap\mathcal{B}_{i}=\{\,\tilde{X}\in\mathcal{B}_{i}\mid\ h_{i}(\tilde{X})=0\}$ for $i=1,2$.
Let $\mathcal{B}:=\mathcal{B}_{1}\cap\mathcal{B}_{2}$ and $)h:\mathcal{B}\to\mathbb{R}^{k_{1}}\times\mathbb{R}^{k_{2}}:\,\tilde{X}\mapsto(h_{1}(\tilde{X}),h_{2}(\tilde{X}$)). Transversality at $X$ is equivalent to $\operatorname{rank}(\mathrm{D}h_{X})=k_{1}+k_{2}$. By continuity of $\mathrm{D}h$, shrinking $\mathcal{B}$ if necessary, we may assume $\operatorname{rank}(\mathrm{D}h_{X})\equiv k_{1}+k_{2}$ on $\mathcal{B}$. Apply Lemma 4 on the manifold $\{\tilde{X}\in\mathcal{B}\mid h(\tilde{X})=0\}$ to give a $\rho_{h}>0$ such that
$\operatorname{dist}(Y,\mathcal{K}_{1}\cap\mathcal{K}_{2})\leq\rho_{h}\|h(Y)\|_{2}\ \ \text{for all}\ Y\in\mathcal{B}.$ (B.2)
After possibly shrinking the neighborhood $\mathcal{B}$, there exist $L_{i}>0$ such that $h_{i}$ are $L_{i}$-Lipschitz on $\mathcal{B}$ ($i=1,2$).
Consequently, give any $Y\in\mathcal{B}$, we can pick $\mathcal{P}_{i}(Y)\in\mathcal{K}_{i}$ with $\|Y-\mathcal{P}_{i}(Y)\|_{2}=\operatorname{dist}(Y,\mathcal{K}_{i})$, and thus we have $\|h_{i}(Y)\|_{2}=\|h_{i}(Y)-h_{i}(\mathcal{P}_{i}(Y))\|_{2}\leq L_{i}\operatorname{dist}(Y,\mathcal{K}_{i})$ for $i=1,2$. Additionally, incorporating $\|h(Y)\|_{2}\leq\|h_{1}(Y)\|_{2}+\|h_{2}(Y)\|_{2}$ into (B.2) yields $\operatorname{dist}(Y,\mathcal{K}_{1}\cap\mathcal{K}_{2})\leq\rho_{h}\!\left(L_{1}\,\operatorname{dist}(Y,\mathcal{K}_{1})+L_{2}\,\operatorname{dist}(Y,\mathcal{K}_{2})\right)$, and taking $C:=\rho_{h}\max\{L_{1},L_{2}\}$ arrives at the conclusion.
We are now in a position to prove Theorem 4.1. The proof proceeds by combining Lemma 4 and Lemma 5 to establish the error bound property of $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ in the auxiliary space $\mathbb{R}^{\bar{q}}$, and then transferring this property to $\mathcal{M}\cap\mathcal{K}$ via the smooth mapping $\phi:\mathbb{R}^{\bar{q}}\to\mathbb{R}^{q}$. As a result, $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1, and thus falls within the scope of Theorem 3.1.
Proof of Theorem 4.1 Since $\mathcal{M}$ satisfies Assumption 1 and $h$ is smooth, we can take a neighborhood $\mathcal{B}_{p}\subseteq\mathbb{R}^{q}$ of $X$ such that $c_{1}$ and $h$ are $L_{c}$- and $L_{h}$-Lipschitz on $\mathcal{B}_{p}$, respectively, and $\operatorname{dist}({Y},\mathcal{M})\leq\rho_{\mathcal{M}}\|c_{1}({Y})\|_{2}$ holds for a constant $\rho_{\mathcal{M}}>0$ and any ${Y}\in\mathcal{B}_{p}$ (see Lemma 4). Therefore, we can find a neighborhood $\mathcal{B}\subseteq\mathcal{B}_{p}$ such that given any $Y\in\mathcal{B}$, there exists $Y_{p}\in\mathcal{M}\cap\mathcal{B}_{p}$ such that $\|Y-Y_{p}\|_{2}=\operatorname{dist}(Y,\mathcal{M})$.
Since $\phi_{\overline{\mathcal{M}}}:\overline{\mathcal{M}}\to\mathcal{M}$ is smooth and open at $x\in\phi^{-1}(X)$, we can shrink $(\mathcal{B}_{p},\mathcal{B})$ until it admits a neighborhood $\overline{\mathcal{B}}_{p}\subseteq\mathbb{R}^{\bar{q}}$ of $x$ such that $\mathcal{M}\cap\mathcal{B}_{p}\subseteq\phi(\overline{\mathcal{M}}\cap\overline{\mathcal{B}}_{p})$ and $\phi$ is $L_{\phi}$-Lipschitz on $\overline{\mathcal{B}}_{p}$. Therefore, we can find a preimage of $Y_{p}$, denoted by $y_{p}\in\overline{\mathcal{M}}\cap\overline{\mathcal{B}}_{p}$. By Lemma 5 and shrinking $(\mathcal{B}_{p},\mathcal{B},\overline{\mathcal{B}}_{p})$ if necessary, we can find a $z_{p}\in(\overline{\mathcal{M}}\cap\overline{\mathcal{K}})\cap\overline{\mathcal{B}}_{p}$ such that $\|y_{p}-z_{p}\|_{2}=\operatorname{dist}(y_{p},\overline{\mathcal{M}}\cap\overline{\mathcal{K}})$, and the transversality of $\overline{\mathcal{M}}\cap\overline{\mathcal{K}}$ reveals that for a constant $C>0$, we have
$\|y_{p}-z_{p}\|_{2}\leq C\left({\operatorname{dist}(y_{p},\overline{\mathcal{M}})+\operatorname{dist}(y_{p},\overline{\mathcal{K}})}\right).$ (B.3)
Note that $\operatorname{dist}(y_{p},\overline{\mathcal{M}})=0$ due to $y_{p}\in\overline{\mathcal{M}}$. Moreover, since $\overline{\mathcal{K}}=\{\tilde{x}\in\mathbb{R}^{\bar{q}}\mid h\circ\phi(\tilde{x})=0\}$ satisfies Assumption 2(i), we apply Lemma 4 on $\overline{\mathcal{K}}$ to obtain $\operatorname{dist}(y_{p},\overline{\mathcal{K}})\leq\rho_{\overline{\mathcal{K}}}\|h(\phi(y_{p}))\|_{2}$ for a constant $\rho_{\mathcal{K}}>0$. Therefore, (B.3) indicates that
$\|y_{p}-z_{p}\|_{2}\leq C\operatorname{dist}(y_{p},\overline{\mathcal{K}})\leq C\rho_{\overline{\mathcal{K}}}\|h(\phi(y_{p}))\|_{2}=C\rho_{\overline{\mathcal{K}}}\|h(Y_{p})\|_{2}$ (B.4)
Consequently, letting $Z_{p}=\phi(z_{p})\in\mathcal{M}\cap\mathcal{K}$, it holds that for all $Y\in\mathcal{B}$,
$\|Y-Z_{p}\|_{2}$ $\leq\|Y-Y_{p}\|_{2}+\|Y_{p}-Z_{p}\|_{2}$
$\leq\rho_{\mathcal{M}}\|c_{1}(Y)\|_{2}+L_{\phi}\|y_{p}-z_{p}\|_{2}$
$\leq\rho_{\mathcal{M}}\|c_{1}(Y)\|_{2}+L_{\phi}C\rho_{\overline{\mathcal{K}}}\|h(Y_{p})\|_{2}$
$\leq\rho_{\mathcal{M}}\|c_{1}(Y)\|_{2}+L_{\phi}C\rho_{\overline{\mathcal{K}}}\left({L_{h}\|Y-Y_{p}\|_{2}+\|h(Y)\|_{2}}\right)$
$\leq\rho_{\mathcal{M}}(1+L_{\phi}L_{h}C\rho_{\overline{\mathcal{K}}})\|c_{1}(Y)\|_{2}+L_{\phi}C\rho_{\overline{\mathcal{K}}}\|h(Y)\|_{2},$
where we employ the the triangle inequality $\|h(Y_{p})\|_{2}-\|h(Y)\|_{2}\leq\|h(Y_{p})-h(Y)\|_{2}$, and the Lipschitz continuity of $\phi$ and $h$. Therefore, setting $\rho=\max\{\rho_{\mathcal{M}}(1+L_{\phi}L_{h}C\rho_{\overline{\mathcal{K}}}),\,L_{\phi}C\rho_{\overline{\mathcal{K}}}\}$ concludes that $\operatorname{dist}(Y,\mathcal{M}\cap\mathcal{K})\leq\rho\|(c_{1}(Y),h(Y))\|_{2}$ for all $Y\in\mathcal{B}$, which together with the locally Lipschitz property of $(c_{1},h)$, verifies that $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1. Applying Theorem 1 to $\mathcal{M}\cap\mathcal{K}=\{\tilde{X}\in\mathbb{R}^{q}\mid c_{1}(\tilde{X})=0,\,h(\tilde{X})=0\}$ gives
$\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)$ $=\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime}(X;\eta)=0,\,h^{\prime}(X;\eta)=0\}$
$=\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime}(X;\eta)=0\}\cap\{\eta\in\mathbb{R}^{q}\mid h^{\prime}(X;\eta)=0\}$
$=\mathrm{T}_{\mathcal{M}}(X)\cap\mathrm{T}_{\mathcal{K}}(X),$
where the last equality holds since both $\mathcal{M}$ and $\mathcal{K}$ satisfy Assumption 1. Similarly, letting $\eta\in\mathrm{T}_{\mathcal{M}\cap\mathcal{K}}(X)$, we have
$\mathrm{T}^{2}_{\mathcal{M}\cap\mathcal{K}}(X;\eta)$ $=\{\zeta\in\mathbb{R}^{q}\mid c_{1}^{\prime\prime}(X;\eta,\zeta)=0,\,h^{\prime\prime}(X;\eta,\zeta)=0\}$
$=\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime\prime}(X;\eta,\zeta)=0\}\cap\{\eta\in\mathbb{R}^{q}\mid h^{\prime\prime}(X;\eta,\zeta)=0\}$
$=\mathrm{T}^{2}_{\mathcal{M}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{K}}(X;\eta).$
Appendix C Tangent sets to $\mathcal{M}_{\leq r}\cap\mathcal{H}$
In this section, which supplements section 4.2.1, we apply Theorem 4.1 to sets of low-rank rectangular matrices in the form of $\mathcal{M}_{\leq r}\cap\mathcal{H}$, recovering the first-order results in li2023normalboundedaffine and yang2025spacedecouple, respectively, and further characterizing the second-order tangent sets. Finally, we consider the case $\mathcal{H}=\mathbb{H}^{n}_{m-1}$ and derive the first- and second-order tangent sets of $\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}$, which represent novel contributions.
The derivation proceeds by checking that $\mathcal{M}_{\leq r}\cap\mathcal{H}$ satisfies Assumption 2, which mainly resorts to the LR parameterization for $\mathcal{M}_{\leq r}$,
$(\mathcal{M}_{\mathrm{LR}},\phi_{\mathrm{LR}})=(\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r},\,(L,R)\mapsto LR^{\top}).$ (C.1)
Then, a straightforward application of Theorem 4.1 leads to the intersection rules
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathcal{H}}(X),$ (C.2)
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X;\eta)$ $=\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{H}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathcal{H}}(X).$
To facilitate the discussion, we note that $\mathcal{M}_{\mathrm{LR}}$ coincides with the whole ambient Euclidean space $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$, and thus the transversality property in Assumption 2(ii) naturally holds. Moreover, $\phi_{\mathrm{LR}}|_{\mathcal{M}_{\mathrm{LR}}}$ is open at $(L,R)\in\mathcal{M}_{\mathrm{LR}}$ if and only if $\operatorname{rank}(L)=\operatorname{rank}(R)=\operatorname{rank}(LR^{\top})$, according to (levin2025effect, Theorem 2.3, Proposition 2.8). Therefore, when $(\mathcal{M}_{\mathrm{LR}},\phi_{\mathrm{LR}})$ is chosen as the smooth parameterization of $\mathcal{M}_{\leq r}$, conditions (ii) and (iii) of Assumption 2 are automatically satisfied. It then remains to verify condition (i) for applying Theorem 4.1 to $\mathcal{M}_{\leq r}\cap\mathcal{H}$. This observation motivates the following corollary, which can be viewed as an instance of Theorem 4.1 specified for $\mathcal{M}_{\leq r}\cap\mathcal{H}$.
Corollary 3
Suppose that $h:\mathbb{R}^{m\times n}\to\mathbb{R}^{q}$ is smooth and $\mathcal{H}=\{\tilde{X}\in\mathbb{R}^{m\times n}\mid h(\tilde{X})=0\}$ satisfies Assumption 1 at point $X\in\mathcal{M}_{\leq r}\cap\mathcal{H}$. Additionally, the differential of the $h\circ\phi_{\mathrm{LR}}$ has constant rank in a neighborhood of
$\overline{\mathcal{H}}:=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid h(\phi_{\mathrm{LR}}(L,R))=0\}=\phi_{\mathrm{LR}}^{-1}(\mathcal{H}).$
Then, the intersection rules (C.2) for the associated tangent sets hold.
C.1 $\mathcal{H}$ as an affine manifold
Consider the case $\mathcal{H}=\mathrm{Aff}(m,n)=\{X\in\mathbb{R}^{m\times n}\,|\ \mathcal{A}(X)-b={0}\}$. Let $A_{1},A_{2},\ldots,A_{q}\in\mathbb{R}^{m\times n}$ be the matrices constituting the mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to\mathbb{R}^{q}$, i.e., $\mathcal{A}(X)_{i}=\langle{A_{i},X}\rangle$ for $i=1,2,\ldots,q$. Given $X\in\mathcal{M}_{\leq r}$ with $\operatorname{rank}(X)=s$ and the SVD $X=U\varSigma V^{\top}$, denoting
$T^{i}_{X}=\left[\begin{matrix}U^{\top}A_{i}V&U^{\top}A_{i}V_{\bot}\\ U_{\bot}^{\top}A_{i}V&0\\ \end{matrix}\right]\ \ \text{and}\ \ E^{i}_{X}=U^{\top}A_{i}V\ \ \text{for}\ i=1,2,\ldots,q,$
Li and Luo li2023normalboundedaffine proposed the following constraint qualification.
Assumption 3
(li2023normalboundedaffine, Assumptions 3.3 and 3.4) When $s=r$, the matrices $T^{i}_{X}$, i=1,2,,qi=1,2,...,q, are linearly independent; when $s<r$, the matrices $E^{i}_{X}$, i=1,2,,qi=1,2,...,q, are linearly independent.
Taking into account the LR parameterization $(\mathcal{M}_{\mathrm{LR}},\phi_{\mathrm{LR}})$ in (C.1), we present the preimage of $\mathcal{H}=\mathrm{Aff}(m,n)$ under the mapping $\phi_{\mathrm{LR}}$ as follows,
$\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathcal{H})=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid\langle{A_{i},LR^{\top}}\rangle-b_{i}=0,\,i=1,2,\ldots,q\}.$ (C.3)
In addition, the set $\overline{\mathcal{H}}$ can be characterized as the level set of $\bar{h}:\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\to\mathbb{R}^{q}:\,\bar{h}(L,R)_{i}:=\langle{A_{i},LR^{\top}}\rangle-b_{i},\ \text{for}\ i=1,2,\ldots,q$. In this view, we then show that $\overline{\mathcal{H}}$ is indeed an embedded submanifold in $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ under Assumption 3.
Proposition 12
Suppose that Assumption 3 holds at all $X\in\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)$. The set $\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathrm{Aff}(m,n))$ is a smooth submanifold embedded in $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ of dimension (mr+nrq)(mr+nr-q).
Proof
We compute the differential of $\bar{h}$ below,
$\left({\mathrm{D}\bar{h}_{(L,R)}[\dot{L},\dot{R}]}\right)_{i}=\langle{A_{i},L\dot{R}^{\top}}\rangle+\langle{A_{i},\dot{L}R^{\top}}\rangle,\ \text{for}\ i=1,2,\ldots,q,$ (C.4)
and it suffices to prove that $\operatorname{rank}(\mathrm{D}\bar{h}_{(L,R)})=q$ for any $(L,R)\in\overline{\mathcal{H}}$, according to (lee2012manifolds, Corollary 5.14). To this end, let $X=LR^{\top}$ with the SVD $X=U\varSigma V^{\top}$.
We first consider the case $\operatorname{rank}(X)=r$, which indicates that $(L,R)$ can be written as $(L,R)=(UB,VC)$ for some invertible $B,C\in\mathbb{R}^{r\times r}$. Taking $(\dot{L},\dot{R})=((U\dot{B}_{1}+U_{\bot}\dot{B}_{2})C^{-\top},(V\dot{C}_{1}+V_{\bot}\dot{C}_{2})B^{-\top})$ in (C.4) yields
$\mathrm{D}\bar{h}_{(L,R)}[\dot{L},\dot{R}]_{i}=\Big\langle T^{i}_{X},\left[\begin{matrix}\dot{C}_{1}^{\top}+\dot{B}_{1}&\dot{C}_{2}^{\top}\\ \dot{B}_{2}&0\\ \end{matrix}\right]\Big\rangle\ \ \text{for}\ i=1,2,\ldots,q.$
The linear independence of $T^{i}_{X}$ and the arbitrariness of $\dot{B}_{1},\dot{C}_{1}\in\mathbb{R}^{r\times r}$, $\dot{B}_{2}\in\mathbb{R}^{(m-r)\times r}$, $\dot{C}_{2}\in\mathbb{R}^{(n-r)\times r}$ reveal that $\operatorname{rank}\mathrm{D}\bar{h}_{(L,R)}=q$.
The second case is $\operatorname{rank}(X)=s<r$. We parameterize $(L,R)$ by $(L,R)=(UB_{1}+U_{\bot}B_{2},VC_{1}+V_{\bot}C_{2})$, and the SVD of $X$ implies that
$U\varSigma V^{\top}=LR^{\top}=UB_{1}C^{\top}_{1}V^{\top}+UB_{1}C_{2}^{\top}V_{\bot}^{\top}+U_{\bot}B_{2}C_{1}^{\top}V^{\top}+U_{\bot}B_{2}C_{2}^{\top}V_{\bot}^{\top}.$
Therefore, we have $B_{1}C_{1}^{\top}=\varSigma$, $B_{1}C^{\top}_{2}=0$, $B_{2}C_{1}^{\top}=0$, and $B_{2}C^{\top}_{2}=0$. Taking $(\dot{L},\dot{R})=(0,V\dot{C}\varSigma^{-1}C_{1})$ with $\dot{C}\in\mathbb{R}^{s\times s}$ in (C.4) yields
$\mathrm{D}\bar{h}_{(L,R)}[\dot{L},\dot{R}]_{i}=\langle{A_{i},L\dot{R}^{\top}}\rangle=\langle{A_{i},U\dot{C}^{\top}V^{\top}}\rangle=\langle{E^{i}_{X},\dot{C}^{\top}}\rangle,\ \text{for}\ i=1,2,\ldots,q.$
The linear independence of $E^{i}_{X}$ and the arbitrariness of $\dot{C}\in\mathbb{R}^{s\times s}$ indicate that $\operatorname{rank}\mathrm{D}\bar{h}_{(L,R)}=q$.
The above proof concludes that the differential of $\bar{h}$ has full rank $q$ in the level set $\overline{\mathcal{H}}$, and thus applying Corollary 3 directly gives (C.2) with $\mathcal{H}=\mathrm{Aff}(m,n)$.
We then delve into the closed-form formula of the tangent cone at a point $X\in\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)$ when $\operatorname{rank}(X)=s$. Recalling (2.13), any $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}}(X)$ can be parameterized as $\eta=UW_{1}V^{\top}+UW_{2}V^{\top}_{\bot}+U_{\bot}W_{3}V^{\top}+U_{\bot}JV^{\top}_{\bot}$ with $\operatorname{rank}(J)\leq r-s$. If, additionally, $\eta$ belongs to $\mathrm{T}_{\mathrm{Aff}(m,n)}(X)$, i.e., $\langle{A_{i},\eta}\rangle=0$ for $i=1,2,\ldots,q$, we have $\langle{U^{\top}A_{i}V,W_{1}}\rangle+\langle{U^{\top}A_{i}V_{\bot},W_{2}}\rangle+\langle{U_{\bot}^{\top}A_{i}V,W_{3}}\rangle+\langle{U_{\bot}^{\top}A_{i}V_{\bot},J}\rangle=0$, and then substitute the expression of $T^{i}_{X}$ to obtain
$\Big\langle T^{i}_{X},\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&0\\ \end{matrix}\right]\Big\rangle+\langle{U_{\bot}^{\top}A_{i}V_{\bot},J}\rangle=0\ \ \text{for}\ i=1,2,\ldots,q.$
Therefore, the tangent cone to $\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)$ admits the following characterizations,
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)}(X)$
$=$ $\ \mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathrm{Aff}(m,n)}(X)$
$=$ $\left\{{\left[U\,\,U_{\bot}\right]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right]\left[V\,\,V_{\bot}\right]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times\left(n-s\right)},\\ W_{3}\in\mathbb{R}^{\left(m-s\right)\times s},\,J\in\mathbb{R}^{\left(m-s\right)\times\left(n-s\right)},\\ \operatorname{rank}(J)\leq r-s,\\[2.84526pt] \Big\langle T^{i}_{X},\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&0\\ \end{matrix}\right]\Big\rangle+\langle{U_{\bot}^{\top}A_{i}V_{\bot},J}\rangle=0,i\in[q]\end{array}\right.}\right\},$ (C.9)
where we denote $q]=\{1,2,\ldots,q\}$. Taking the polar operation on the above equality yields the Fréchet normal cone as a byproduct,
$\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)}(X)=\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}}(X)+\hat{\mathrm{N}}_{\mathrm{Aff}(m,n)}(X),$ (C.10)
where $\hat{\mathrm{N}}_{\mathcal{M}_{\leq r}}(X)$ is presented in (2.14) and $\hat{\mathrm{N}}_{\mathrm{Aff}(m,n)}(X)=\{\sum_{i=1}^{q}c_{i}A_{i}\mid c_{i}\in\mathbb{R},\,i\in[q]\}$. We remark that normal cone (C.10) recovers (li2023normalboundedaffine, Theorem 3.7), and the developed tangent cone (C.9) serves as a new result.
Finally, given any $\eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)}(X)$, we note that $\mathrm{T}^{2}_{\mathrm{Aff}(m,n)}(X;\eta)=\mathrm{T}_{\mathrm{Aff}(m,n)}(X)$ by definition (2.2) and the affine structure of $\mathrm{Aff}(m,n)$. Hence, an explicit characterization of $\mathrm{T}^{2}_{\mathcal{M}_{\leq r}\cap\mathrm{Aff}(m,n)}(X;\eta)$ is attainable by intersecting the result in (3.6) with the fixed subspace $\mathrm{T}_{\mathrm{Aff}(m,n)}(X)$.
C.2 $\mathcal{H}$ as an orthogonally invariant manifold
Yang et al. yang2025spacedecouple considered $\mathcal{M}_{\leq r}\cap\mathcal{H}$ with the $\mathcal{H}$ encompassing $\mathrm{S_{F}}(m,n)$ and $\mathrm{Ob}(m,n)$ as specific instances. They resorted to the concept of “orthogonal invariance” as follows.
Assumption 4
(yang2025spacedecouple, Assumption 1) The set $\mathcal{H}$ is the level set of a smooth and orthogonally invariant mapping $h:\,\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{q}$ in the sense that
$\mathcal{H}=\{X\in\mathbb{R}^{m\times n}\mid h(X)=0\},\ \ \text{and}\ \ h(XQ)=h(X),\ \text{for all}\ Q\in\mathcal{O}(n).$ (C.11)
Moreover, $h$ has full rank $q$ in $\mathcal{H}$, i.e., $\operatorname{rank}(\mathrm{D}h_{X})=q$ for all $X\in\mathcal{H}$.
We utilize the parameterization $(\mathcal{M}_{\mathrm{LR}},\phi_{\mathrm{LR}})$ to lift $\mathcal{H}$ through the mapping $\phi_{\mathrm{LR}}$:
$\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathcal{H})=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid\bar{h}(L,R):=h(LR^{\top})=0\}.$ (C.12)
Then, the aim is to show that $\mathrm{D}\bar{h}$ has full rank in the level set $\overline{\mathcal{H}}$, and thus $\overline{\mathcal{H}}$ is an embedded submanifold; see the following proposition.
Proposition 13
Suppose that $\mathcal{H}$ satisfies Assumption 4. The set $\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathcal{H})$ is a smooth submanifold embedded in $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ of dimension (mr+nrq)(mr+nr-q).
Proof
The differential of $\bar{h}$ can be computed as
$\mathrm{D}\bar{h}_{(L,R)}[\dot{L},\dot{R}]=\mathrm{D}h_{X}[L\dot{R}^{\top}]+\mathrm{D}h_{X}[\dot{L}R^{\top}],$ (C.13)
where we denote $X=LR^{\top}$. Suppose that $\operatorname{rank}(X)=s\leq r$ and let the SVD of $X$ be $X=U\varSigma V^{\top}$. We parameterize $(L,R)$ by $(L,R)=(UB_{1}+U_{\bot}B_{2},VC_{1}+V_{\bot}C_{2})$, and the SVD of $X$ implies that
$U\varSigma V^{\top}=LR^{\top}=UB_{1}C^{\top}_{1}V^{\top}+UB_{1}C_{2}^{\top}V_{\bot}^{\top}+U_{\bot}B_{2}C_{1}^{\top}V^{\top}+U_{\bot}B_{2}C_{2}^{\top}V_{\bot}^{\top}.$
Therefore, it holds that $B_{1}C_{1}^{\top}=\varSigma$ and $B_{1}C_{2}^{\top}=0$. Moreover, using (yang2025spacedecouple, Proposition 1) shows that $\{BV^{\top}_{\bot}\mid B\in\mathbb{R}^{m\times(n-s)}\}\subseteq\operatorname{ker}(\mathrm{D}{h}_{X})$. This observation, together with the full-rankness of $\mathrm{D}h_{X}$, reveals that any $b\in\mathbb{R}^{q}$ admits a preimage $\eta$ in the form of $\eta=\tilde{B}V^{\top}\in\mathbb{R}^{m\times n}$ for some $\tilde{B}\in\mathbb{R}^{m\times s}$, i.e., $\mathrm{D}h_{X}[\tilde{B}V^{\top}]=b$. Taking $(\dot{L},\dot{R})=(\tilde{B}\varSigma^{-1}B_{1},0)$ in (C.13) shows that $\mathrm{D}\bar{h}_{(L,R)}[\dot{L},\dot{R}]=b$. Consequently, the arbitrariness of $b\in\mathbb{R}^{q}$ implies that $\mathrm{D}\bar{h}$ has the full rank $q$ in $\overline{\mathcal{H}}$, which completes the proof by invoking (lee2012manifolds, Corollary 5.14).
Consequently, applying Corollary 3 to $\mathcal{M}_{\leq r}\cap\mathcal{H}$ yields the intersection rules for the tangent sets in (C.2). Motivated by this result, we now derive the closed-form expressions for the cases $\mathcal{H}=\mathrm{S_{F}}(m,n)$ and $\mathcal{H}=\mathrm{Ob}(m,n)$. Specifically, we note that $\mathrm{T}_{\mathrm{S_{F}}(m,n)}(X)=\{\eta\in\mathbb{R}^{m\times n}\mid\mathrm{tr}(X^{\top}\eta)=0\}$, and thus
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathrm{S_{F}}(m,n)}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathrm{S_{F}}(m,n)}(X)$
$=\left\{{\left[U\,\,U_{\bot}\right]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right]\left[V\,\,V_{\bot}\right]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times\left(n-s\right)},\\ W_{3}\in\mathbb{R}^{\left(m-s\right)\times s},\\ J\in\mathbb{R}^{\left(m-s\right)\times\left(n-s\right)},\\ \operatorname{rank}(J)\leq r-s,\,\mathrm{tr}(\varSigma W_{1})=0\end{array}\right.}\right\}$.
Similarly, from $\mathrm{T}_{\mathrm{Ob}(m,n)}(X)=\{\eta\in\mathbb{R}^{m\times n}\mid\operatorname{diag}(X\eta^{\top})=0\}$, we obtain
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathrm{Ob}(m,n)}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathrm{Ob}(m,n)}(X)$
$=\left\{{\left[U\,\,U_{\bot}\right]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right]\left[V\,\,V_{\bot}\right]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times\left(n-s\right)},\\ W_{3}\in\mathbb{R}^{\left(m-s\right)\times s},\\ J\in\mathbb{R}^{\left(m-s\right)\times\left(n-s\right)},\\ \operatorname{rank}(J)\leq r-s,\\ \operatorname{diag}(U\varSigma(W_{1}^{\top}U^{\top}+W_{3}^{\top}U_{\bot}^{\top}))\!=\!0,\!\!\!\end{array}\right.}\right\}$
$=\left\{{\left[U\,\,U_{\bot}\right]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right]\left[V\,\,V_{\bot}\right]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times\left(n-s\right)},\\ W_{3}\in\mathbb{R}^{\left(m-s\right)\times s},\\ J\in\mathbb{R}^{\left(m-s\right)\times\left(n-s\right)},\\ \operatorname{rank}(J)\leq r-s,\\ UW_{1}+U_{\bot}W_{3}\in\mathrm{T}_{\mathrm{Ob}(m,s)}(U\varSigma)\end{array}\right.}\right\}$.
The explicit formulas of the tangent cones recover the results in (cason2013iterative, Theorem 6.1) and (yang2025spacedecouple, Theorem 1).
C.3 $\mathcal{H}$ as the product of hyperbolic manifolds
Let $\mathcal{L}=\mathrm{Diag}(-1,1,\ldots,1)\in\mathbb{R}^{m\times m}$. For $x,y\in\mathbb{R}^{m}$, we define the Lorentzian inner product as $\langle x,y\rangle_{\mathcal{L}}:=x^{\top}\mathcal{L}y=-x_{1}y_{1}+\sum_{i=2}^{m}x_{i}y_{i}$. Then, we consider the upper sheet of an (m1)(m-1)-dimensional hyperboloid to define hyperbolic manifold:
$\mathbb{H}_{m-1}=\{x\in\mathbb{R}^{m}\mid\langle x,x\rangle_{\mathcal{L}}=-1,\ x_{1}>0\}.$
Stacking $n$ vectors in $\mathbb{H}_{m-1}$ gives rise to the product manifold,
$\mathbb{H}^{n}_{m-1}=\{X\in\mathbb{R}^{m\times n}\mid X_{i}\in\mathbb{H}_{m-1}\ \text{for}\ i=1,2,\ldots,n\},$ (C.14)
where $X_{i}$ extracts the $i$-th column of $X$.
The hyperbolic manifold is a smooth manifold with negative constant curvature, and it has attracted recent interest in the machine learning community for learning hyperbolic embeddings of entities nickel2018learninghierarchieslorentz. For computational efficiency, Jawanpuria et al. jawanpuria2019lowrankhyperbolic proposed learning hyperbolic embeddings within a latent low-dimensional subspace. Specifically, they searched for a low-rank matrix in $\mathbb{R}^{m\times n}$ with columns encoding (m1)(m-1)-dimensional hyperbolic embeddings corresponding to $n$ data points, which, in turn, motivates our study on the geometry of the feasible region $\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}$.
To align with the spirit of Assumption 2, we then treat $\mathbb{H}^{n}_{m-1}$ as the zero set of a sequence of functions $h_{i}:\mathbb{R}^{m\times n}\to\mathbb{R}:\,X\mapsto\langle{X_{i},X_{i}}\rangle_{\mathcal{L}}+1$ ($i=1,2,\ldots,n$), that is, $\mathbb{H}^{n}_{m-1}=\{X\in\mathbb{R}^{m\times n}\mid h_{i}(X)=0,\,X_{i,1}>0\ \text{for}\ i=1,2,\ldots,n\}$, where $X_{i,1}$ denotes the first element of $X_{i}$. Let $h:=(h_{1},h_{2},\ldots,h_{n})$, and a direct computation tells that $\mathrm{D}h$ has full rank on $\mathbb{H}^{n}_{m-1}$.
Furthermore, define the smooth functions $\bar{h}_{i}:\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\to\mathbb{R}$ by
$\bar{h}_{i}(L,R):=\langle{(LR^{\top})_{i},(LR^{\top})_{i}}\rangle_{\mathcal{L}}+1,\ \text{for}\ i=1,2,\dots,n,$
and the mapping $\bar{h}:=(\bar{h}_{1},\bar{h}_{2}\dots,\bar{h}_{n})$. We note that $\bar{h}=h\circ\phi_{\mathrm{LR}}$, and are interested in the preimage of $\mathcal{H}=\mathbb{H}^{n}_{m-1}$ through the smooth mapping $\phi_{\mathrm{LR}}$:
$\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathbb{H}^{n}_{m-1})=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid\bar{h}(X)=0\}\cap\mathcal{Z},$ (C.15)
where $\mathcal{Z}:=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid(LR^{\top})_{i,1}>0\ \text{for}\ i=1,2,\ldots,n\}$. Notice that $\mathcal{Z}$ is an open subset, and the following proposition reveals that $\overline{\mathcal{H}}$ is a smooth manifold.
Proposition 14
The set $\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathbb{H}^{n}_{m-1})$ is a smooth submanifold embedded in $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ of dimension (mr+nrn)(mr+nr-n).
Proof
According to (lee2012manifolds, Corollary 5.14) and the characterization (C.15), it suffices to prove that $\operatorname{rank}(\mathrm{D}\bar{h}_{(L,R)})=n$ for every $(L,R)\in\overline{\mathcal{H}}$. In preparation, write $R^{i}\in\mathbb{R}^{1\times r}$ for the $i$-th row of $R$ and note that $x_{i}=LR^{i\top}\in\mathbb{R}^{m}$. We then compute the gradients of $\bar{h}_{i}$ with respect to $R$ as follows:
$(\nabla_{R}\bar{h}_{i})^{i}$ $=2x_{i}^{\top}\mathcal{L}L\in\mathbb{R}^{1\times r},$ (C.16)
$(\nabla_{R}\bar{h}_{i})^{j}$ $=0\ \ \text{for}\ j=1,\ldots,i-1,i+1,\ldots,n,$
where we use the superscripts to extract the corresponding rows from a matrix. Moreover, $\nabla_{L}\bar{h}_{i}=2\mathcal{L}x_{i}R^{i}\in\mathbb{R}^{m\times r}$. Hence, given a point $(L,R)\in\overline{\mathcal{H}}$, we have $x_{i}^{\top}\mathcal{L}x_{i}=(LR^{i\top})^{\top}\mathcal{L}(LR^{i\top})=R^{i}(L^{\top}\mathcal{L}L)R^{i\top}=-1$, and thus it holds that
$(\nabla_{R}\bar{h}_{i})^{i\top}=2L^{\top}\mathcal{L}x_{i}=2(L^{\top}\mathcal{L}L)R^{i\top}\neq 0,\ \ \text{for}\ i=1,2,\ldots,n.$ (C.17)
We now show that the differentials of $\bar{h}_{1},\bar{h}_{2}\dots,\bar{h}_{n}$ are linearly independent, for which it suffices to consider only the $\nabla_{R}\bar{h}_{i}$. In detail, the computation (C.16) reveals that each $\nabla_{R}\bar{h}_{i}$ is supported exclusively on the $i$-th row. Therefore, if there exists $\alpha\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{n}\alpha_{i}\,\nabla_{R}\bar{h}_{i}=0$, then the $i$-th row of the sum equals $2\alpha_{i}x_{i}^{\top}\mathcal{L}L=0$, which forces $\alpha_{i}=0$ for all $i$ by (C.17). This concludes that $\{\nabla_{R}\bar{h}_{i}\}_{i=1}^{n}$ are linearly independent, and thus the full gradients $\{(\nabla_{L}\bar{h}_{i},\nabla_{R}\bar{h}_{i})\}_{i=1}^{n}$ are also linearly independent. In other words, we have $\operatorname{rank}(\mathrm{D}\bar{h}_{(L,R)})=n$ on $\overline{\mathcal{H}}$.
The above proof shows that $\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}$ falls within the scope of Theorem 4.1—or more specifically, Corollary 3—and therefore we obtain
$\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}}(X)$ $=\mathrm{T}_{\mathcal{M}_{\leq r}}(X)\cap\mathrm{T}_{\mathbb{H}^{n}_{m-1}}(X)$
$=\left\{{\eta=\left[U\,\,U_{\bot}\right]\left[\begin{matrix}W_{1}&W_{2}\\ W_{3}&J\\ \end{matrix}\right]\left[V\,\,V_{\bot}\right]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathbb{R}^{s\times s},\,W_{2}\in\mathbb{R}^{s\times\left(n-s\right)},\\ W_{3}\in\mathbb{R}^{\left(m-s\right)\times s},\\ J\in\mathbb{R}^{\left(m-s\right)\times\left(n-s\right)},\\ \operatorname{rank}(J)\leq r-s,\\ \operatorname{diag}(X^{\top}\mathcal{L}\eta)=0\end{array}\right.}\right\},$
where the second equality holds by substituting the formula $\mathrm{T}_{\mathbb{H}^{n}_{m-1}}(X)=\{\eta\in\mathbb{R}^{m\times n}\mid\operatorname{diag}(X^{\top}\mathcal{L}\eta)={0}\}$ (see nickel2018learninghierarchieslorentz). Similarly, the intersection rule for the second-order tangent set also holds:
$\mathrm{T}^{2}_{\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{M}_{\leq r}}(X;\eta)\cap\mathrm{T}^{2}_{\mathbb{H}^{n}_{m-1}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{M}_{\leq r}\cap\mathbb{H}^{n}_{m-1}}(X).$
Appendix D Tangent sets to low-rank symmetric sets
This section serves as a supplement to sections 4.2.2-4.2.3, and the organization is outlined as follows. Appendix D.1 applies Theorem 3.1 to $\mathcal{S}_{j}$ for $j\in\{1,2,\ldots,r+1\}$, characterizing the corresponding tangent sets. Appendix D.2 then aggregates these results to derive the tangent sets to $\mathcal{S}(n)$. Finally, in Appendix D.3, we identify $\mathcal{S}^{+}_{\leq r}(n)$ with $\mathcal{S}_{r+1}$, and thus translate the obtained results on $\mathcal{S}_{r+1}$ to yield the tangent sets to $\mathcal{S}^{+}_{\leq r}(n)$; moreover, applying Theorem 4.1 derives the tangent sets to $\mathcal{S}^{+}_{\leq r}\cap\mathcal{U}$ when $\mathcal{U}$ is an affine set, a case that has garnered increasing interest in recent work boumal2020deterministic; levin2025effect.
We first verify that each $\mathcal{S}_{j}$ satisfies Assumption 1, which is defined as follows,
$\mathcal{S}_{j}=\{X\in\mathcal{S}(n)\mid\lambda_{j}(X)=0,\,\lambda_{j+n-r-1}(X)=0\},\ \ \text{for}\ j=1,2,\ldots,r+1.$ (D.1)
[…]
$\bar{\varSigma}_{P}(i,i)=\begin{cases}\max\{0,\lambda_{i}(\tilde{X})\},\ \ \ \ \ \text{if}\ 1\leq i\leq j-1,\\ 0,\hskip 69.70924pt\ \text{if}\ j\leq i\leq j+n-r-1,\\ \min\{0,\lambda_{i}(\tilde{X})\},\ \ \ \ \ \ \text{if}\ j+n-r\leq i\leq n.\end{cases}$ (D.2)
Therefore, we have
$\!\!\!\operatorname{dist}(\tilde{X},\mathcal{S}_{j})^{2}\leq\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}^{2}=$ $\!\ \sum_{i=1}^{j-1}|\lambda_{i}(\tilde{X})-\max\{0,\lambda_{i}(\tilde{X})\}|^{2}+\!\sum_{i=j}^{j+n-r-1}\lambda_{i}(\tilde{X})^{2}\!\!\!$ (D.3)
$+\sum_{i=j+n-r}^{n}|\lambda_{i}(\tilde{X})-\min\{0,\lambda_{i}(\tilde{X})\}|^{2}.$
If $\lambda_{j}(\tilde{X})\geq 0$, it holds that $\sum_{i=1}^{j-1}|\lambda_{i}(\tilde{X})-\max\{0,\lambda_{i}(\tilde{X})\}|^{2}=0$. Otherwise, if $\lambda_{j}(\tilde{X})<0$, there exists an index $e$ ($1\leq e\leq j$) such that $e$ is the smallest index with $\lambda_{e}(\tilde{X})<0$, implying that
$\sum_{i=1}^{j-1}|\lambda_{i}(\tilde{X})-\max\{0,\lambda_{i}(\tilde{X})\}|^{2}=\sum_{i=e}^{j-1}\lambda_{i}(\tilde{X})^{2}\leq n|\lambda_{j}(\tilde{X})|^{2}.$ (D.4)
[…]
$\sum_{i=j+n-r}^{n}|\lambda_{i}(\tilde{X})-\min\{0,\lambda_{i}(\tilde{X})\}|^{2}\leq n|\lambda_{j+n-r-1}(\tilde{X})|^{2}.$ (D.5)
Consequently, we collect the estimates (D.3), (D.4), and (D.5) to obtain
$\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}^{2}$ $\leq n|\lambda_{j}(\tilde{X})|^{2}+\sum_{i=j}^{j+n-r-1}\lambda_{i}(\tilde{X})^{2}+n|\lambda_{j+n-r-1}(\tilde{X})|^{2}$
$\leq 3n\max\{|\lambda_{j}(\tilde{X})|^{2},|\lambda_{j+n-r-1}(\tilde{X})|^{2}\},$ (D.6)
which reveals that $\mathcal{S}_{j}$ given by (D.1) satisfies the error bound condition. Additionally, Weyl’s inequality weyl1912inequality indicates the Lipschitz continuity of the mapping $)\tilde{X}\mapsto(\lambda_{j}(\tilde{X}),\lambda_{j+n-r-1}(\tilde{X}$)), which justifies the application of Theorem 3.1 to $\mathcal{S}_{j}$.
D.1 Tangent sets to $\mathcal{S}_{j}$
Applying Theorem 3.1 to $\mathcal{S}_{j}$ reduces to finding the zeros of the directional derivatives of eigenvalue mappings. To this end, we note that the explicit expressions for the directional derivatives of $\lambda_{i}$ ($i=1,2,\ldots,n$) are provided in torki2001secondtoeigen; zhang2013secondordersingular. In preparation, we recall some necessary notation from zhang2013secondordersingular and section 3.2.
[…]
$X=\bar{U}\bar{\Lambda}\bar{U}^{\top},$
where $\bar{U}\in\mathcal{O}(n)$ and $\bar{\Lambda}=\operatorname{Diag}(\lambda_{1}(X),\ldots,\lambda_{n}(X))$. We slightly abuse the notation of $\mu$} and $\alpha_{k}$ by adapting them to this symmetric scenario. In detail, suppose that $X$ admits $t$ distinct eigenvalues, $\mu_{1}>\mu_{2}>\ldots>\mu_{t}$, based on which we categorize the index set $\{1,2,\ldots,n\}$ into $\alpha_{k}=\left\{i\mid\lambda_{i}(X)=\mu_{k},1\leq i\leq n\right\}$ for $k=1,\ldots,t$. Moreover, the definitions of $N_{k}$ and $\beta_{j}^{k}$ for $j=1,2,\ldots,N_{k}$ follow those in section 3.2 when $m=n$. The index mappings also require a minor adaptation as follows:
$q_{a}$ $:\{1,\ldots,n\}\rightarrow\{1,\ldots,t\},\ q_{a}(i)=k,\text{ if }i\in\alpha_{k},$ (D.7)
$l$ $:\{1,\ldots,n\}\rightarrow\mathbb{N},\ l(i)=i-\kappa_{q_{a}(i)-1},$
$q_{b}$ $:\{1,\ldots,n\}\rightarrow\mathbb{N},\ q_{b}(i)=e,\ \text{if}\ l(i)\in\beta_{e}^{q_{a}(i)},$
$l^{\prime}$ $:\{1,\ldots,n\}\rightarrow\mathbb{N},\ l^{\prime}(i)=l(i)-\kappa_{q_{b}(i)-1}^{(q_{a}(i))},$
where $\kappa_{i}:=\sum_{j=1}^{i}|\alpha_{j}|$ and $\kappa_{i}^{(k)}:=\sum_{j=1}^{i}|\beta_{j}^{k}|$. We then outline the results of (zhang2013secondordersingular, Theorem 2.1). Given $i\in\alpha_{k}$ and $\eta\in\mathcal{S}(n)$, the first-order directional derivative of $\lambda_{i}$ at $X$ along $\eta$ can be computed by
$\lambda^{\prime}_{i}(X;\eta)=\lambda_{l(i)}(\bar{\eta}_{\alpha_{k}\alpha_{k}}),$ (D.8)
where $\bar{\eta}_{\alpha_{k}\alpha_{k}}=\bar{U}_{\alpha_{k}}^{\top}\eta\bar{U}_{\alpha_{k}}$. Given, in addition, a direction $\zeta\in\mathcal{S}(n)$, the second-order directional derivative is expressed by
[…]
where $Q^{k}\in\mathcal{O}^{|\alpha_{k}|}(\bar{\eta}_{\alpha_{k}\alpha_{k}})$, and the subscript $\beta^{k}_{q_{b}(i)}$ extracts the columns of $Q^{k}$ indexed by the set $\beta^{k}_{q_{b}(i)}$.
We now proceed to derive the tangent sets to $\mathcal{S}_{j}$ by determining the zeros of $\lambda^{\prime}$ and $\lambda^{\prime\prime}$ given in (D.8) and (D.9), respectively; see the following proposition.
Proposition 15
Given $j\in\{1,2,\ldots,r+1\}$, $X\in\mathcal{S}_{j}$ with $\operatorname{rank}{(X)}=s$, and the spectral decomposition $X=U\varLambda U^{\top}$ with $U\in\mathrm{St}(n,s)$. Let $s_{+}\leq s$ count the positive eigenvalue of $X$. The tangent cone to $\mathcal{S}_{j}$ can be characterized by
$\mathrm{T}_{\mathcal{S}_{j}}(X)=\left\{{[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&J\\ \end{matrix}\right][U\ U_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(s),\\ W_{2}\in\mathbb{R}^{s\times(n-s)},\\ J\in\mathcal{S}(n-s),\\ \lambda_{j-s_{+}}(J)=0,\\ \lambda_{j+n-r-1-s_{+}}(J)=0\end{array}\right.}\right\}$. (D.10)
Additionally, given a direction $\eta\in\mathrm{T}_{\mathcal{S}_{j}}(X)$ parameterized in the above manner with $\operatorname{rank}(J)=\ell-s$ for some $s\leq\ell\leq r$. Let $\ell_{+}\leq\ell-s$ count the positive eigenvalue of $J$. Let the spectral decomposition of $U_{\bot}JU^{\top}_{\bot}$ be $U_{\bot}JU^{\top}_{\bot}=U_{\eta}\varSigma_{\eta}U_{\eta}^{\top}$ with $U_{\eta}\in\mathrm{St}(n,\ell-s)$. Take ${U}_{\eta\bot}$ such that $U\ U_{\eta}\ U_{\eta\bot}]\in\mathcal{O}(n)$, and denote $U^{+}=[U\ U_{\eta}]$. It holds that
$\mathrm{T}^{2}_{\mathcal{S}_{j}}(X;\eta)=\left\{{2\eta X^{\dagger}\eta+[U^{+}\ U_{\eta\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&L\\ \end{matrix}\right][U^{+}\ U_{\eta\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(\ell),\\ W_{2}\in\mathbb{R}^{\ell\times(n-\ell)},\\ L\in\mathcal{S}(n-\ell),\\ \lambda_{j-s_{+}-\ell_{+}}(L)=0,\\ \lambda_{j+n-r-1-s_{+}-\ell_{+}}(L)=0\end{array}\right.}\right\}$. (D.11)
Proof
It suffices to apply Theorem 3.1 to $\mathcal{S}_{j}(n)$, and then to obtain the corresponding tangent sets as the zeros of the directional derivatives of the eigenvalue mappings $)\tilde{X}\mapsto(\lambda_{j}(\tilde{X}),\lambda_{j+n-r-1}(\tilde{X}$)). In fact, computing the zeros of (D.8) and (D.9) proceeds in parallel with the analysis in sections 3.2 and 3.3.
The condition $X\in\mathcal{S}_{j}$ implies that $\lambda_{j}({X})=\lambda_{j+n-r-1}({X})=0$, and thus the indices $q_{a}(j)=q_{a}(j+n-r-1)$ point to the zero eigenvalue of $X$. Therefore, taking $i=j$ in (D.7) yields the index set $\alpha_{k}$ with $k:=q_{a}(j)$, which corresponds to the set of all zero eigenvalues of $X$. Hence we have $\bar{\eta}_{\alpha_{k}\alpha_{k}}=U_{\bot}^{\top}\eta U_{\bot}$ in this scenario. By definition of the index mapping (D.7), we have $l(j)=j-s_{+}$ and $l(j+n-r-1)=j+n-r-1-s_{+}$, with $s_{+}\leq s$ counting the positive eigenvalues of $X$. By (D.8), the condition $\lambda^{\prime}_{j}(X;\eta)=\lambda_{j+n-r-1}^{\prime}(X;\eta)=0$ is equivalent to
$\lambda_{j-s_{+}}(U_{\bot}^{\top}\eta U_{\bot})=\lambda_{j+n-r-1-s_{+}}(U_{\bot}^{\top}\eta U_{\bot})=0,$ (D.12)
which yields the expression (D.10) by denoting $J=U_{\bot}^{\top}\eta U_{\bot}$.
The derivation of the second-order tangent set (D.11) parallels the proof of Proposition 1, which starts by identifying the values of the associated index mappings. Given the direction $\eta\in\mathrm{T}_{\mathcal{S}_{j}}(X)$ with the associated $\operatorname{rank}(J)=\ell-s$. We notice from (D.12) that the singular value of $J=\bar{\eta}_{\alpha_{k}\alpha_{k}}$ indexed by $l(j)=j-s_{+}$ is zero, and thus $q_{b}(j)$ points to the zero eigenvalue of $J$. Therefore, taking $i=j$ in (D.9) reveals that the index set $\beta_{q_{b}(j)}^{k}$ corresponds to the set of all zero eigenvalues of $J$. Hence we can identify the quantity in (D.9) as $\bar{U}_{\alpha_{k}}Q^{k}_{\beta^{k}_{q_{b}(j)}}=U_{\eta\bot}$ without loss of generality. Moreover, the index $l^{\prime}(j)$ counts the position of $\lambda_{l(j)}(J)$ in the zero eigenvalues of $J$, indicating that $l^{\prime}(j)=l(j)-\ell_{+}=j-s_{+}-\ell_{+}$. A similar analysis can be implemented by letting i=j+nr1i=j+n-r-1 in (D.7) to obtain $l^{\prime}(j+n-r-1)=j+n-r-1-s_{+}-\ell_{+}$. Consequently, by (D.9), the condition $\lambda^{\prime\prime}_{j}(X;\eta,\zeta)=\lambda^{\prime\prime}_{j+n-r-1}(X;\eta,\zeta)=0$ is equivalent to
$\lambda_{j-s_{+}-\ell_{+}}(U_{\eta\bot}^{\top}[\zeta-2\eta X^{\dagger}\eta]U_{\eta\bot})=\lambda_{j+n-r-1-s_{+}-\ell_{+}}(U_{\eta\bot}^{\top}[\zeta-2\eta X^{\dagger}\eta]U_{\eta\bot})=0,$
which yields the expression (D.11) by denoting $L=U_{\eta\bot}^{\top}[\zeta-2\eta X^{\dagger}\eta]U_{\eta\bot}$.
D.2 Tangent sets to $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}$
We first investigate the geometry of $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}$ when $\mathcal{U}$ is the trivial ambient space, i.e., $\mathcal{U}=\mathcal{S}(n)$. To this end, we collect the derived tangent sets to $\mathcal{S}_{j}$ for $j\in\{1,2,\ldots,r+1\}$, thereby obtaining those to $\mathcal{S}_{\leq r}(n)=\bigcup_{j=1}^{r+1}\mathcal{S}_{j}$.
Proof of Proposition 3 The condition $\operatorname{rank}(X)=s\leq r$ implies that $X\in\mathcal{S}_{j}$ if and only if $j\in\mathcal{J}(X):=\{s_{+}+1,\ldots,r+1-s_{-}\}$, where $s_{+}$ and $s_{-}$ count the positive and negative eigenvalues of $X$, respectively. Therefore, we have
$\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X)=\bigcup_{j\in\mathcal{J}(X)}\mathrm{T}_{\mathcal{S}_{j}}(X),$ (D.13)
Enlightened by (D.10), we observe that
$\mathcal{S}_{\leq r-s}(n-s)=\bigcup_{j\in\mathcal{J}(X)}\{J\in\mathcal{S}(n-s)\mid\lambda_{j-s_{+}}(J)=\lambda_{j+n-r-1-s_{+}}(J)=0\},$
which, together with (D.10) and (D.13), produces the expression (4.13).
Furthermore, given the direction $\eta\in\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X)$ parameterized in the form of (4.13) with $\operatorname{rank}(J)=\operatorname{rank}(U^{\top}_{\bot}\eta U_{\bot})=\ell-s$. According to (D.10), it holds that $X\in\mathcal{S}_{j}$ and $\eta\in\mathrm{T}_{\mathcal{S}_{j}}(X)$ if and only if $j\in\mathcal{J}^{\prime}(X;\eta):=\{s_{+}+\ell_{+}+1,\ldots,r+1-s_{-}-\ell_{-}\}$, where $\ell_{+}$ and $\ell_{-}$ count the positive and negative eigenvalues of $J=U^{\top}_{\bot}\eta U_{\bot}\in\mathcal{S}(n-s)$, respectively. Therefore, we have
$\mathrm{T}^{2}_{\mathcal{S}_{\leq r}(n)}(X;\eta)=\bigcup_{j\in\mathcal{J}^{\prime}(X;\eta)}\mathrm{T}^{2}_{\mathcal{S}_{j}}(X;\eta),\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X).$ (D.14)
Then, the relations $_{+}+s_{-}=s, ++=s\ell_{+}+\ell_{-}=\ell-s$, and the observation that
$\bigcup_{j\in\mathcal{J}^{\prime}(X;\eta)}\{L\in\mathcal{S}(n-\ell)\mid\lambda_{j-s_{+}-\ell_{+}}(L)=\lambda_{j+n-r-1-s_{+}-\ell_{+}}(L)=0\}=\mathcal{S}_{\leq r-\ell}(n-\ell)$
conclude (4.14) by substituting (D.11) into (D.14).∎
D.2.1 $\mathcal{U}$ as a Frobenius sphere
We then impose an additional constraint $\mathcal{U}$ on $\mathcal{S}_{\leq r}(n)$ by considering the representative example $\mathcal{U}=\{X\in\mathcal{S}(n)\mid\|X\|_{\mathrm{F}}^{2}-1=0\}$ cason2013iterative; li2020jotaspectral. The main principle is to apply Theorem 3.1 to $\mathcal{S}_{j}\cap\mathcal{U}$, and then assemble the results to obtain the tangent sets of the union $\mathcal{S}_{\leq r}(n)\cap\mathcal{U}=\bigcup_{j=1}^{r+1}(\mathcal{S}_{j}\cap\mathcal{U})$ according to the rule (2.4).
We first verify that each
$\mathcal{S}_{j}\cap\mathcal{U}=\{X\in\mathcal{S}(n)\mid\lambda_{j}(X)=0,\,\lambda_{j+n-r-1}(X)=0,\,h(X)=0\}$
[…]
$\operatorname{dist}(\tilde{X},\mathcal{S}_{j}\cap\mathcal{U})$ $\leq\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}+\|\tilde{X}_{P}-\tilde{X}^{\diamond}_{P}\|_{\mathrm{F}}$
$=\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}+|\|\tilde{X}_{P}\|_{\mathrm{F}}-1|$
$\leq 2\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}+|\|\tilde{X}\|_{\mathrm{F}}-1|,$ (D.15)
where we use the triangle inequality $|\|\tilde{X}_{P}\|_{\mathrm{F}}-1|\leq|\|\tilde{X}\|_{\mathrm{F}}-1|+|\|\tilde{X}\|_{\mathrm{F}}-\|\tilde{X}_{P}\|_{\mathrm{F}}|\leq|\|\tilde{X}\|_{\mathrm{F}}-1|+\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}$. Consequently, we can confirm the error bound property of $\mathcal{S}_{j}\cap\mathcal{U}$ by incorporating (D.6) and $|\|\tilde{X}\|_{\mathrm{F}}-1|\leq|h(X)|$ into (D.15). Applying Theorem 3.1 to $\mathcal{S}_{j}\cap\mathcal{U}$ and taking into account the rule (2.4) show that
$\mathrm{T}_{\mathcal{S}_{\leq r}(n)\cap\mathcal{U}}(X)=\bigcup_{j=1}^{r+1}\mathrm{T}_{\mathcal{S}_{j}\cap\mathcal{U}}(X)=\bigcup_{j=1}^{r+1}(\mathrm{T}_{\mathcal{S}_{j}}(X)\cap\mathrm{T}_{\mathcal{U}}(X))=\mathrm{T}_{\mathcal{S}_{\leq r}(n)}(X)\cap\mathrm{T}_{\mathcal{U}}(X).$
Similarly, given any $\eta\in\mathrm{T}_{\mathcal{S}_{\leq r}(n)\cap\mathcal{U}}(X)$, we have
$\mathrm{T}^{2}_{\mathcal{S}_{\leq r}(n)\cap\mathcal{U}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{S}_{\leq r}(n)}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{U}}(X;\eta).$
D.3 Tangent sets to $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$
We have clarified in section 4.2.3 that the set $\mathcal{S}^{+}_{\leq r}(n)$ coincides with $\mathcal{S}_{r+1}$ defined in (D.1), and thus Proposition 4 directly inherits the results derived in Proposition D.10.
The discussion then proceeds to the intersection $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ with a nontrivial $\mathcal{U}$. Building upon the results of Theorem 4.1, we employ the well-known Burer–Monteiro parameterization burer2003BM:
$(\mathcal{M}_{\mathrm{BM}},\phi_{\mathrm{BM}})=(\mathbb{R}^{n\times r},\,R\mapsto RR^{\top}).$ (D.16)
It is worth noting that $\mathcal{M}_{\mathrm{BM}}$ equals the whole ambient Euclidean space $\mathbb{R}^{n\times r}$. Moreover, $\phi_{\mathrm{BM}}|_{\mathcal{M}_{\mathrm{BM}}}$ is open at all $R\in\mathcal{M}_{\mathrm{BM}}$, according to (levin2025effect, Theorem 2.3, Proposition 2.7). Therefore, Theorem 4.1 can be adapted to the positive semidefinite setting, yielding the following corollary, which is analogous to Corollary 3.
Corollary 4
Suppose that $h:\mathcal{S}(n)\to\mathbb{R}^{q}$ is smooth and $\mathcal{U}=\{\tilde{X}\in\mathcal{S}(n)\mid h(\tilde{X})=0\}$ satisfies Assumption 1 at the point $X\in\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$. Additionally, the differential of the mapping $h\circ\phi_{\mathrm{BM}}$ has constant rank in a neighborhood of
$\overline{\mathcal{U}}:=\{R\in\mathbb{R}^{n\times r}\mid h(\phi_{\mathrm{BM}}(R))=0\}=\phi_{\mathrm{BM}}^{-1}(\mathcal{U}).$
Then, the following intersection rules hold,
$\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X)$ $=\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)}(X)\cap\mathrm{T}_{\mathcal{U}}(X),$ (D.17)
$\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X;\eta)$ $=\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{U}}(X;\eta)\ \ \text{for any}\ \eta\in\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X).$
D.3.1 $\mathcal{U}$ as an affine set
We apply Corollary 4 to $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ when $\mathcal{U}$ is an affine set. In fact, the study of $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ with $\mathcal{U}=\ \{X\in\mathcal{S}(n)\mid\mathcal{A}(X)-b=0\}$ boumal2020deterministic; levin2025effect has been motivated by low-rank SDPs with linear equality constraints, which have witnessed a wide range of applications. Specifically, let $A_{1},A_{2},\ldots,A_{q}\in\mathcal{S}(n)$ be the matrices constituting the linear mapping $\mathcal{A}:\mathcal{S}(n)\to\mathbb{R}^{q}$, i.e., $\mathcal{A}(X)_{i}=\langle{A_{i},X}\rangle$ for $i=1,2,\ldots,q$; and the tangent cone to $\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ was first derived in levin2025effect based on the standard assumption below.
Assumption 5
(boumal2020deterministic, Assumption 1.1) The set $\mathcal{U}=\{X\in\mathcal{S}(n)\mid\mathcal{A}(X)-b=0\}$ defined by $A_{1},A_{2},\ldots,A_{q}\in\mathcal{S}(n)$ satisfies at least one of the following conditions.
  • (i)
    $\{A_{1}R,A_{2}R,\ldots,A_{q}R\}$ are linearly independent in $\mathbb{R}^{n\times r}$;
  • (ii)
    $\{A_{1}R,A_{2}R,\ldots,A_{q}R\}$ span a subspace of constant dimension in $\mathbb{R}^{n\times r}$ for all $R$ in an open neighborhood of $\overline{\mathcal{U}}=\{R\in\mathbb{R}^{n\times r}\mid\mathcal{A}(RR^{\top})-b=0\}$.
Denote $h:\mathcal{S}(n)\to\mathbb{R}^{q}:X\mapsto\mathcal{A}(X)-b$. Under Assumption 5, the analysis in (boumal2020deterministic, Appendix A) shows that the differential of $h\circ\phi_{\mathrm{BM}}$ at $R$ has rank equal to the dimension of the space spanned by $\{A_{1}R,A_{2}R,\ldots,A_{q}R\}$, which validates the application of Corollary 4. Therefore, the intersection rules (D.17) hold with $\mathcal{U}$ as an affine set satisfying Assumption 5.
Finally, given $X\in\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}$ with $\operatorname{rank}{(X)}=s$ and the spectral decomposition $X=U\varLambda U^{\top}$, we present the closed-form formula of the tangent cone:
$\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X)$ $=\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)}(X)\cap\mathrm{T}_{\mathcal{U}}(X)$
$=\left\{{\eta=[U\ U_{\bot}]\left[\begin{matrix}W_{1}&W_{2}\\ W_{2}^{\top}&J\\ \end{matrix}\right][U\ U_{\bot}]^{\top}\left|\,\begin{array}[]{l}W_{1}\in\mathcal{S}(s),\\ W_{2}\in\mathbb{R}^{s\times(n-s)},\\ J\in\mathcal{S}^{+}_{\leq r-s}(n-s),\\ \langle{A_{i},\eta}\rangle=0\ \text{for}\ i\in[q]\end{array}\right.}\right\},$
which recovers the result in (levin2025effect, Corollary 4.12). In addition, given any $\eta\in\mathrm{T}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X)$, we have $\mathrm{T}^{2}_{\mathcal{U}}(X;\eta)=\mathrm{T}_{\mathcal{U}}(X)$ according to definition (2.2) and the affine structure of $\mathcal{U}$. Therefore, the second-order tangent set
$\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)\cap\mathcal{U}}(X;\eta)=\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)}(X;\eta)\cap\mathrm{T}^{2}_{\mathcal{U}}(X;\eta)$
can be characterized by substituting the formula of $\mathrm{T}^{2}_{\mathcal{S}^{+}_{\leq r}(n)}(X;\eta)$ developed in Proposition 4.
Appendix E Proof of Lemma 3
We prove Lemma 3 in section 7.3, borrowing the idea from (olikier2022continuity, Lemma 4.2).
Proof of Lemma 3 Let $\underline{X}_{i}\in\mathcal{P}_{\mathcal{M}_{s}}(X_{i})$ and $\underline{Y}_{i}\in\mathcal{P}_{\mathcal{M}_{k-\ell}}(Y_{i})$. Throughout the proof, we write $P_{i}^{X}:=\underline{X}_{i}\underline{X}_{i}^{\dagger}\in\mathbb{R}^{m\times m}$ and $\Pi_{i}^{X}:=\underline{X}_{i}^{\dagger}\underline{X}_{i}\in\mathbb{R}^{n\times n}$ for the orthogonal projections onto $\operatorname{im}\underline{X}_{i}$ and $\operatorname{im}\underline{X}_{i}^{\top}$, respectively; likewise, $P_{i}^{Y}:=\underline{Y}_{i}\underline{Y}_{i}^{\dagger}$ and $\Pi_{i}^{Y}:=\underline{Y}_{i}^{\dagger}\underline{Y}_{i}$ represent the projections onto $\operatorname{im}\underline{Y}_{i}$ and $\operatorname{im}\underline{Y}_{i}^{\top}$, respectively. Furthermore, we define the orthogonalization of a matrix $Z$ with full column rank by $\operatorname{orth}(Z):=Z(Z^{\top}Z)^{-1/2}$. In the analysis, if needed, we may take a subsequence, still indexed by $i$ for simplicity.
Noticing that $(\underline{X}_{i},\underline{X}_{i}^{\dagger})\to(X,X^{\dagger})$ as $i\to\infty$, we set $U_{i}:=\operatorname{orth}(P_{i}^{X}\,U)\in\mathrm{St}(m,s)$ and $V_{i}:=\operatorname{orth}(\Pi_{i}^{X}\,V)\in\mathrm{St}(n,s)$ to obtain $U_{i}\to U$ and $V_{i}\to V$. Next, complete $U_{i}$ and $V_{i}$ to orthonormal bases of $\operatorname{im}X_{i}$ and $\operatorname{im}X_{i}^{\top}$ by selecting $\tilde{U}_{i\bot}\in\mathrm{St}(m,{\underline{r}}-s)$ and $\tilde{V}_{i\bot}\in\mathrm{St}(n,{\underline{r}}-s)$, respectively, such that $\operatorname{im}[U_{i}\ \tilde{U}_{i\bot}]=\operatorname{im}X_{i}$ and $\operatorname{im}[V_{i}\ \tilde{V}_{i\bot}]=\operatorname{im}X_{i}^{\top}$. By compactness of the Stiefel manifolds, after taking a subsequence if necessary, there exist $\tilde{U}_{\bot}\in\mathrm{St}(m,{\underline{r}}-s)$ and $\tilde{V}_{\bot}\in\mathrm{St}(n,{\underline{r}}-s)$ with $\tilde{U}_{i\bot}\to\tilde{U}_{\bot}$ and $\tilde{V}_{i\bot}\to\tilde{V}_{\bot}$.
In a similar manner, we can set $U_{iY}:=\operatorname{orth}(P_{i}^{Y}\,U_{Y})\in\mathrm{St}(m,k-\ell)$, $V_{iY}:=\operatorname{orth}(\Pi_{i}^{Y}\,V_{Y})\in\mathrm{St}(n,k-\ell)$, and find $\tilde{U}_{iY\bot}\in\mathrm{St}(m,\ell-\bar{r})$ and $\tilde{V}_{iY\bot}\in\mathrm{St}(n,\ell-\bar{r})$ such that $\operatorname{im}[U_{iY}\ \tilde{U}_{iY\bot}]=\operatorname{im}Y_{i}$ and $\operatorname{im}[V_{iY}\ \tilde{V}_{iY\bot}]=\operatorname{im}Y_{i}^{\top}$. Moreover, we have $U_{iY}\ \tilde{U}_{iY\bot}]\to[U_{Y}\ \tilde{U}_{Y\bot}]$ and $V_{iY}\ \tilde{V}_{iY\bot}]\to[V_{Y}\ \tilde{V}_{Y\bot}]$ for some $\tilde{U}_{Y\bot}\in\mathrm{St}(m,\ell-\bar{r})$ and $\tilde{V}_{Y\bot}\in\mathrm{St}(n,\ell-\bar{r})$.
[…]
$\breve{U}_{i\bot}:=\operatorname{orth}\!\big((I_{m}-P_{i}^{X}-\tilde{U}_{i\bot}\tilde{U}_{i\bot}^{\top}-P_{i}^{Y}-\tilde{U}_{iY\bot}\tilde{U}_{iY\bot}^{\top})\,\breve{U}_{\bot}\big)\in\mathrm{St}\!\big(m,\,m-k+\bar{r}-{\underline{r}}\big),$
$\breve{V}_{i\bot}:=\operatorname{orth}\!\big((I_{n}-\Pi_{i}^{X}-\tilde{V}_{i\bot}\tilde{V}_{i\bot}^{\top}-\Pi_{i}^{Y}-\tilde{V}_{iY\bot}\tilde{V}_{iY\bot}^{\top})\,\breve{V}_{\bot}\big)\in\mathrm{St}\!\big(n,\,n-k+\bar{r}-{\underline{r}}\big).$
Using the continuity of the orthogonalization, we obtain $\breve{U}_{i\bot}\to\breve{U}_{\bot}$ and $\breve{V}_{i\bot}\to\breve{V}_{\bot}$ as $i\to\infty$. Note that $\breve{U}_{i\bot}$ are orthogonal to both $\operatorname{im}X_{i}$ and $\operatorname{im}Y_{i}$; similarly, $\breve{V}_{i\bot}$ are orthogonal to both $\operatorname{im}X_{i}^{\top}$ and $\operatorname{im}Y_{i}^{\top}$. Therefore, we have
$U_{i}\ \tilde{U}_{i\bot}\ \breve{U}_{i\bot}\ \tilde{U}_{iY\bot}\ U_{iY}]\in\mathcal{O}(m),\qquad[V_{i}\ \tilde{V}_{i\bot}\ \breve{V}_{i\bot}\ \tilde{V}_{iY\bot}\ V_{iY}]\in\mathcal{O}(n)$
for each $i$. Letting $i\to\infty$ and collecting the limits of each component, we conclude that the sequences satisfy the required properties. ∎